### v1

# SMILES has to go : Representation of Molecules via Algebraic Data Types

Abstract
This paper proposes a novel representation of molecules through Algebraic Data Types (ADTs). The representation has useful properties primarily by including type information. The representation uses the Dietz representation enabling representation of organometallics with multi-centre, multi-atom bonding and delocalised electrons, resonant structures and co-ordinate data of atoms. Furthermore, this representation goes further than any other in the literature, providing a natural data structure to represent shells, subshells and orbitals. Perks of the representation include it’s natural inclusion in reaction descriptions and the ability to make molecules instances of algebraic groups. The representation is further motivated as providing guarantees for those wishing to do Bayesian machine learning (probabilistic programming) over molecular structures. A criticism of competing and commonly used representations such as SMILES and SELFIES is provided and solutions are proposed to the weaknesses of these along with an open source library, written in Haskell. An example of integrating the library with LazyPPL – a lazy probabilistic programming library written in Haskell – is provided, conceptually justifying the efficiency of the representation over string based representations and recent work such as SELFIES [Krenn et al., 2020]. This library distinguishes between the data and the type of data – enabling a separation of concerns between interface and object. I solve three problems associated with the future of SELFIES, molecular programming language, 3D information, syntactic invalidity and Dietz representation.
Keywords: Molecular Representation, Cheminformatics, Algebraic Data Types
1Oxford University, Department of Computer Science, oliver.goldstein@cs.ox.ac.uk
1 Scientific Contribution
  1. 1.
    Introduction of a minimal grammatical representation of molecules. This is combined with detailed criticism of previous molecular representations, along with proposed solutions for each issue identified. Perks of the representation are explored including naturally modelling reactions and group theoretic properties of molecules.
  2. 2.
    Conceptual justification of efficient Bayesian Inference over this representation demonstrated by an example with a Haskell implementation using lazy trace Metropolis Hastings and 3D molecular generation [Paquet and Staton, 2021].
2 Introduction
The choice of molecular representation is consequential for the cheminformatician, and there is an ongoing search for computational representations which produce superior results in machine learning (ML) tasks such as novel molecule discovery, molecular property prediction, and reaction predictions. Usable computational representations which are capable of representing certain types of molecules, or capable of adequately representing reactions in a computational format, are still being searched for. Currently, string based representations are the predominant type of representation used and studied by cheminformatics for such tasks, but they are not without problems.
SMILES [Weininger, 1988] is the most commonly used string based representation, particularly in the case of machine learning [Gómez-Bombarelli et al., 2018], [Zheng et al., 2019] & [Deng et al., 2022]. Its well-known problems include the fact that not all valid SMILES strings are valid molecules. SMILES strings also cannot represent complicated bonding systems (such as in organometallics), resonance structures, any kind of delocalised bonding, and certain kinds of complex stereoisomerism, problems shared with other representations which use molecular graphs. In addition, small changes in SMILES strings can lead to large changes in a molecule’s structure, and adjacent atoms in the molecule are often far apart in the string making mathematical measures of similarity difficult to calculate [Fang et al., 2022],
Different fixes for the problems of string based representations have been attempted or proposed, but these usually require the creation of brand new grammars, of which there is no shortage.
In this paper, we summarise and critique modern computational representations, and proposed representations, specifically with a focus for ML applications. In particular, we critique string based representations and their numerous extensions, including some limitations which are under-discussed by the cheminformatics community, such as the limitations of string based representations in the application of probabilistic programming.
Then, as a solution to these identified problems, we introduce a novel representation of molecules: the Algebraic Data Type (ADT).
Our ADT is written in the pure, type-safe, functional programming language Haskell, and implements a modified Dietz representation, with additional (and optional) molecule properties such as 3D coordinate information and quantum information. It easily represents resonance structures, delocalised bonds and organometallics, and due to the inclusion of quantum information, can be used for quantum chemistry applications.
We also demonstrate how the ADT can be easily extended to represent reaction networks, and we provide a simple example of a probabilistic model which generates molecules and scores them with an arbitrary distance metric for a chosen molecule.
Despite the immediate benefits presented by our ADT, this paper does not does not intend to provide an actual implementation of a sound and complete molecular representation. It merely suggests ADTs as a form of representation, and instead, emphasis is placed on the advantages of ADTs and its potential extensibility and use with probabilistic programming. We leave it to the cheminformatics community to discover the ways in which ADTs would be most useful.
2.1 Types of representation
This paper distinguishes between different usages of ‘molecular representation’.
  1. 1.
    Storage representations refer to the format of storage of raw molecular data, such as molfiles or Structured Data Files (SDF) files [Wigh et al., 2022].
  2. 2.
    Transmission representations refer to how molecules are sent between cheminformaticians and depends on the medium of communication used.
  3. 3.
    Computational representations refer to how a molecule is represented in a programming language. Examples of these are bit-vectors or matrices.
2.2 Modern string representations
As this paper is focused on modern molecular representations and their application for machine learning, the history of molecular representation and an in depth discussion of other representaions is not found here, but can be found in e.g. Krenn et al. [2022] or Wigh et al. [2022].
The types molecular representations considered here are Strings, Graphs, Fingerprints and Variational Auto Encoders, and Computation Tables, each considered in a recent review of molecular representation [Wigh et al., 2022].
2.2.1 Graphs
A molecular graph is defined as “a connected, labelled, and undirected graph. The vertices are atoms, and the edges are bonds. A vertex is labelled with an atom type (chemical element) and possibly the number of unshared electrons; an edge is labelled with a bond order or a bond type (single, double, aromatic, etc.)[Dietz, 1995]
As edges in graphs (as opposed to hypergraphs) are between two vertices only, they are limited in how they can represent delocalised and multi-centre bonding. Although it is possible to label edges with, for instance, rational bond orders, the contribution of individual electrons to multiple covalent bonds would not be described by such a graph. Attempting to express delocalised bonding by labelling edges of a molecular graph with bond types would require a sufficient number of predefined bond types, which would need to be expended for each new bonding relation not described. As Dietz said “Enhancing the expressiveness by including a new bond type for every exceptional case is certainly not a very elegant solution[Dietz, 1995]
Tautomers are unable to be expressed by molecular graphs in a single graph, instead requiring multiple structures to represent the different forms of the molecule.
Furthermore graphs detrimentally interact with tautomers, according to [David et al., 2020] as evidenced by the following: “[For] applications where one is limited to using a single static representation for a molecule that is in fact rearranging on the timescale of the problem (e.g. tautomers), then a single molecular graph representation would not be appropriate and could even be detrimental to solving the problem.”.
2.2.2 Hypergraphs
A hypergraph is a graph where edges may connect any number of vertices.
In a molecular hypergraph, edges with two verticies represent localised bonding, and edges with more than two verticies represent delocalised bonding, and are labelled with the number of electrons shared between the corresponding atoms.
From Dietz:“A hyperedge containing more than two atoms gives us no information about the binary neighborhood relationships between them. That means we have no information at our disposal conceming the way in which the electrons are delocalized over these atoms” [Dietz, 1995]. Hypergraphs would indicate an electronic “soup” with out knowledge of the manner electrons are delocalised. Consequently, they are unsuitable as as a representation.
2.2.3 Dietz representaion
The Dietz representation was designed to overcome limitations of graph based and hypergraph representation. Specifically designed as a computational representation which is as expressive as a structural formulae diagram, and which provides constitutional and stereochemistry features with ease. It has so far seen little use. It uses paddlewheels to represent 3D information, but we ignore this as not relevant so far.
In the paper, it captures diborane and ferrocene with ease, and has a general model for delocalisation, which does not require explicit specificication of bond types such as “covalent" or “aromatic".
2.2.4 SMILES
SMILES (Simplified Molecular Input Line Entry System)[Weininger, 1988] represents the structure of chemical molecules using ASCII strings. Conversion between molecular graphs and text occurs by encoding atoms and bonds with characters and symbols, following a specified grammar to indicate connectivity and structure. Despite known problems, SMILES remains the predominant contemporary representation of molecules[Gómez-Bombarelli et al., 2018], [Semenova et al., 2020] & [Deng et al., 2022]
Not all valid smiles strings correspond to valid molecules. The problem of “syntactic invalidity” means that programmers may parse technically valid smiles strings but fail to produce molecules, and machine learning algorithms can output syntactic nonsense. Machine learning algorithms must re-sample when they encounter syntactic nonsense during the input or output [Gómez-Bombarelli et al., 2018].
In SMILES, small changes in the string can lead to large changes in molecular structure and adjacent atoms in the molecule can be far apart in the string [Fang et al., 2022]. As well as impacting human understanding, this can make it difficult to write or even find a similarity metric of SMILES strings whose magnitude corresponds to a relevant similarity of molecules.
As SMILES strings correspond to molecular graphs, they inherit the weaknesses of graphs as a form of representation. SMILES strings cannot represent delocalised bonds, resonance forms, or tautomerism. Where resonance or tautomerism is represnted, separate structures representing different forms or kukule structures must represented, or chosen between. Ionic bonds, polycentric bonds and metal-metal bonds similarly cannot be represented.
SMILES strings are unable to represent multiple types of stereochemistry. Atropisomers, complex molecules with multiple stereocentres, and non-tetrahedral stereocentres cannot be represented. Without 3D information, it is not possible to represent many forms of stereochemistry, and tasks such as calculating dihedral bond angles are not possible.
Given multiple ways to represent one molecule, several problems arise. The first is that machine learning algorithms need to discern the equality of differing representations. The second is that the symmetry of a representation can be broken. Consider the symmetry of the Cyclopentadienyl anion which becomes [cH-]1cccc1 in SMILES, this does not indicate the rotational symmetry of the molecule. In particular, in SMILES no indication is given as to whether [cH-]1cccc1 = 1cccc1[cH-]. However, the Figure 2 clearly shows rotational symmetry and no signs at all of “one-sidedness”.
Figure 1: An example of a non symmetric SMILES molecule representation known as the Cyclopentadienyl anion: [cH-]1cccc1
Figure 2: Consider the complexity in reading and understanding existing well known molecules such as morphine depicted here [Dietz, 1995]
According to [Wigh et al., 2022], an issue that hasn’t even been considered with Daylight’s SMILES is that its canonicalization algorithm is proprietary, and as such implementations vary between companies and research teams [David et al., 2020]. This caused the community to standardise the implementation with OpenSMILES. Despite recent work which suggests that machine learning models which are able to output nonsensical SMILES strings perform better [Skinnider, 2024], this work has only been tested on SELFIES strings. Furthermore, as a scientific result, there is no analytic proof that these models necessarily perform better and the generality of the result “Models that generate invalid outputs outperform models that do not” is debatable, especially to cheminformaticians who are not in the machine learning community. Indeed even in the machine learning community 3D coordinate data and dihedral angles are often essential [Zhou et al., 2022], [Xu et al., 2022], [Dayalan et al., 2006] and is not included in standard SMILES strings. The title of the article is too strong for the result provided “Invalid SMILES are beneficial rather than detrimental to chemical language models”. Instead we believe the title should be specific and epistemically conservative, suggesting that SMILES strings may be better than SELFIES strings on same tasks, where SELFIES is another representation which this paper criticises. Furthermore, it has nothing to say about the readability of the molecules – and the subsequent performance increases for both bench chemists and analytic chemistry in general.
2.2.5 SELFIES
This work brings together problems with representations that previous literature reviews have failed to notice [David et al., 2020] & [Wigh et al., 2022], such as the overloading of notation or complexity of the CFG with SELFIES, or the breaking of symmetry with SMILES. Previous reviews have also left out [Srinivasan et al., 1996].
SELFIES [Krenn et al., 2020] attempts to solve problems with invalidity of SMILES strings, however, it remains unable to represent molecules with resonant structures and delocalised bonding. Furthermore, it is not easy to extend requiring significant knowledge of a complex and seemingly arbitrary grammar and as such is not easy to extend e.g. incorporating 3D atomic coordinate data. This means there is no way to calculate dihedral angles or bond lengths, which are often critical in machine learning tasks [Schütt et al., 2017]. Furthermore, it also unnaturally introduces the overloading of terms, or dual semantics for a single token, in the language. This will lead to machine learning algorithms requiring more data to discern molecular equality, not to mention difficulties in the readibility of one dimensional strings. Furthermore, there are difficulties in writing or finding a similarity metric that has guaranteed properties, guaranteed by a type and whose magnitude corresponds to a relevant similarity of molecules represented by SMILES strings. This could be problematic in using Variational Auto-encoders (VAEs) based off of SMILES strings to calculate the similarity of a molecule. This is where small changes in the representation significantly changes the properties of the molecule. Similarity metrics in the ADT proposed here can be defined transparently. The ADT also includes coordinate data and indeed, the 3D information is crucial for molecular property prediction [Zhou et al., 2022], [Xu et al., 2022].
SELFIES [Krenn et al., 2020] solves arguably the single most pressing issue with SMILES. They create a “100% ROBUST” representation, which would translate to sound and complete in computer science terminology. In particular, they present a Context Free Grammar (CFG) where all sequences of terminal symbols represent molecules (soundness) and all molecules can be represented (completeness). The CFG of SELFIES can be translated into sum and product notation as every CFG can be represented by an algebraic data type [Syrotiuk, 1984]. In addition, pattern matching, programming with or understanding their CFG would be immensely challenging, because of the significant complexity of the grammar.
Their CFG has undesirable properties from the point of view of readability and also machine learning. The authors overload standard terms in their CFG with numbers after branch or ring statements. This is seen in Figure 3. This could be related to their desire to use SELFIES with machine learning algorithms that mutate and evaluate the resulting molecule, such as genetic algorithms. However, they still wish to preserve semantic coherence under mutation, rather than having syntactic incoherence [^1] In this article, they talk about combining SELFIES with genetic algorithms: https://aspuru.substack.com/p/molecular-graph-representations-and. Figure 3 describes the overridden numeric denotation of a symbol if it appears after a branch or ring token, which breaks the straightforward readability of the grammar. The machine learning algorithm would additionally have to learn this strange pattern itself requiring more data in order to do so, and once again possibly incorrectly generalising.
Figure 3: This is how the overloading takes place. This must expand with the number of elements and types of thing in the SELFIES language. Source: [Krenn et al., 2020]
In addition, resonant structures are not directly expressible in the SELFIES language. Whilst, aromatic rings like benzene are expressible, resonant structures – which involve delocalised electrons over a set of atoms, are not generally representable. Whilst SELFIES does support rudimentary stereochemistry, SELFIES does not represent three dimensional information. It is hard to see how to easily extend the grammar ensuring that it remains to be sound. This is a problem for stereoisomers like cisplatin and transplatin, one being an approved anti-cancer drug, the other having no activity in human subjects. At this point, the soundness and completeness properties become trivial by excluding potentially important data. This could be especially relevant from the point of view of cheminformaticians, who may well need access to tautomers to solve protein-ligand binding affinity [Pospisil et al., 2003].
Furthermore, E-Z stereochemistry does not provide accurate dihedral or torsional angles [Grell et al., 2006]. Proteins rely on these angles for many tasks. In [Dayalan et al., 2006], the importance of this information for protein structure prediction is revealed: “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [Anderson, 2003].
New syntaxes and grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with SMILES. Consider the widely used following syntaxes from just one vendor called Daylight222www.daylight.com. SMILES was introduced and subsequently standardised in OpenSMILES. CurlySMILES was introduced to annotate SMILES strings. CHUCKLES, its extension, CHORTLES, as well as BigSMILES were introduced to represent polymeric molecules [David et al., 2020]. SMIRKS was introduced as an extension of SMARTS to represent chemical reactions [Mobley et al., 2018] – used in the Open Reaction Database. InChi V1, not part of Daylight, was introduced and can represent basic prototautemerism [Heller et al., 2013], yet is unfortunately only probabilistically reliable. SMIRKS allows for string representation of reactions. It does not represent reaction features like temperature or pressure or catalysts. One feature it lacks is the ability to indicate the directionality of the reaction, which took a new language to be implemented ten years after the release of SMIRKS, called RinChi David et al. [2020], Wigh et al. [2022]. This was further extended years later, with another language called ProcAuxInfo, which enables the storage of metadata relating to yields, temperature, concentration, and other reaction conditions. Although, this represents definite progress, the profusion of new syntaxes and grammars, and their creative backronyms, create avoidable complexity for the cheminformatician. [Apodaca, 2022] introduces Balsa, a fully-specified language designed as a subset of SMILES. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
2.2.6 Extensions of SMILES
Despite the introduction of DeepSMILES, which fixed the bracketing issue in SMILES strings, the syntactic invalidity problem was not solved completely [O’Boyle and Dalke, 2018].
2.3 Other proposed representations
2.3.1 Fingerprints
Fingerprints, which typically involve encoding key features of a molecule into a vector representation, offer convenience for machine learning applications. However, a significant drawback of fingerprints is their inability to consistently represent tautomers, which are molecules that have the same atomic composition but different structural arrangements. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [Martin, 2009]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
Moreover, while fingerprints can capture the presence of specific substructures within molecules, they are susceptible to bit collisions and vector sparsity, which limit their overall representational power [Fang et al., 2022]. Bit collisions occur when multiple distinct substructures are mapped to the same bit in the fingerprint, leading to ambiguity and potential false positives. Vector sparsity arises when the fingerprint representation is high-dimensional, but only a small fraction of the bits are actively used, resulting in inefficient memory usage and computational overhead.
A fundamental issue with numeric vector representations, such as fingerprints, is the lack of inherent semantics associated with the individual numbers. Without additional contextual information or a predefined mapping, the numeric values in the vector do not carry intrinsic meaning. This lack of semantics can hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
In contrast, algebraic data types (ADTs) provide a structured and type-safe representation that inherently encodes the semantics of the molecular data. By defining the data structure and its associated types, ADTs constrain the space of possible values and operations, ensuring that only meaningful and valid transformations can be applied. This type-driven approach helps prevent errors and inconsistencies that may arise from arbitrary manipulation of numeric vectors.
Furthermore, the type information associated with ADTs can guide and constrain the machine learning process. The type constraints serve as a form of prior knowledge, restricting the search space and focusing the learning algorithm on meaningful and plausible solutions. This can lead to more efficient and effective learning, as the model can leverage the structure and semantics embedded in the ADT representation.
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [Capecchi et al., 2020], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint. These algebraic properties such as group theoretic properties Variational Auto Encoders (VAEs) also have problems, [Gómez-Bombarelli et al., 2018] claims obtaining structured output to be a hard and currently unsolved problem, even for a simple graph data structure.
2.4 Molecules and programming languages
Inductive Logic Programming (ILP), has seen usage in representing basic molecules Srinivasan et al. [1996], but the focus was doing rule induction over graphical data structures, with molecules being an example, rather than presenting ILP as a representational format for cheminformaticians. It is perhaps for this reason that it has had little uptake within the Cheminformatics community.
Molecular representations using Inductive Logic Programming (ILP) have been explored in previous research, such as the work by Srinivasan et al. [Srinivasan et al., 1996]. However, the application of machine learning techniques in conjunction with ILP-based molecular representations has not been extensively discussed in this early work. This lack of emphasis on machine learning aspects may explain the absence of the paper from recent reviews of molecular representations, such as the comprehensive survey by Wigh et al. [Wigh et al., 2022].
In the approach presented by Srinivasan et al. [Srinivasan et al., 1996], molecules are represented as logical rules using an extension of the Prolog programming language. Prolog is a declarative language that enables the expression of logical relationships between variables using first-order logic. Programs written in Prolog consist of a complex network of facts and rules, which are constructed using logical primitives such as conjunction, disjunction, and universal quantification. However, the data types supported by Prolog are limited to lists, integers, floats, and strings, with minimal support for user-defined data types.
Prolog’s inference mechanism allows for querying the knowledge base by providing examples and evaluating their truth values based on the presence or absence of counterexamples (negation as failure). Nevertheless, it is important to note that Prolog and its extensions are not purely declarative, as they incorporate side effects through built-in predicates such as read, get, and assert. These side effects can succeed or fail independently of the problem being solved, introducing challenges in reasoning about the program’s behavior, particularly when dealing with complex rule bases.
The specific extension of Prolog used in [Srinivasan et al., 1996] is called Progol, which employs inductive logic programming techniques to derive rules that describe the given data. It is worth clarifying that Progol should not be confused with ProbLog, a probabilistic logic programming language that focuses on modeling the probabilities of examples being true based on the probabilities of the underlying relations.
While the work by Srinivasan et al. [Srinivasan et al., 1996] demonstrates the feasibility of representing molecules using logical rules in an ILP framework, the integration of machine learning techniques with these representations was not explored.
This makes the proposal here the first to suggest to cheminformaticians to use programming languages more directly in representation. The notion that “molecular programming languages” could be possible, occurred in an article about the future of SELFIES [Krenn et al., 2022]. It is mentioned that molecular graphs could be executed, but the language is ambiguous and short. It is unclear whether they suggest using molecules to represent programming languages via the Church-Turing thesis, whether they want to compute with graph structures or are attempting to compute in a substrate other than silicon  [^3] https://molecularprogrammers.org/.
Investigations of type safe functional programming for cheminformatics has occurred before. Recent work has suggested the use of OCaml in writing software for cheminformatics [Berenger et al., 2019], [Höck and Riedl, 2012]. However, the emphasis in both of these articles as well as the Haskell library Ouch [T, 2021], was not on extensible molecular representations but on functional programming. This is highlighted by the papers introduction of lengthy tutorials on functional programming. Furthermore they do not show how Bayesian Inference can be performed on drug structures over algebraic data types which is essential in motivating efficiency improvements in not requiring exploration of syntactically nonsensical regions of molecular space.
2.4.1 RDKit as representation
One might also ask whether Cheminformatics libraries support representation. Behemoth Cheminformatics programming libraries, such as RDKit, [Landrum et al., 2013] and lightweight molecular manipulation libraries that exist on top of these e.g. datamol [^4] https://datamol.io/ and https://www.rdkit.org/ (which builds on NumPY [Oliphant et al., 2006], Pandas, RDKit and Python 3+), Datamol explicitly [^5] https://doc.datamol.io/stable/tutorials/Preprocessing.html don’t advocate themselves as providing a prototypical representation. These libraries parse representations to computational representations and then apply functions to them, before outputting them to strings like SMILES.
The computational representation of RDKit in C++ is complex, taking up over 400 lines of highly complex code and the molecule class taking up 873 lines of code. There are no clear invariants, or guarantees about the management of side effects in snippets shown below. The ubiquitous prescence of null pointers and highly abstruse C++ keywords, scattered throughout the code, insinuates points of failure of the program. It is unclear, whether and when the program may crash and what the consequences might be. The following snapshot of lines $792$ to $821$ in ROMol.h [^6] https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html which represents the molecule class, provides a brief glimpse into the extreme complexity of code:
Taken from https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html
--Line 792
    #ifdef RDK_USE_BOOST_SERIALIZATION
[…]
   template <class Archive>
   void save(Archive &ar, const unsigned int version) const;
   template <class Archive>
   void load(Archive &ar, const unsigned int version);
   BOOST_SERIALIZATION_SPLIT_MEMBER()
 #endif
  private:
   MolGraph d_graph;
   ATOM_BOOKMARK_MAP d_atomBookmarks;
   BOND_BOOKMARK_MAP d_bondBookmarks;
   RingInfo *dp_ringInfo = nullptr;
   CONF_SPTR_LIST d_confs;
   std::vector<SubstanceGroup> d_sgroups;
   std::vector<StereoGroup> d_stereo_groups;
   std::unique_ptr<boost::dynamic_bitset<>> dp_delAtoms = nullptr;
   std::unique_ptr<boost::dynamic_bitset<>> dp_delBonds = nullptr;
   friend RDKIT_GRAPHMOL_EXPORT std::vector<SubstanceGroup> &getSubstanceGroups(
       ROMol &);
   friend RDKIT_GRAPHMOL_EXPORT const std::vector<SubstanceGroup>
       &getSubstanceGroups(const ROMol &);
   void clearSubstanceGroups() { d_sgroups.clear(); }
  protected:
   unsigned int numBonds{0};
-- Line 821
Figure 4: This is just a small fragment of the molecule class in RDKit. It demonstrates the sheer complexity and obfuscation of a concept that can be represented more simply in the ADT proposed here.
The proposed ADT-based representation of molecules in Haskell offers a more concise, expressive, and type-safe approach. By leveraging the power of algebraic data types and the strong static typing of Haskell, we can define molecular structures in a declarative and compositional manner. The resulting code is more readable, maintainable, and less prone to runtime errors. The use of pure functions and the absence of side effects in the ADT representation enable easier reasoning about the behavior of the program and facilitate formal verification techniques.
Moreover, the ADT-based representation provides a clear separation of concerns between the data structure and the operations performed on it. This modular design allows for greater flexibility and extensibility, as new operations can be defined independently of the underlying data representation. The use of higher-order functions and type classes in Haskell further enhances the reusability and composability of the code.
While cheminformatics libraries like RDKit and datamol have been widely adopted and have proven their utility in various applications, their complex OOP-based representations can hinder the development and maintenance of robust and reliable software. The proposed ADT-based approach offers a compelling alternative that prioritizes simplicity, type safety, and mathematical rigor.
2.5 Probabilistic Programming
Languages that are said to be probabilistic are defined as those that support both random sampling from distributions and the conditioning of values of a variable via observations [Gordon et al., 2014]. At the core of probabilistic programming are two essential operations: sampling and scoring. Sampling allows for the generation of random values from probability distributions, while scoring enables the assignment of likelihoods to observed data given a probabilistic model. These operations form the foundation for performing Bayesian inference, where the goal is to infer the posterior distribution of model parameters given observed data. Furthermore, if programs featuring the ADT are written in do notation, as the probability probability / measure monad is commutative [Staton, 2017], consecutive sampling statements can be done independently and in parallel.
In the proposed approach, the Molecule algebraic data type (ADT) serves as the basis for constructing a generative model of molecules. By representing molecules as structured data using the ADT, we can define probability distributions over various molecular properties, such as atom types, bond types, and spatial coordinates.
We can define probability distributions over other molecular properties, such as bond types and electron configurations, using appropriate probability distributions.
Given a dataset of observed molecules, we can perform Bayesian inference to estimate the posterior distribution of the model parameters. One popular inference algorithm is Trace Metropolis-Hastings (Trace-MH), which is a Markov Chain Monte Carlo (MCMC) method adapted for probabilistic programming.
Trace-MH operates by proposing modifications to the execution trace of the probabilistic program, which represents a particular instantiation of the model parameters and random variables. The proposed trace is then accepted or rejected based on the Metropolis-Hastings acceptance criterion, which compares the likelihood of the proposed trace to the current trace.
Furthermore, the use of probabilistic programming allows for the incorporation of prior knowledge and the quantification of uncertainty in molecular modeling. By specifying informative priors over molecular properties and parameters, domain expertise can be incorporated into the modeling process. Additionally, the posterior distributions obtained through Bayesian inference provide a measure of uncertainty associated with the inferred molecular structures and properties, enabling more robust and reliable predictions.
The proposed approach also facilitates the integration of probabilistic programming with other machine learning techniques, such as deep learning. By leveraging the automatic differentiation capabilities of modern deep learning frameworks, it becomes possible to train complex neural network models within the probabilistic programming framework. This enables the development of hybrid models that combine the expressiveness of deep learning with the principled handling of uncertainty provided by probabilistic programming.
The proposed approach of using algebraic data types (ADTs) to represent molecules offers a significant advantage in terms of efficiency and meaningful exploration of chemical space. By directly synthesizing molecules in accordance with the grammar defined by the ADTs, the machine learning process can operate directly on the program or grammar itself, without the need for distilling the representation into another format.
This direct operation on the grammatical representation has important implications for the efficiency and effectiveness of the machine learning process. One of the key benefits is that it eliminates the need to explore nonsensical or invalid regions of chemical space. The grammar inherently constrains the space of possible molecules to those that are semantically meaningful and consistent with the rules and structure defined by the ADTs.
In traditional approaches, where molecules are represented using string-based formats like SMILES or graph-based representations, the machine learning algorithms often generate or explore a large number of candidate molecules, many of which may be invalid or chemically infeasible. This exploration of nonsensical regions of chemical space is computationally wasteful and can hinder the efficiency and effectiveness of the learning process.
By operating directly on the grammatical representation, the machine learning algorithm can focus its exploration on the space of valid and meaningful molecules. The ADTs provide a structured and constrained representation that ensures the generated molecules adhere to the specified rules and constraints. This targeted exploration of the chemical space leads to more efficient and meaningful learning, as the algorithm can avoid wasting computational resources on exploring invalid or nonsensical molecules.
Furthermore, the direct operation on the grammar allows for the seamless integration of domain knowledge and chemical constraints into the learning process. The ADTs can be designed to incorporate specific rules, symmetries, and constraints that are relevant to the problem at hand. By encoding this knowledge directly into the representation, the machine learning algorithm can leverage this information to guide its search and generate molecules that are more likely to exhibit desired properties or meet specific criteria.
The probabilistic conditioning operations, such as sampling and scoring, can be directly applied to the grammatical representation, enabling the algorithm to assign probabilities and likelihoods to different molecular structures and properties. This probabilistic treatment allows for the quantification of uncertainty and the incorporation of prior knowledge into the learning process.
Moreover, the use of a grammatical representation facilitates the interpretability and explainability of the generated molecules. Since the molecules are synthesized directly from the grammar, they inherently possess a structured and meaningful representation. This enables chemists and domain experts to easily interpret and understand the generated molecules, as they align with the defined rules and constraints of the ADTs.
In contrast, when molecules are distilled into other representations, such as fixed-length vectors or arbitrary graph structures, the interpretability and chemical meaning of the generated molecules can be lost or obscured. The direct operation on the grammar preserves the semantic structure and allows for a more intuitive understanding of the generated molecules.
One of the key benefits of using probabilistic programming for retrosynthetic analysis is the ability to directly generate and manipulate molecular graphs or trees. In traditional approaches, molecules are often represented using string-based formats like SMILES or graph-based representations, which require additional preprocessing and conversion steps. However, with probabilistic programming languages like LazyPPL, molecules can be directly synthesized as algebraic data types (ADTs) that adhere to a well-defined grammar. This allows for the direct generation and manipulation of molecular structures without the need for intermediate representations.
Previous implementations of Gaussian Processes have been restricted by the grammar [Moss and Griffiths, 2020]. By using Gaussian processes in combination with probabilistic programming, we can model the probability distributions over molecular properties and reactivity. Gaussian processes provide a flexible and expressive framework for capturing the complex relationships between molecular features and their corresponding properties. They allow for the incorporation of prior knowledge and the quantification of uncertainty in the predictions.
In the context of retrosynthetic analysis, Gaussian process regression can be used to predict the likelihood of a particular synthetic route or the feasibility of a retrosynthetic step. By training the Gaussian process model on a dataset of known reactions and their outcomes, it can learn the underlying patterns and relationships between molecular structures and reactivity. This learned model can then be used to guide the exploration of the chemical space and prioritize promising retrosynthetic pathways.
The integration of probabilistic programming with retrosynthetic analysis enables a more efficient and targeted exploration of the chemical space. By directly growing molecular trees or graphs and subjecting them to probabilistic conditioning operations, we can focus on the most relevant and plausible regions of the chemical space. The scoring and sampling mechanisms provided by probabilistic programming languages allow for the assignment of probabilities to different retrosynthetic routes and the selection of the most promising candidates.
2.6 Introduction to the proposal
The proposed representation here, is embedded in a pure, type-safe, functional programming language called Haskell. This purity makes reasoning about, and debugging code easier. Types, together with type inference allows one to have correctness guarantees at compile-time, eliminating runtime errors. None of the concepts presented here are conceptually specific to Haskell and are found in other functional programming languages.
The proposed representation is also motivated by established and the conceptually well motivated interest of Bayesian machine learning for drug design [Moss and Griffiths, 2020], [Semenova et al., 2020] & [Mervin et al., 2021]. Integration of probabilistic programming with Haskell e.g. [Paquet and Staton, 2021], [Ścibior et al., 2015], [Nguyen et al., 2022] has been done before laying the groundwork for usage with chemical data structures.
Algebraic Data Types (ADTs) provide a natural and intuitive way to represent molecular structures because they implicitly define a grammar for the data. ADTs are composed of simple, atomic data types combined using constructs such as products (tuples or records) and sums (variants or unions). This compositional nature allows for the creation of complex, hierarchical data structures that can accurately capture the intricate details of molecular structures. By defining molecules using ADTs, we inherently specify the allowed components and their valid combinations, effectively creating a grammar for the representation. This grammar ensures that only semantically meaningful structures can be constructed, preventing the formation of invalid or nonsensical molecules. Furthermore algebraic data types can be made instances of groups, allowing for a uniform, verifiable language to describe rotational symmetries of molecules.
One perk of this representation is the ability to be more general than a Lewis structure as it allows “individual electrons” to be in multiple covalent bonds. This is achieved with a representation published in 1995 [Dietz, 1995]. This allows for modelling delocalised electrons thus allowing for resonance structure, organometallics and multi-center or electron deficient bonds. This is advocated for in recent papers such as those discussing the future of molecular representations[Krenn et al., 2022]. Modelling organometallics has already proved useful in the drug design process, with Platinum based drugs being used to prevent cancer [dilruba2016platinum]. One result of this is the ability to calculate fractional and zero bond orders, often impossible in other common representations. Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbitals with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
The ADT introduced is fairly extensible: removing the need for completely different syntaxes to be created for each different kind of molecule. Haskell’s record syntax further enables access of only parts of the molecule needed for a task e.g. in machine learning. This means redundant data is discarded, making machine learning models less wasteful, thus more efficient and interpretable. This is difficult to do with some representations, where dissecting or parsing the string is difficult, so the whole molecule is fed in, including redundant data.
The representation is easily extensbile allowing chemists to adjust their data structure straightforwardly in line with their needs.
[…]
Probabilistic programming languages (PPLs) have emerged as a powerful paradigm for performing Bayesian inference, enabling the representation and manipulation of probability distributions over complex data structures. PPLs provide a unified framework for specifying probabilistic models, performing inference, and making predictions based on observed data. The program calculates the posterior provided a model and an inference strategy, where inference is broadly defined by Gordon et al. [Gordon et al., 2014] as the problem of computing an explicit probability distribution implicit in the program. The idea that a program can represent a probabilistic model goes back to 1978 [Saheb-Djahromi, 1978] and programming languages for Bayesian modelling have existed since at least 1994 [Gilks et al., 1994]. The goal of probabilistic programming as mentioned by [Gordon et al., 2014] is to hide inference inside the compiler to allow the programmer to focus on writing models and have inference automatically done for them.
In the context of molecular representation, probabilistic programming offers a natural way to capture the inherent uncertainty and variability associated with molecular structures and properties. By combining the expressive power of functional programming with the ability to represent and reason about probability distributions, PPLs enable the development of rich, probabilistic models of molecules.
Further efficiency guarantees are provided by using a lazy language. Haskell as a lazy language, means that it only evaluates functions when needed. This can provide efficiency improvements when simulations are performed. This is very helpful when sampling from molecular orbitals which are represented naturally in this library by higher order functions over random samples. In particular, combining it with LazyPPL [Paquet and Staton, 2021], a recently developed lazy probabilistic programming language that only samples when necessary may provide efficiency improvements.
3 Algebraic Data Types
An algebraic data type (ADT) in Haskell represents data with additional structure and they differentiate functions from data. An ADT provides a structured way to organize data by utilizing type constructors, such as sums and products, to create complex types from simpler ones. This structured organization of data adheres to algebraic laws. Typeclasses in Haskell exemplify this capability by denoting which data types support specific operations and specification of the equational laws that these operations must satisfy can be stated. By making a data type an instance of a typeclass, we declare that the data type supports the operations defined by the typeclass and claim that these operations adhere to the specified equational laws. This allows for the representation of complex algebraic structures, such as algebraic groups within the type system. Moreover, typeclasses enable a high level of abstraction and code reuse, promoting modularity and extensibility in the context of molecular representation, where they can be used to define the algebraic structures and operations relevant to molecules, such as groups for rotational symmetries.
For example, the Maybe type is an ADT that can encapsulate a computation that may or may not yield a value, signified by its constructors Nothing or Just a, respectively. This provides a type-safe way to handle operations that could fail, or denote an abscence of a value rather than using zero – a concept that in other languages might be handled by null references or exceptions, which may introduce errors.
As mentioned, in Haskell, algebraic data types (ADTs) are constructed using sum and product type constructors, analogous to addition and multiplication in algebra. A sum type is a choice between different types (denoted by the | symbol), akin to an or; for example, a Maybe type represents a value that may be Nothing (like zero in addition) or Just a where a is any type. This is reflective of a disjoint sum in algebra, where a value can be one thing or another, but not both simultaneously. This allows the grammar to distinguish between empty values and zero values. As will be discussed, the orientation field of the Orbital data type is wrapped in a Maybe type, which allows for the natural expression of whether an orbital has a specific orientation or not. For $s$ orbitals, the orientation is Nothing because $s$ orbitals are spherically symmetric and have no spatial orientation. In contrast, $p$ orbitals have specific orientations, and their orientation can be represented as Just Coordinate, where Coordinate is a type that can represent orientation in the $x,y,z$ axes. On the other hand, a product type combines several values into one composite type, similar to how factors multiply to a product; a tuple (a, b) is an instance of a product type containing both a and b.
In contrast to Haskell’s strong, static type system, languages like Python employ dynamic typing, which can lead to a different class of runtime errors. Large-scale machine learning applications often written in Python, such as those using the PyTorch framework, must contend with the possibility of type-related errors during execution. Whilst this is a well known advantage of using Haskell, it is not the focus of the paper.
The ADT in this article provides a minimal representation for cheminformaticians to use and extend. One desirable characteristic of this representation, is that it allows electrons to participate in more than one covalent or hydrogen bond, which is a consequence of using the representation set up by Dietz, which this article will refer to as the Dietz Representation [Dietz, 1995]. Ferrocene, an important organometallic that is difficult to faithfully represent in current formats now has a straightforward and readable representation. The Dietz format also allows for resonance structures (with multiple contributing graphs) to be exhibited [Dietz, 1995]. Indeed, implementing precisely this was proposed as an advancement of future molecular representations by [Krenn et al., 2022]. Furthermore, with the 3D coordinates of each atom, the ADT can be used to calculate the dihedral angles, bond angles and bond lengths, which are also unavailable in SMILES strings. The addition of these adds minimally to the space complexity, yet can be crucial for predicting the physical properties of drugs. Whilst not done here, the representation could be extended to implement single-atom tautomerism (which for example is more general than prototautemorism, representable with hydrogen in the inChi V1 framework) [Pospisil et al., 2003].
As a consequence of being an ADT that derives Show and Read instances, molecules can be leveraged as a transmission representation by directly serialising them to Haskell compatible syntax. These can act as storage files which can be read from, when a program requires said molecule.
Record syntax in Haskell offers a robust mechanism for defining algebraic data types. They simplify the process of working with data by automatically generating getter and setter functions, thus reducing boilerplate code. This feature not only enhances code readability but also supports the seamless extension of data types without impacting existing functionalities.
3.1 Record Syntax
Record syntax allows for defining named fields within data types, facilitating direct access to individual parts of a data structure. The use of these names acts as a way to prevent confusion about whether one has a specific atom or coordinate in a data structure representing atoms, for instance. Here is a simple example demonstrating record syntax:
data Atom = AtomS { atomID :: Integer,
                    symbol :: AtomicSymbol }
          | AtomC { atomID :: Integer,
                    symbol :: AtomicSymbol,
                    coordinate :: Coordinate }
data AtomicSymbol = C | O | H
data Coordinate = Coordinate { x :: Double,
                               y :: Double,
                               z :: Double }
In this example, Atom is a sum type with two constructors: AtomS and AtomC. The AtomS constructor represents an atom with a symbol and has two fields: atomID of type Integer and symbol of type AtomicSymbol. The AtomC constructor represents an atom with a symbol and a coordinate, and it has three fields: atomID of type Integer, symbol of type AtomicSymbol, and coordinate of type Coordinate. Record syntax automatically generates getter functions for each field.
Example 1: Updating the coordinate of an atom
setCoordinate :: Coordinate -> Atom -> Atom
setCoordinate newCoordinate atomWC@(AtomC _ _ _) = atomWC { coordinate = newCoordinate }
setCoordinate _ atom = atom
The setCoordinate function updates the coordinate of an AtomC. It takes a new Coordinate and an Atom as arguments. If the atom is an AtomC, it uses the @ pattern to bind the entire AtomC value to the variable atomWC. This allows us to refer to the original AtomC value when updating its fields. We then use record syntax to create a new AtomC by updating only the coordinate field of atomWC with the newCoordinate, while keeping the atomID and symbol fields unchanged. If the atom is an AtomS, it returns the original atom unchanged since AtomS does not have a coordinate field.
Example 2: Retrieving the x-coordinate of an atom
getXCoordinate :: Atom -> Maybe Double
getXCoordinate (AtomC _ _ coordinate) = Just (x coordinate)
getXCoordinate _ = Nothing
The getXCoordinate function retrieves the x-coordinate of an AtomC. It uses pattern matching to extract the coordinate field from the AtomC constructor. If the atom is an AtomC, it uses record syntax to access the x field of the coordinate and returns it wrapped in Just. If the atom is an AtomS, it returns Nothing since AtomS does not have a coordinate field.
Record syntax in Haskell provides a basic level of extensibility for data types, allowing the addition of new fields without requiring changes to existing functions that operate on the data type. This capability enables the evolution of data models with minimal impact on the codebase. This extensibility is particularly valuable in the context of molecular representation, as it allows for the incorporation of new scientific knowledge and the adaptation of the representation to specific domains or applications. However, for more sophisticated and flexible data manipulation, lenses offer a powerful approach. Lenses are first-class getter & setter functions that can be composed to access and modify deeply nested fields or manipulate multiple fields simultaneously, simplifying the process of working with complex or nested data structures by focusing directly on the target fields, bypassing the need for explicit navigation.
3.2 The Molecule ADT
The following presents the heart of this paper, the molecule data type:
{verbnobox}
[] data Molecule = Molecule atoms :: [Atom], bonds :: M.Map (Integer, Integer) BondType deriving (Show, Read)
data Atom = Atom atomID :: Integer, atomicAttr :: ElementAttributes, coordinate :: Coordinate, shells :: Shells deriving (Show, Read)
data BondType = HydrogenBond | CovalentBond delocNum :: Integer, atomIDs :: Maybe [Integer] | IonicBond deriving (Eq, Read, Show)
data AtomicSymbol = O | H | N | C | B | Fe | F | Cl | S | Br | P | I deriving (Eq, Read, Show)
data ElementAttributes = ElementAttributes symbol :: AtomicSymbol, atomicNumber :: Integer, atomicWeight :: Double deriving (Eq, Read, Show)
data Coordinate = Coordinate x :: Double , y :: Double , z :: Double deriving (Show, Eq, Read)
Figure 5: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent HydrogenBond, CovalentBond (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over), or IonicBond.
The Molecule data type consists of two fields: atoms and bonds. The atoms field is a list of Atom values, representing the individual atoms that make up the molecule. Each Atom is identified by a unique atomID of type Integer, ensuring that each atom can be distinctly referenced within the molecule. The Integer type in Haskell represents arbitrary-precision integers, allowing for the representation and manipulation of integers without any predefined size limits, thus avoiding the issue of overflow that can occur with fixed-size types like Int. Unlike Int, which is limited to a fixed range determined by the computer’s architecture, Integer can handle values as large as the available memory permits.
The Atom data type contains several important attributes. The atomicAttr field of type ElementAttributes captures the fundamental properties of an atom, such as its atomic symbol (AtomicSymbol), atomic number (atomicNumber), and atomic weight (atomicWeight). These attributes provide essential information about the nature and characteristics of each atom in the molecule.
In addition to the atomic attributes, each Atom also has a coordinate field of type Coordinate, which represents the spatial position of the atom within the molecule. Each Coordinate has an x, y and z field. This information is crucial for understanding the geometric structure and spatial arrangement of the atoms, enabling the analysis of molecular shape, conformation, and reactivity.
Furthermore, the shells field of type Shells represents the electronic configuration of the atom, capturing the arrangement of electrons in the various atomic orbitals. This information is useful for understanding the chemical behavior and bonding properties of the atom, as well as its participation in chemical reactions. This will be discussed in the section on the Orbital type.
The BondType data type is a sum type that captures the different types of chemical bonds that can exist between atoms. It includes HydrogenBond, CovalentBond, and IonicBond as possible variants. The CovalentBond constructor takes additional fields, delocNum and atomIDs, which represent the delocalization number and the optional ring structure associated with the covalent bond. This allows for the representation of delocalized electrons and cyclic structures within the molecule. This allows for the accurate modeling of the dynamic distribution of electrons across multiple bonds within a molecule, capturing the essence of the Dietz representation. By incorporating this information, the data type can effectively represent resonant structures and account for the delocalization of electrons, which is crucial for understanding the electronic properties and reactivity of molecules. The atomIDs field, of type Maybe [Integer], provides a means to represent the interconnected bonds within a cyclic structure. It is an optional field, indicated by the Maybe type constructor, allowing for the presence or absence of a list of atomIDs across which electrons are delocalised associated with the covalent bond. This list of integers enables for instance the representation of cyclic structures within the molecule. This design choice of using the delocNum and atomIDs fields in the CovalentBond constructor offers a flexible framework capable of accurately depicting complex molecules, such as organometallic compounds like ferrocene. It provides a foundation for computing fractional bond orders, which is essential for molecules with resonant bonding characteristics, such as ozone ($O_{3}$).
3.2.1 Example of a Molecule : Benzene
{verbnobox}
[] benzene :: Molecule benzene = Molecule atoms = [ Atom atomID = 1, atomicAttr = elementAttributes C, coordinate = Coordinate 0.0 1.3948 0.0, shells = carbon , Atom atomID = 2, atomicAttr = elementAttributes C, coordinate = Coordinate 1.2079 0.6974 0.0, shells = carbon , Atom atomID = 3, atomicAttr = elementAttributes C, coordinate = Coordinate 1.2079 (-0.6974) 0.0, shells = carbon , Atom atomID = 4, atomicAttr = elementAttributes C, coordinate = Coordinate 0.0 (-1.3948) 0.0, shells = carbon , Atom atomID = 5, atomicAttr = elementAttributes C, coordinate = Coordinate (-1.2079) (-0.6974) 0.0, shells = carbon , Atom atomID = 6, atomicAttr = elementAttributes C, coordinate = Coordinate (-1.2079) 0.6974 0.0, shells = carbon , Atom atomID = 7, atomicAttr = elementAttributes H, coordinate = Coordinate 0.0 2.4732 0.0, shells = hydrogen , Atom atomID = 8, atomicAttr = elementAttributes H, coordinate = Coordinate 2.1431 1.2366 0.0, shells = hydrogen , Atom atomID = 9, atomicAttr = elementAttributes H, coordinate = Coordinate 2.1431 (-1.2366) 0.0, shells = hydrogen , Atom atomID = 10, atomicAttr = elementAttributes H, coordinate = Coordinate 0.0 (-2.4732) 0.0, shells = hydrogen , Atom atomID = 11, atomicAttr = elementAttributes H, coordinate = Coordinate (-2.1431) (-1.2366) 0.0, shells = hydrogen , Atom atomID = 12, atomicAttr = elementAttributes H, coordinate = Coordinate (-2.1431) 1.2366 0.0, shells = hydrogen ]
, bonds = M.fromList \$ getSymmetricBonds [ ((1,2), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((1,6), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((1,7), CovalentBond delocNum = 2, atomIDs = Nothing) , ((2,3), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((2,8), CovalentBond delocNum = 2, atomIDs = Nothing) , ((3,4), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((3,9), CovalentBond delocNum = 2, atomIDs = Nothing) , ((4,5), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((4,10), CovalentBond delocNum = 2, atomIDs = Nothing) , ((5,6), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((5,11), CovalentBond delocNum = 2, atomIDs = Nothing) , ((6,12), CovalentBond delocNum = 2, atomIDs = Nothing) ]
Figure 6: Representation of the benzene molecule using the proposed algebraic data type (ADT) system. The atoms field contains a list of 12 Atom values, representing the six carbon atoms and six hydrogen atoms. Each atom is specified with its unique atomID, atomicAttr (including the AtomicSymbol, atomicNumber, and atomicWeight), coordinate (3D spatial position), and shells (electronic configuration). The bonds field is represented using a Map from tuples of atom IDs to BondType values, capturing the delocalized nature of the electrons in the aromatic ring and the explicit representation of the ring structure.
The provided code demonstrates the representation of the benzene molecule using the proposed algebraic data type (ADT) system. Benzene is a classic example of an aromatic compound, consisting of a ring of six carbon atoms with alternating single and double bonds (or delocalised electrons around a ring), and six hydrogen atoms attached to each carbon.
In this representation, the benzene molecule is defined as a value of type Molecule. The atoms field contains a list of 12 Atom values, representing the six carbon atoms and six hydrogen atoms in the molecule. Each atom is specified with its unique atomID, atomicAttr (including the AtomicSymbol, atomicNumber, and atomicWeight), coordinate (3D spatial position), and shells (electronic configuration).
The carbon atoms (atomIDs 1-6) are positioned in a hexagonal arrangement, with coordinates that define the vertices of the hexagon. The hydrogen atoms (atomIDs 7-12) are positioned radially outward from each carbon atom, with coordinates that place them at a typical carbon-hydrogen bond distance.
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as CovalentBond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6], explicitly specifying the cyclic structure formed by the carbon atoms.
The bonds between the carbon atoms and their respective hydrogen atoms are also represented as CovalentBond values, but with a delocNum of 2, indicating a typical single bond between carbon and hydrogen. The atomIDs field for these bonds is therefore set to Nothing, as they are not part of the aromatic ring.
The code also includes the definitions of elementAttributes for hydrogen (H), nitrogen (N), and carbon (C). These elementAttributes specify the AtomicSymbol, atomicNumber, and atomicWeight for each element. For example, elementAttributes C = ElementAttributes C 6 12.011 represents the carbon element with an atomic number of $6$ and an atomic weight of $12.011$. These predefined elementAttributes are used in the atomicAttr field of each Atom in the benzene molecule, providing the necessary information about the elemental properties of the atoms. The representation could be extended to represent reactions or polymers, without requiring a completely new and separate syntax like BigSMILES. Instead, lists whose syntax and semantics represents reflects repetition of subunits.
3.3 The Orbital ADT
{verbnobox}
[] – Orbital types for each subshell data So = So deriving (Show, Eq, Read) data P = Px | Py | Pz deriving (Show, Eq, Read) data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read) data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
– The Orbital has a type and a number of electrons. data Orbital subshellType = Orbital orbitalType :: subshellType , electronCount :: Int , orientation :: Maybe Coordinate deriving (Show, Eq, Read)
– Each SubShell consists of a list of Orbitals. – We ensure that all orbitals within a subshell are of the same type – by using the subshellType parameter. newtype SubShell subshellType = SubShell orbitals :: [Orbital subshellType] deriving (Show, Eq, Read)
– Each Shell has a principal quantum number and consists of a list of SubShells. data Shell = Shell principalQuantumNumber :: Int , sSubShell :: Maybe (SubShell So) , pSubShell :: Maybe (SubShell P) , dSubShell :: Maybe (SubShell D) , fSubShell :: Maybe (SubShell F) deriving (Show, Eq, Read)
– An Atom consists of a list of Shells. type Shells = [Shell]
Figure 7: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells.
The representation captures quantum mechanical properties of atoms and can be made to adhere to the principles governing the arrangement of electrons in an atom. Each shell is represented as a distinct entity, characterized by its principal quantum number. This aligns with the quantum mechanical description of atomic structure, where electrons occupy discrete energy levels or shells. By organizing the electronic configuration into separate shells, the representation accurately reflects the hierarchical nature of atomic orbitals and the increasing energy and complexity of the electronic structure as the principal quantum number increases.
The use of a list to represent the shells implies that there could be a countably infinite number of shells, which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
Within each shell, there are subshells ($s,p,d$, and $f$), which correspond to the different types of atomic orbitals. Any interface to this data structure must respect the maximum number of electrons that can occupy each subshell, with $s$ subshells containing up to two electrons, $p$ subshells containing up to six electrons, $d$ subshells containing up to ten electrons, and $f$ subshells containing up to fourteen electrons. However, since the current implementation does not utilize dependent types, the correctness of the number of electrons in each subshell cannot be guaranteed at compile time and is left as future work. The responsibility of ensuring the proper electronic configuration lies with the user of the ADT.
Each atomic orbital, contains the subshell type, electron count, and spatial orientation of the orbital. These properties can be used to construct the actual basis functions, but they are not the basis functions themselves. Furthermore, the representation allows for the specification of the spatial orientation of each orbital within a subshell through the use of the Coordinate type. This information is crucial for describing the shape and symmetry of atomic orbitals, particularly for $p$, $d$, and $f$ orbitals that have distinct orientations. However, it is important to note that the actual spatial representation of the orbitals would require the use of the appropriate mathematical functions, such as Slater-type orbitals (STOs), spherical harmonics, or Gaussian functions. Slater-type orbitals are commonly used in quantum chemistry as they provide a more accurate representation of the electron density near the nucleus compared to Gaussian functions. However, these mathematical functions are not explicitly included in this ADT system, which focuses on the symbolic representation of the electronic configuration.
By designing the appropriate interface functions that operate on the Shells type, one can ensure that rules such as the Aufbau principle, Hund’s rule, and the Pauli exclusion principle are properly enforced when manipulating the electronic configuration. These interface functions would act as a layer of abstraction, encapsulating the logic and rules governing the distribution of electrons within the atom. By providing a well-defined set of operations that adhere to these principles, the interface can maintain the integrity and consistency of the electronic configuration, thereby accurately representing the charge of the molecule. The interface functions would handle the creation, modification, and querying of the Shells data structure, ensuring that any changes made to the electronic configuration are valid and compliant with the quantum mechanical principles. This approach allows for a clear separation of concerns, where the Shells type focuses on the data representation, while the interface functions handle the logic and enforcement of the principles governing the electronic configuration.
The Aufbau principle can be enforced by ensuring that electrons are added to orbitals in the correct order of increasing energy. Hund’s rule can be implemented by maximizing the number of unpaired electrons in degenerate orbitals before pairing them. The Pauli exclusion principle can be enforced by limiting the maximum number of electrons in each orbital to two, with opposite spins. The use of type information can greatly assist in enforcing these principles and ensuring the validity of the electronic configuration. For example, the SubShell data type could be parameterized by the maximum number of orbitals allowed for each subshell type. This type-level information would prevent the creation of invalid subshells with more orbitals than permitted by the quantum mechanical rules. Similarly, the electronCount field of the Orbital data type can be constrained to a range of values (e.g., 0 to 2) using type-level naturals or refined types. This ensures that the Pauli exclusion principle is upheld, and no orbital contains more than two electrons.
However, it does not explicitly account for the spin of electrons or the coupling between orbital and spin angular momenta, which play a significant role in determining the detailed electronic structure of atoms, particularly for elements with partially filled subshells.
3.3.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
{verbnobox}
[] hydrogen :: Shells hydrogen = [ Shell principalQuantumNumber = 1 , sSubShell = Just (SubShell [Orbital So 1 Nothing]) – 1s^1 , pSubShell = Nothing , dSubShell = Nothing , fSubShell = Nothing ]
carbon :: Shells carbon = [ Shell principalQuantumNumber = 1 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 1s^2 , pSubShell = Nothing , dSubShell = Nothing , fSubShell = Nothing , Shell principalQuantumNumber = 2 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 2s^2 , pSubShell = Just (SubShell [Orbital Px 1 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0))]) – 2p^2 , dSubShell = Nothing , fSubShell = Nothing ]
oxygen :: Shells oxygen = [ Shell principalQuantumNumber = 1 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 1s^2 , pSubShell = Nothing , dSubShell = Nothing , fSubShell = Nothing , Shell principalQuantumNumber = 2 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 2s^2 , pSubShell = Just (SubShell [Orbital Px 2 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0)), Orbital Pz 1 (Just (Coordinate 0 0 1))]) – 2p^4 , dSubShell = Nothing , fSubShell = Nothing ]
Figure 8: Examples of electronic configurations for hydrogen, carbon, and oxygen atoms using the algebraic data type representation in Haskell.
For the hydrogen atom, the electronic configuration is represented by a single shell with a principal quantum number of 1. The shell contains only an $s$ subshell (sSubShell) with one orbital occupied by a single electron, denoted as Orbital So 1 Nothing. The Nothing value for the orientation indicates that s orbitals have a spherical symmetry and do not have a specific orientation. This represents the ground state configuration of hydrogen, which is 1s1.
The carbon atom’s electronic configuration is represented by two shells. The first shell, with a principal quantum number of 1, consists of an $s$ subshell with two electrons (Orbital So 2 Nothing), corresponding to the 1s2 configuration. The second shell, with a principal quantum number of $2$, comprises an $s$ subshell with two electrons (Orbital So 2 Nothing) and a $p$ subshell (pSubShell) with two orbitals, each containing one electron. The p orbitals are specified as Orbital Px 1 (Just (Coordinate 1 0 0)) and Orbital Py 1 (Just (Coordinate 0 1 0)), representing the 2px and 2py orbitals, respectively. The Coordinate values (1, 0, 0) and (0, 1, 0) indicate the orientation of the p orbitals along the x and y axes, respectively. This configuration, 1s22s22p2, describes the ground state of carbon.
Similarly, the oxygen atom’s electronic configuration is represented by two shells. The first shell is identical to that of carbon, with an $s$ subshell containing two electrons. The second shell consists of an $s$ subshell with two electrons and a $p$ subshell with four electrons distributed across three orbitals: Orbital Px 2 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0)), and Orbital Pz 1 (Just (Coordinate 0 0 1)). The Coordinate values (1, 0, 0), (0, 1, 0), and (0, 0, 1) represent the orientation of the $p$ orbitals along the $x$, $y$, and $z$ axes, respectively. This describes the ground state configuration of oxygen, which is 1s22s22p4. The oxygen atom’s electronic configuration differs from carbon in the occupancy of the p subshell, where oxygen has four electrons distributed across three p orbitals, while carbon has two electrons in two p orbitals.
The inclusion of the Coordinate type in the representation allows for the specification of the spatial orientation of the orbitals. This is particularly relevant for $p$, $d$, and $f$ orbitals, which have distinct orientations in three-dimensional space. By providing the coordinates, the representation captures the directional nature of these orbitals, enabling a more accurate description of the electronic structure and the spatial distribution of electrons in the atom.
3.4 Reaction Networks, Group Theory and Serialisation
3.4.1 Reaction Networks
The proposed algebraic data type for representing molecules can be extended to model chemical reactions and reaction networks. By defining a Reaction data type that captures the reactants, products, and conditions under which a reaction occurs, we can create a powerful tool for studying and simulating chemical processes. A chemical reaction network comprises a set of reactants, a set of products, and a set of reactions.
Reactions are commutative but not associative and given further time could be incorporated into the type. This would be called an abelian or commutative magma.
Here’s an example of how the Reaction data type can be defined:
data Reaction = Reaction {
    reactants :: [(Double, Molecule)]
  , products :: [(Double, Molecule)]
  , conditions :: [Condition]
  , reactionRate :: Double
data Condition = TempCondition {temperature :: Double, times :: Times}
               | PressureCondition {pressure :: Double, times :: Times}
newtype Times = Times (StartTime Double, EndTime Double)
In this representation, a Reaction consists of a list of reactants and products, each represented as a tuple of a stoichiometric coefficient (of type Double) and an Atom. The conditions field is a list of Conditions, which can be either temperature or pressure conditions, specified along with the time range over which they apply.
Using this Reaction data type, we can define a ChemicalReactionNetwork as a collection of reactions:
newtype ChemicalReactionNetwork = ChemicalReactionNetwork {
    reactions :: [Reaction]
This allows for the modeling of complex chemical systems and the study of their behavior under different conditions. For example, consider the reaction for the formation of water from hydrogen and oxygen:
exampleReaction :: Reaction
exampleReaction = Reaction {
    reactants = [(2.0, h2), (1.0, o2)]
  , products = [(2.0, h2o)]
  , conditions =
    [ TempCondition 500.0 (Times (0.0, 10.0))
    , PressureCondition 1.0 (Times (0.0, 10.0))
    reactionRate = 0.1
This reaction represents the formation of two water molecules from two hydrogen molecules and one oxygen molecule, under the conditions of a temperature of $500K$ and a pressure of $1$ atm, over a time range of 00 to $10$ seconds.
3.4.2 Group Theory
Moreover, the ADT can be instantiated as a group to incorporate group theory concepts, which play a crucial role in understanding molecular symmetry and its impact on chemical properties. In particular, molecules can be made instances of rotation groups, such as the cyclic group $C3$, which represents the rotational symmetry of molecules with threefold rotational axes.
Here’s an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass:
class Group g where
    mul :: g -> g -> g     -- Group operation: combining two operations
    inv :: g -> g          -- Inverse of an operation
    e   :: g -> g          -- Identity operation
data MolRot = MolRot Molecule Coordinate Double deriving (Show)
instance Group MolRot where
    mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) =
        combineRotations mol1 axis1 angle1 axis2 angle2
    inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle)
    e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0
In this representation, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e).
By making MolRot an instance of Group, we can combine rotations using the mul function, find the inverse of a rotation using inv, and obtain the identity rotation using e. This allows for the study of molecular symmetry and its implications on chemical properties and reactivity.
However, it’s important to note that proving that a data type satisfies all the group axioms (closure, associativity, identity, and inverse) would require the use of dependent types, which is beyond the scope of this representation. Nonetheless, by defining the necessary operations and providing sensible implementations, we can still leverage the power of typeclasses in representing these properties.
3.4.3 Reading & Writing
{verbnobox}
[] – Writing the molecule to a file writeMoleculeToFile :: FilePath -> Molecule -> IO () writeMoleculeToFile filePath molecule = writeFile filePath (show molecule)
– Reading the molecule from a file readMoleculeFromFile :: FilePath -> IO Molecule readMoleculeFromFile filePath = do contents <- readFile filePath return (read contents)
– Example usage main :: IO () main = do – Write the methane molecule to a file writeMoleculeToFile "methane.hs" methane
– Read the methane molecule from the file molecule <- readMoleculeFromFile "methane.hs"
– Print the molecule read from the file print molecule
Figure 9: The provided Haskell code demonstrates how molecules can be easily read from and written to files.
The code defines two functions:
  1. 1.
    writeMoleculeToFile :: FilePath -> Molecule -> IO () – This function takes a file path and a Molecule as input and writes the molecule to the specified file. It uses the show function to convert the molecule to a string representation and then writes it to the file using the writeFile function.
  2. 2.
    readMoleculeFromFile :: FilePath -> IO Molecule – This function takes a file path as input and reads a Molecule from the specified file. It uses the readFile function to read the contents of the file as a string and then uses the read function to parse the string back into a Molecule.
The main function provides an example usage of these functions. It first writes the methane molecule to a file named methane.hs using the writeMoleculeToFile function. Then, it reads the molecule back from the same file using the readMoleculeFromFile function and prints the molecule to the console.
This code demonstrates the ease of serializing and deserializing molecules to and from files in Haskell. By leveraging the Show and Read instances defined for the Molecule data type and its related types, molecules can be seamlessly converted to and from their string representations, allowing them to be stored in files and retrieved later.
This functionality is particularly useful when working with large molecules or when you need to persist molecules across different program runs. It provides a convenient way to save and load molecular data, enabling further analysis, visualization, or processing of the molecules as needed.
3.4.4 A simple example of probablistic programming
Consider the following code:
{verbnobox}
[] – Observed molecule (e.g., H2O) observedMolecule :: Molecule observedMolecule = h2o
– A generative model for a molecule moleculeModel :: Meas Molecule moleculeModel = do let numAtoms = 3 atoms <- replicateM numAtoms \$ do symbol <- sample \$ uniformD [C, N, O, H] coord <- sampleCoordinate let attr = elementAttributes symbol let shellsVar = elementShells symbol return \$ Atom atomID = 0, atomicAttr = attr, coordinate = coord, shells = shellsVar let molecule = Molecule atoms = atoms, bonds = randomBonds 3
– Score the model based on the distance to the observed molecule let distance = distMolecule molecule observedMolecule score \$ normalPdf 0 1 distance
return molecule
– Sample a coordinate from a normal distribution sampleCoordinate :: Meas Coordinate sampleCoordinate = do x <- sample \$ normal 0 1 y <- sample \$ normal 0 1 z <- sample \$ normal 0 1 return \$ Coordinate x y z
– Calculate the Euclidean distance between two coordinates euclideanDistance :: Coordinate -> Coordinate -> Double euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) = sqrt \$ (x1 - x2)^2 + (y1 - y2)^2 + (z1 - z2)^2
– Calculate the distance between two molecules distMolecule :: Molecule -> Molecule -> Double distMolecule mol1 mol2 = let coords1 = map coordinate (atoms mol1) coords2 = map coordinate (atoms mol2) distances = [euclideanDistance c1 c2 | c1 <- coords1, c2 <- coords2] in minimum distances
– Example usage main :: IO () main = do samples <- mh 0.1 moleculeModel let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples putStrLn \$ "Sampled molecules: " ++ show molecules putStrLn \$ "Weights: " ++ show weights
Figure 10: Haskell code for generating and conditioning molecules using a probabilistic model. The model generates molecules and scores them based on their distance to an observed molecule (e.g., H2O). Metropolis-Hastings sampling is used to generate samples from the conditioned model.
Explanation:
  • We define observedMolecule as the molecule we want to condition our model on. In this example, we use the predefined h2o molecule, but you can replace it with any other molecule of interest.
  • We define a generative model moleculeModel using the Meas monad. This model generates a Molecule by first sampling the number of atoms numAtoms from a discrete uniform distribution using uniformD. In this case, we set numAtoms to a fixed value of 3.
  • We then use replicateM to generate numAtoms number of Atoms. For each Atom, we sample the symbol uniformly from the list [C, N, O, H] using uniformD, and sample the coordinate using the sampleCoordinate function.
  • We use record syntax to create each Atom. The atomID is set to a default value of 0 (which can be updated later), the atomicAttr is set using the elementAttributes function based on the sampled symbol, and the shells are set using the elementShells function, also based on the symbol. The charge is assumed to be neutral at this point.
  • After generating the list of atoms, we create a Molecule using record syntax. The atoms field is set to the generated list of atoms, and the bonds field is set using the randomBonds function, which randomly initializes a 3x3 adjacency matrix.
  • The sampleCoordinate function is a helper function that samples a Coordinate by sampling the $x$, $y$, and $z$ values from normal distributions using normal.
  • The randomBonds function generates a random 3x3 adjacency matrix representing the bonds between atoms. Currently it is not implemented for brevity, but the implementation would be straightforward.
  • In the moleculeModel, after generating the molecule, we calculate the distance between the generated molecule and the observed molecule using the distMolecule function.
  • We use the score function to compute the likelihood of the distance given a normal distribution with mean 00 and standard deviation $1$. This scoring step conditions the model on the observed molecule, giving higher weights to generated molecules that are closer to the observed molecule.
  • The distMolecule function takes two molecules as input and calculates the distance between them. It extracts the coordinates of the atoms in each molecule and computes the Euclidean distance between all pairs of coordinates using euclideanDistance. The minimum distance among all pairs is returned as the distance between the molecules.
  • In the main function, we generate samples using Metropolis-Hastings and drop the first one thousand samples as a burn in. We then print the remaining molecules and weights.
By having an observed molecule and using the distMolecule function, we can condition our generative model on any molecule of interest. The scoring mechanism computes the distance between the generated molecule and the observed molecule, guiding the sampling process towards molecules that are more similar to the observed one.
The Prob monad is used for representing probability distributions, while the Meas monad is used for representing unnormalized probability distributions (measures). The sample function is used to sample from a probability distribution (Prob), and the score function is used to score the model based on some condition.
By conditioning the generative model on the observed molecule, we restrict the samples to only those molecules that are similar to the specified molecule. The Metropolis-Hastings algorithm then explores the space of possible molecules and generates samples according to their weights (probabilities) based on the conditioned model.
4 Discussion
In a recent review of molecular representation, four main types of molecular representation are considered [Wigh et al., 2022]. These are Strings, Graphs, Fingerprints and Variational Auto Encoders. Here, relative advantages and disadvantages of all four approaches will be investigated and evaluated against the proposal here.
4.0.1 Unsorted Advantages compared to other representaions
In the Molecule ADT, whilst the same problem may exist with atom IDs for instance, there is a structured way to represent symmetries with group theoretic typeclasses.
4.1 The three types of representation
4.1.1 Storage
The ADT representation may have some drawbacks in terms of storage efficiency compared to more compact file formats like molfiles. The verbose nature of the ADT syntax could result in larger file sizes, especially for large molecular datasets. This could impact storage requirements and data transfer times. Nevertheless, the benefits of the ADT’s expressiveness and extensibility may outweigh the storage efficiency concerns in many cases, particularly when the accurate representation of complex molecular features is crucial.
4.1.2 Transmission
For transmission purposes, the ADT representation can be serialized into a human-readable format using the show and read functions in Haskell. This allows for the easy exchange of molecular data between cheminformaticians using plain text formats, such as .hs files. In the context of molecular representation transmission, it is important to distinguish between the serialized format of the data and the underlying data type itself. While the proposed algebraic data type (ADT) representation can be serialized into a human-readable string format using the show and read functions in Haskell, it is crucial to recognize that the ADT itself is not merely a string.
The ADT representation is a structured and type-safe data type that encapsulates the complex features and properties of molecules. The type annotations within the ADT provide a rich semantic context that goes beyond the limitations of plain strings. These type annotations ensure that the data is well-formed, consistent, and adheres to the specified structure and constraints of the molecular representation.
In the modern era of cheminformatics, it is essential to distinguish between the data itself and the type of data being used. Treating molecular representations as mere strings can lead to ambiguity, inconsistency, and potential errors in interpretation and manipulation. Strings lack the inherent structure and semantics that are crucial for accurately capturing and reasoning about molecular properties and behaviors.
By leveraging the power of ADTs and type systems, cheminformaticians can work with molecular representations in a more precise, reliable, and meaningful way. The type annotations serve as a form of documentation and validation, ensuring that the data conforms to the expected structure and constraints. This enables more robust and error-free manipulation of molecular data, reducing the chances of introducing inconsistencies or invalid states.
Moreover, the use of ADTs facilitates the development of type-safe and composable functions and algorithms that operate on molecular data. Type-driven development allows for the creation of reusable and modular code, promoting code clarity, maintainability, and scalability. This is particularly valuable in the context of complex cheminformatics workflows, where the ability to reason about the types and structure of data can lead to more efficient and reliable implementations.
In contrast, treating molecular representations as plain strings can hinder the development of robust and maintainable cheminformatics software. Strings provide no guarantees about the structure or validity of the data, leading to the need for extensive runtime checks and error handling. This can result in more brittle and error-prone code, making it harder to reason about the correctness and behavior of the software. Therefore, in this age of advanced computational tools and techniques, it is imperative to embrace the distinction between data and the type of data being used.
4.1.3 Computational
The ADT representation offers several computational advantages over other molecular representations like strings, graphs, bond electron matrices, and fingerprints. The structured nature of the ADT allows for efficient and type-safe manipulation of molecular data within a programming environment. The use of algebraic data types enables pattern matching and recursive processing, which can lead to concise and expressive code for complex molecular operations. The strong static typing of the ADT representation catches potential errors at compile-time, reducing the chances of runtime failures and increasing the overall reliability of cheminformatics software. In contrast, string-based representations like SMILES or SELFIES may require additional parsing and validation steps to ensure the correctness of the molecular structure. Graph-based representations, while capturing the connectivity of atoms, may lack the expressiveness to represent more nuanced features like stereochemistry or resonance structures. Bond electron matrices and fingerprints, although useful for specific applications like similarity searching or machine learning, may not provide a complete and detailed representation of the molecular structure.
Apart from the obvious benefits of being able to represent molecules which cannot currently be represented using string based representations (even with the creation of new and complicated grammars), our ADT confers other crucial benefits:
  1. 1.
    By directly synthesizing molecules in accordance with the grammar defined by the ADT, machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation into another format.
  2. 2.
    Our ADT is relatively extensible, meaning that the cheminformatics community can modify the grammar of the ADT in accordance with their needs, (e.g. to represent polymers or tautomeric molecules).
4.2 Strings
4.2.1 The future of SELFIES and of molecular representation
Some of the authors of SELFIES happen to have produced [Krenn et al., 2022], which points to future representational projects for molecules. In the future of SELFIES, [Krenn et al., 2022] they suggest that strings can represent Turing complete programming languages. Why not just drop the word strings (which have limitations as a data type) and use pre-existing well developed programming languages? They title project 8 as; “Future Project 8: Introducing “molecular programming languages” which they mention means adding 3D information to the representation, with a meta ‘for’ keyword, so despite the title, is in no way an actual suggestion of a practical molecular programming language, in the usual sense of programming language.
The authors of [Krenn et al., 2022] describe programming languages as strings, but indeed more formally they are sets of strings. Clarification may be worthwhile to ensure integrity of discourse between communities. Strings in this article are a data type that support a very limited number of operations, such as [CharAt, Concatenation, Contains, Equality, Find, indexof,
instrlength, locate, Lowercase, partition, replace, reverse, split]. Whilst software engineers write programs with text, it would be misleading to suggest that one reflects on applying string operations to the text of our programs. This would likely break the syntax of the code, rendering the code without formal denotation, let alone well-formed structure. The programmer thinks about and reasons on the object level of the program itself in accordance with some framework determined by the Church-Turing thesis, not on its textual syntactic representation. They mention “Python is of course a Turing-complete language, which means, strings can encode the most powerful computational algorithms.” which suggests strings are of prime importance [^7] The authors also write like that in this substack article: https://aspuru.substack.com/p/molecular-graph-representations-and. Looking at programming languages through the Church Turing thesis, or data and functions which reflects the division that Von Neumann suggested, is more faithful to the jargon used by the programming languages community, and can help to avoid confusion. It is important to remember that all of the “power” of strings occurs with Abacus Machines or the lambda calculus, which is not just a string. Not to mention that one dimensional strings are difficult to read. They also suggest strings as the “esperanto” of chemistry, which seems absurd thinking about strings such as those in Figure LABEL:morphine.
They also propose “BEYOND ORGANIC CHEMISTRY: COMPLICATED BONDS”, which advocates using the Dietz representation, which is precisely the one in use in this articles ADT.
They also suggest “Future Project 9: A 100% robust programming language” by finding a syntax for a programming language that however you mutate the source program code, a valid program is always produced. It seems unclear how to do this except with a trivial interpreter or parser. It seems this is an attempt to generalise their own robustness of mutation to the most general data structure; programs themselves. Despite their title, this does not correspond to what has been proposed in this article.
5 Future Work
The representation proposed is for cheminformaticians, not chemists, and as such the representations used by chemists, need not even be viewed on computer screens. For chemists, inspiration can be taken from work on Tangible Bits or Morphological Computation [Ishii and Ullmer, 1997], resulting in virtual reality or high frequency sound waves to levitate balls and sticks representing Lewis structures [Marzo et al., 2015].
Addition of optional tautomeric data to the representation would be very helpful. In [Pospisil et al., 2003] and [Sitzmann et al., 2010] the importance of tautomers is made clear. In [Sitzmann et al., 2010], a database of 680 million tautomers is established. They are in medications from sildenafil (Viagra) [Pajka et al., 2022], to Warfarin, to Remdesivir (used for Covid 19), Tetracyclines and antibiotics. A central problem in drug design is predicting ligand binding to proteins [Boyles et al., 2020]. Tautomers are important because tautomers affect the conformation, which affects the protein ligand attachment. [Pospisil et al., 2003] states “Tautomers are often disregarded in computer-aided molecular modeling applications. Little is known about the different tautomeric states of a molecule and they are rarely registered in chemical databases. Tautomeric forms of a molecule differ in shape, functional groups, surface, and hydrogen-bonding pattern. Calculation of physical–chemical properties and molecular descriptors differ from one tautomeric state to the other as it is demonstrated with an example of the log P calculation, similarity index, and the complementarity pattern to the targeted protein. Considering tautomery in ligand–protein interactions therefore has a significant impact on the prediction of the ligand binding using various docking techniques. This article points [to] hitherto unaddressed issue of tautomerism in computer-aided drug design
One avenue for future work is the integration of the proposed algebraic data type (ADT) representation with geometric deep learning techniques to exploit the inherent symmetries and invariances present in molecular structures [Atz et al., 2021]. By leveraging the mathematical properties of ADTs, such as group invariances and equivariances and their verification in Haskell through typeclasses, one could develop neural network architectures that are specifically tailored to the unique characteristics of those molecules. For instance, the atomIDs used in the ADT representation are essentially arbitrary labels and should not influence the learning process. Neural networks can be designed to be invariant to permutations of atomIDs, ensuring that molecularly equivalent structures are treated identically regardless of the specific ID assignment. Similarly, certain molecular features, such as the atomSymbol, are both translation and rotation invariant, while others, may only be translation invariant. By incorporating these symmetries directly into the neural network architecture, we can significantly reduce the amount of training data required and improve the generalization capability of the models. Geometric deep learning techniques, such as equivariant neural networks and graph neural networks, have shown great promise in capturing the complex spatial and relational structure of molecules.
To further demonstrate the practical utility and superiority of the proposed algebraic data type (ADT) representation, an important direction for future work is to empirically validate its performance on real-world machine learning tasks in cheminformatics. By benchmarking the ADT representation against existing approaches, such as SMILES and SELFIES, on tasks like molecular property prediction, virtual screening, and de novo drug design, we can quantitatively assess its effectiveness and identify areas for further improvement. Moreover, the current ADT representation can be extended to incorporate additional quantum mechanical properties, such as spin and angular momentum, into the electronic configuration of atoms. This enhancement would enable a more accurate and comprehensive description of the electronic structure, potentially leading to improved predictions of molecular properties and reactivity. To further strengthen the type safety and invariance guarantees of the representation, exploring the use of dependently typed programming languages, such as Agda, could be highly beneficial. Dependent types allow for the encoding of more sophisticated invariants and constraints directly into the type system, ensuring that only valid and consistent molecular structures are expressible. This increased level of type safety can help catch potential errors at compile-time and provide stronger guarantees about the correctness of the representation. Another important extension to the ADT representation is the inclusion of polymers, which are crucial in many areas of chemistry and materials science. By leveraging the power of lists and recursive data structures, the representation can be augmented to handle polymeric molecules seamlessly. This extension would enable the modeling and analysis of complex macromolecular systems, opening up new possibilities for polymer design and optimization. Finally, to enhance the modularity and composability of the ADT representation, exploring techniques like modular syntax trees and the “Data Types a la Carte” approach [Swierstra, 2008] could prove highly valuable. By decomposing the molecular representation into smaller, reusable components, we can create a more flexible and adaptable framework. This modular design would allow for the easy integration of new features, the creation of domain-specific languages for molecular manipulation, and the development of reusable libraries and tools for cheminformatics. By pursuing these avenues of future work, we can further strengthen the capabilities and impact of the ADT-based molecular representation, positioning it as a powerful and versatile tool for advancing the field of cheminformatics and accelerating the discovery of novel molecules with desired properties.
Some of the most important drugs in history, such as Warfarin, Allopurinol (a medication to treat gout which has nine tautomers), and antibiotics, display tautomerism [Porter, 2010], but existing transmission representations such as SMILES, and others which rely on molecular graphs, cannot represent tautomeric molecules in a single structure. Furthermore graph structures alone cannot straightforwardly be extended to resonant structures, tautomers, 3D information and organometallics [David et al., 2020] & [Wigh et al., 2022]. Dietz [1995]. This could be useful further work.
Further work could explore several avenues to enhance the proposed algebraic data type (ADT) representation and its applications. Integrating Haskell’s Lens library [Steckermeier, 2015] could simplify the manipulation of complex molecular structures by providing a composable and modular way to access and modify nested data fields. Implementing backpropagation using differentiable programming [Elliott, 2009] would enable the training of deep learning models directly on the ADT representation, leveraging the structure and semantics of the molecular data. To ensure the practical value and completeness of the ADT, empirical verification through extensive feedback and collaboration with chemists and cheminformaticians is crucial. This process would involve presenting the ADT to domain experts, gathering their insights, and iteratively refining the representation based on their requirements. Assessing the ADT’s performance on a prediction task using the QM9 chemical benchmark dataset [Pinheiro et al., 2020] would provide valuable insights into its effectiveness in capturing relevant chemical information compared to existing state-of-the-art methods. Furthermore, integrating the ADT with quantum sampling techniques could lead to breakthroughs in predicting molecular properties by efficiently simulating complex quantum systems. However, realizing this goal would require close collaboration between experts in quantum computing, chemical informatics, and molecular modeling to develop efficient quantum circuits and optimize the mapping between the ADT and quantum hardware. These future directions highlight the potential of the ADT representation to revolutionize various aspects of computational chemistry and materials science.
6 Conclusion
In this work, we have presented a novel approach to molecular representation using algebraic data types (ADTs) embedded in a pure, type-safe, functional programming language like Haskell. This ADT-based representation offers several compelling advantages over existing string-based representations such as SMILES and SELFIES. Firstly, the ADT representation inherently defines a grammar for molecular structures, ensuring that only semantically meaningful and valid molecules can be constructed. This eliminates the need for additional validation steps and reduces the likelihood of generating nonsensical or impossible structures. Secondly, the ADT representation is highly extensible, allowing for the seamless incorporation of additional molecular features and properties without requiring the creation of entirely new syntaxes or grammars. This flexibility enables the representation to evolve and adapt to the ever-expanding knowledge and requirements of cheminformatics. Thirdly, the ADT representation naturally supports the inclusion of 3D atomic coordinate data, enabling the calculation of crucial geometric properties such as dihedral angles and bond lengths. This is particularly advantageous for machine learning tasks that rely on 3D structural information, as opposed to SMILES and SELFIES which primarily focus on 2D connectivity. Furthermore, the ADT representation allows for the accurate modeling of resonance structures, tautomers, and organometallic compounds by incorporating the concept of delocalized electrons and multi-center bonding, inspired by the Dietz representation. This level of generality surpasses the capabilities of SMILES and SELFIES, which struggle to represent such complex molecular features in a single, unified structure. Additionally, the integration of the ADT representation with probabilistic programming paradigms opens up new possibilities for performing Bayesian inference and uncertainty quantification over molecular structures. By combining the expressive power of ADTs with the ability to represent and reason about probability distributions, we can develop rich, probabilistic models of molecules that capture the inherent uncertainty and variability associated with molecular properties. This probabilistic treatment is particularly valuable for drug discovery and design, where the ability to quantify uncertainty and incorporate prior knowledge is crucial for making informed decisions. Moreover, the use of a lazy functional programming language like Haskell offers potential efficiency improvements by leveraging lazy evaluation strategies. This is especially beneficial when performing simulations or sampling from molecular orbitals, as computations are only performed when strictly necessary. Finally, the ADT representation, being embedded in a statically-typed language, benefits from strong type safety guarantees. This helps catch potential errors at compile-time, reducing the chances of runtime failures and improving the overall reliability and maintainability of cheminformatics software. In conclusion, the proposed ADT-based molecular representation offers a powerful and flexible alternative to existing string-based representations like SMILES and SELFIES. By leveraging the strengths of functional programming, type safety, and extensibility, this approach has the potential to revolutionize the way we represent, manipulate, and reason about molecular structures in the field of cheminformatics.
7 Data, Funding, Acknowledgements and Abbreviations
7.1 Availability of data and materials
The library with instructions for its basic usage can be found at the following URL: https://github.com/oliverjgoldstein/chemalgprog
7.2 Funding
UK EPSRC DTP Scholarship.
7.3 Competing interests
There are no competing interests.
7.4 Authors Contributions
Oliver Goldstein provided the impetus and concept, as well as the entirety of the implementation.
7.5 Acknowledgements
I am grateful for conversations with Michael Harding about chemistry, and Sam March for helping with the restructuring and re-writing of parts of the article. I am grateful to Jack Liell-Cock for his helpful discussions about Haskell and graphs. Furthermore, I am grateful for the email exchanges with Dr. Richard Apodeca and his blog.
7.6 Author’s information
Oliver Goldstein is a Research Student at Oxford University reachable at oliverjgoldstein@gmail.com or oliver.goldstein@reuben.ox.ac.uk.
7.7 List of abbreviations
ADT(s) = Algebraic Data Type(s)
SMILES = Simplied Molecular Input Line Entry System