### v2

# SMILES & SELFIES have to go: Representation of Molecules via Algebraic Data Types

Abstract
The Algebraic Data Type (ADT) can be used as a computational framework for molecular representation for the purpose of advancing tasks in cheminformatics. This can include generative modles in the context of Bayesian machine learning via probabilistic programming. The ADT that we put forward, implements the ‘Dietz’ representation for molecular constitution via multigraphs of electron valence information, and uses 3D coordinate data to provide stereochemical information, easily enabling the representation of complex molecular phenomena such as organometallics, multi-center bonds, delocalized electrons, and resonant structures. Unlike traditional string-based representations such as SMILES and SELFIES, the ADT is much more flexible, yet retains desirable qualities from type-safety to seamless integration with Bayesian Probabilistic Programming. An extensive criticism of both SMILES and SELFIES, in this article, is given, along with criticisms of the so-called Future of SELFIES. An open-source library implemented in Haskell demonstrates the ADT along with experimental extensions demonstrating its use in reaction modelling, group theoretic applications, and integration with LazyPPL, a lazy probabilistic programming library. Also provided as an extension is the ability to represent electronic structures, including shells, subshells, and orbitals, significantly expanding its representational scope compared to other representations in the literature. These features position the proposed ADT as a robust alternative to existing molecular representations, addressing limitations such as inadequate support for 3D information and syntactic invalidity while offering a platform for innovative cheminformatics research. Accompanying discussions about the meaning of a ‘representation’ are included. The fully functioning GitHub library can be found at https://github.com/oliverjgoldstein/chemalgprog.
keywords:
Molecular Representation, Cheminformatics, Algebraic Data Types
1 Introduction
The choice of molecular representation directly affects cheminformatics tasks such as novel molecule discovery and molecular property prediction [1]. More expressive representation enables the exploration of larger regions of chemical space, both as input to and output from machine learning models. However, representational ability alone does not determine the effectiveness of a molecular representation in ML applications. Computational efficiency of optimisation algorithms, model choice and the amount of data provided are also important factors, along with the ease—or even possibility of integrating representations with techniques such as geometric deep learning [2] or Bayesian machine learning [3]. In cheminformatics, we believe that when it comes to molecular representation, “the medium is the message.”.  [^1] From Marshall McLuhan. More specifically, the context of communication decides the operations which the transported data supports.
Despite decades of research, string-based representations such as SMILES [4] and SELFIES [5] remain the most common representations for ML tasks [6, 7, 8], despite well documented short-comings in representational ability and with syntaxes themselves. For instance, standard SMILES strings cannot represent complex bonding systems, resonance structures, or delocalized bonding, and do not include any 3D information, with only a limited ability to describe stereochemical features. In addition, not all SMILES strings correspond to valid molecules, meaning that the output of ML models is, in many cases, gibberish [5], and small changes in SMILES strings can result in large differences in molecular structure, making property prediction difficult. Attempts to build upon strings have led to a profusion of new grammars and syntaxes to represent molecular phenomena, a problem in itself, such as BIGSMILES or CURLYSMILES etc.
While proposing future work to address the limitations of string-based representations such as SELFIES, [5] suggest developing a “molecular programming language”. Reasoning that strings can encode Turing-complete programming languages (e.g., Python source code), they suggest SELFIES could itself become a programming language, extended with a syntax that allows for recursion (for/while loops) and logical statements. Although this reasoning appears to conflate strings as a data structure with strings as a sequence of ASCII characters, we ask the question: if molecules can be represented as data within well established programming languages, we ask, why not simply use one?
In this paper, we critique contemporary computational representations, their application and challenges with ML, and some of the proposed solutions relating to string-based representations — we then present a solution with a novel framework: the Algebraic Data Type (ADT), written in the functional and typed language, Haskell.
The ADT has features of the Dietz representation for constitution [9], using multigraphs of atomic valence information. It contains optional features such as 3D atomic coordinate data to provide functions such as conformation energy prediction and configuration of molecules, as well as quantum information in the form of orbitals, which is a first in the literature.
The ADT is able to represent complex bonding systems (e.g. organometallics), resonance structures, and delocalised bonding. We further demonstrate extending the ADT in modelling reaction networks, and by using the ADT as an instantiation of a mathematical group. We also construct a probabilistic model in LazyPPL that generates and scores molecules based on arbitrary metrics to show the ADT’s ability to be used with probabilistic programming.
We also consider the impact the ADT has beyond computation, and consider the impact of the ADT as a representation for storage (and retrieval) of molecular information, and for transmission of molecular information between cheminformaticians.
While not intended as a definitive implementation of molecular representation, the ADT demonstrates the potential of algebraic approaches to reshape computational chemistry and invites the cheminformatics community to explore its capabilities and extend its utility.
2 Forms of Molecular Representation
In this paper, we distinguish three different usages of the term “molecular representation”:
  1. 1.
    Storage representations refer to the format for storage of molecular data on-disk, that is, without interpretation by a computer program.
  2. 2.
    Transmission representations relate to the way molecules are sent between cheminformaticians, and may depend on the medium of communication used.
  3. 3.
    Computational representations refer to how molecules are represented inside programming language or program, specifically, the data type or types which are used to represent the program.
For instance, a molecule may be stored in a database as a molfile, be processed with cheminformatics library such as RDKit[10] into another internal representation, and be then transformed into a molecular graph, or a fingerprint, for processing in a machine learning model[11, 12].
String based representations, such as the ubiquitous SMILES[4], may be used for all three. They are available in databases, may be downloaded and transmitted as strings, and also used as direct input for machine learning models[11, 12, 1], which as we later discuss, we regard as a mistake.
We regard the distinction between underlying data, and the data type to be crucial in cheminformatics, as it is the data type which constrains the operations which may be performed on data, reflects the structure and semantics of the data, and affects the efficiency and utility of computational tasks. Failure to make this distinction can lead ambiguity, inconsistency, and potential errors in interpretation and manipulation for meaning, as we later discuss in detail.
3 Contemporary Molecular Representations
In this paper we consider Molecular Graphs, Molecular Hypergraphs, the Dietz representaion (via multigraphs), string based representations and their numerous extensions, fingerprints, Variational Auto Encoders (VAEs), and computation tables.
3.1 Molecular Graphs
Many common molecular representations are built around molecular graphs and, as such, inherit strengths and limitations of molecular graphs for molecular representation [11]. String based representations such as SMILES and SELFIES use a number of rules to decode ASCII strings which correspond to molecular graphs, and MDF molfiles have “blocks” which correspond to sets of atoms and sets of bonds.
Definition 1 (Graph).
Formally, in set-theoretic notation, a graph, $G=(V,E)$ is a pair of sets $V$ and $E$, where $V$ is a set of vertices and $E$ is a set of pairs of vertices {u,v}, called edges.
A molecular graph is an instance of a graph in which the vertices are atoms, and the edges are bonds between atoms. Molecular graphs are connected, undirected, and labeled, where vertices may be labeled with information such as chemical element, and charge; and edges may be labeled with bond type (e.g. single, double or aromatic.), bond order, or 3D information (e.g. R/S isometry) [9, 11].
One benefit of representing molecules computationally as graphs is that insights from chemical graph theory can be directly leveraged by machine learning models. However, not all molecules can be represented by graph models, including molecules with complex bonding, organometallics, resonant structures and tautomerism.
Edges in graphs are between two vertices only, so molecular graphs have a limited ability to represent delocalised and multi-centre bonding, including ionic bonds and organometalics [11, 12]. Although it is possible to label edges with, e.g. rational bond orders, the contribution of individual electrons to multiple covalent bonds would not be described by such a graph. Attempting to express delocalised bonding by labeling edges of a molecular graph with bond types would require a sufficient number of predefined bond types, which would need to be extended for each new bonding relation not described.
Evidence of this approach extends back decades, as Dietz said ([9]):
“Enhancing the expressiveness by including a new bond type for every exceptional case is certainly not a very elegant solution.”
Using a single structure, graph models cannot represent tautomers, so instead require a work-around such as using multiple structures to represent the different forms of molecules.
Hypergraph and multi-graph models have been proposed to solve some issues with graph models [13], but they have seen limited use [11].
3.2 Molecular Hypergraphs
A hypergraph, is a graph where edges may connect any number of vertices. In a molecular hypergraph, edges with two vertices represent localised bonding, and edges with more than two verticies represent delocalised bonding, and are labelled with the number of electrons shared between the corresponding atoms.
From Dietz:
“A hyperedge containing more than two atoms gives us no information about the binary neighborhood relationships between them. That means we have no information at our disposal conceming the way in which the electrons are delocalized over these atoms” [9].
Hypergraphs would indicate an electronic “soup” with out knowledge of the manner electrons are delocalised. Hypergraphs remain uncommon.
3.3 Multigraphs
Multigraphs are not widely used as a standard computational representation in cheminformatics or chemical graph theory, despite the protestations of Dietz [9]. Existing tools and frameworks generally rely on labeled graphs or weighted graphs to encode bond multiplicities and types.
Theoretical discussions of multigraphs in chemical graph theory arise in the context of molecules with multiple bond types or delocalized bonding, but computational implementations of multigraph representations are not widespread.
Definition 2 (Multi-Set).
A multiset (or bag) is a generalization of a set where elements are allowed to appear more than once. Formally, a multiset $M$ over a universe $U$ is defined as a function:
$M:U\to\mathbb{N}$
where $M(x)$ denotes the multiplicity of element $x$, i.e., the number of times $x$ appears in the multiset. A finite multiset can be represented as:
$M=\{a_{1}^{m_{1}},a_{2}^{m_{2}},\dots,a_{n}^{m_{n}}\}$
where each element $a_{i}$ appears $m_{i}$ times.
The cardinality of a multiset $M$ is given by:
$|M|=\sum_{x\in U}M(x)$
which counts the total number of elements, including repetitions.
Definition 3 (Multi-Graph).
Formally, in set-theoretic notation, a multi-graph $G=(V,E)$ is an ordered pair of a set of vertices, $V$, and a multi-set, $E$, of edges.
By using a multigraph approach, it is possible to label bonds with information about the binary bonding relationships between atoms, as well as describing the delocalised electrons that persist over a subset of atoms, which is the approach taken by Dietz [9].
3.4 Dietz representation
The Dietz representation was designed to provide a computational framework capable of representing any chemical structure that can be described by a structural formula. It was developed to overcome the limitations of graph-based and hypergraph-based methods, including string representations that correspond to molecular graphs [9].
The paper describes the “constitution” of a structure in terms of the unshared valence electrons of atoms and the bonding relationships between them. Additionally, the terms “conformation” and “configuration” are used to describe different aspects of the relative spatial relationships between atoms. Conformation refers to the spatial arrangement of atoms that can change without breaking chemical bonds, such as rotations around single bonds, while configuration refers to the fixed spatial arrangement of atoms, including stereochemistry, which requires breaking and reforming bonds to alter. These spatial relationships are illustrated using a unique “paddle wheel” representation.
As the proposed ADT uses atomic coordinate data to describe 3D information of molecules, only the “constitution” of the Dietz representation is discussed here.
The constitution of a structure is represented in set theoretic notation, with two sets. Our ADT includes type information, which prevents invalid molecules as will be discussed. The first set contains all atoms in the structure as triples, containing the number of unshared valence electrons of the atom, a numerical index, and the atomic symbol. The second set describes “bonding systems”, or bonding relationships between atoms, and contains pairs consisting of the number of shared valance electrons, and a set of edges representing bonds over which the electrons are shared.
Definition 4 (Constitution).
The constitution of a molecule is an ordered pair:
$C=(V,B),$
where $V$ is the set of vertices, representing the atoms in the molecule, and $B$ is the set of bonding systems, representing the bonding relationships between atoms.
Each vertex in V is a triple:
$V=\{(u,j,A)\mid u\in\mathbb{N}_{0},\,j\in\mathbb{N},\,A\in\text{Symbols}\},$
where:
  • u is the number of unshared valence electrons for the atom,
  • j is a unique index, and
  • A is the atomic symbol (e.g., $H,C,O$).
Each bonding system in B is a pair:
$B=\{(s,\{\{x,y\}\mid x,y\in V\})\mid s\in\mathbb{N}\},$
where:
  • s is the number of shared electrons in the bonding system, and
  • {{x, y}} is a set of unordered pairs of vertices representing bonds over which the electrons are shared.
Localised bonding is represented by a bonding relationship containing a single set of edges. For instance, a single covalent bond between two atoms would be represented as follows: $(2,\{\{a_{1},a_{2}\}\})$, where $a_{i}$ is the index of each atom.
Delocalised bonding is represented by a bonding relationship containing multiple set of edges. For instance, the delocalised pi electrons of a benzene ring would be described by a bonding relationship of the form $(6,\{\{c_{1},c_{2}\},\{c_{2},c_{3}\},\{c_{3},c_{4}\},\{c_{4},c_{5}\},\{c_{5},c_{6}\},\{c_{6},c_{1}\}\})$, where $c_{i}$ is the numerical index of each carbon atom in the ring.
This representation is equivalent to a multigraph with additional labelling or metadata to describe the groups of edges with shared electrons.
Example 1 (Constitution of Hydrogen).
The constitution of hydrogen
$V(H_{2})=\{(0,1,H),(0,2,H)\},$
$B(H_{2})=\{(2,\{\{1,2\}\})\}\vskip 6.0pt plus 2.0pt minus 2.0pt$
There are two hydrogen atoms, labelled 1 and 2, each with zero unshared valance electrons in the atomic core, and a single covalent bond with two shared valance electrons.
Example 2 (Constitution of Benzene).
The constitution of Benzene:
$V(Benzene)$ =$=\{(0,1,H),(0,2,H),(0,3,H),(0,4,H),(0,5,H),(0,6,H),$
[…]
The delocalised pi electrons of the benzene ring are described in a single bonding system, as the final element of the set, and separately from all other covalent bonds in the structure.
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original paper. The lack of labeling requirements for delocalised bonding stands in contrast to string based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
Despite satisfying the three requirements identified in [1] for representation beyond organic chemistry—namely, the ability to represent delocalized bonding, explicit labeling of terminal hydrogen atoms, and advanced stereochemistry such as in cumulenes—Krenn et al. describe the Dietz representation (apparently mischaracterized as using hypergraphs) as one that “leads to complicated nested sets of brackets that may be hard to comprehend” [1].
Although set-theoretic notation may be unfamiliar, the definition or “grammar” of the constitution of a molecule can be efficiently described. This stands in contrast to SMILES strings or SELFIES strings (which as an additional challenge contains overloading of symbols in the grammar [1]).
As Krenn [1] puts it:
The same applies to other weak interactions such as hydrogen bonds, raising important questions as to which interactions should actually be captured in a digital molecular representation as a “bond” (and which should not) and how to automatically detect them from a set of atomic coordinates, ultimately leading to a rather arbitrary distinction between bonded and non-bonded. To quote Democritus: “Nothing exists except atoms and empty space; everything else is only opinion”.
3.5 SMILES
SMILES (Simplified Molecular Input Line Entry System, [4]) encodes molecular graphs as ASCII strings through a series of rules, representing atoms and bonds as characters and symbols. Introduced in the 1980s, SMILES was designed to facilitate chemical database searching and to enhance portability, and has achieved significant popularity in cheminformatics [11, 12, 1, 14, 15, 16, 17]. However, increasing demands of modern machine learning and cheminformatics reveal challenges that extend beyond the scope of SMILES’ original design.
As molecular graphs, SMILES inherit the limitations of this representational framework. While SMILES can represent aromaticity through specific syntax[4], it cannot encode delocalized bonding beyond such cases. Resonance and tautomerism must either be simplified into a single structure or require multiple distinct structures to represent each form.
The encoding of stereochemistry in SMILES has notable constraints. While tetrahedral centers are well-supported through isomeric SMILES, a common extension, the encoding of more complex stereochemical features, such as dynamic or rotational stereoisomerism seen in atropisomers, remains unsupported [18]. Isomeric SMILES can represent certain forms of axial chirality in cumulenes and handle non-tetrahedral stereocenters, including allene-like, trigonal-bipyramidal, and octahedral configurations [18]. However, these representations rely on 2D graph-based encoding and lack support for nuanced 3D spatial arrangements required for complex organometallic systems or multi-center bonding scenarios. The absence of three-dimensional (3D) spatial data further limits SMILES’ utility for applications requiring dihedral angles or precise molecular conformations, a critical aspect of protein-ligand modeling and other machine learning tasks such as property prediction [19, 20, 21].
Beyond these limitations, SMILES suffers from unique structural and functional shortcomings such as syntactic invalidity [5, 11, 12], non-local encoding of features, and non-unique representations of molecules, all of which can impair machine learning models as the models must implicitly learn the SMILES syntax, which can be difficult, and require larger training sets and wasted computation [11, 22].
“Syntactic invalidity” means that not all valid SMILES strings correspond to valid molecules. In Machine learning contexts, algorithms may output syntactic nonsense, and must re-sample or require larger training sets to ensure robustness [11, 6, 5]. Even though one paper claims that invalid SMILES should be seen as beneficial to chemical language models [23], it should be noted that the perceived improvement is only relative to comparisons between SMILES and SELFIES models, and likely reflects the models’ enhanced ability to infer SMILES syntax when generating valid molecules.
In SMILES, small changes in the string can lead to drastic changes in molecular structure, and vice versa – adjacent atoms in the molecule can be far apart in the string [22]. As well as impacting readability, the non-local encoding of structural information can make it difficult to write or even find a similarity metric of SMILES strings whose magnitude corresponds to a relevant similarity of molecules [22].
SMILES representations are non-unique: a single molecule can be expressed by multiple valid strings [4]. As well as complicating database searches [12], machine learning algorithms must discern the equality of different representations. Although some machine leaning approaches have attempted to utilize multiple SMILES for the same molecule in their training sets, and have claimed improvements to their models as a result [24, 25], this is again likely explained by the ability of the model to infer the SMILES syntax more easily. In other words, it exposes a weakness of SMILES for machine learning, not an advantage.
Representations in SMILES also do not give indications as to molecular symmetry. By example, the Cyclopentadienyl anion in SMILES is [cH-]1cccc1, but this does not indicate the rotational symmetry of the molecule, as is clearly shown in Fig. 1. The non-uniqueness of SMILES strings can be seen as 1cccc1[cH-] is also a valid representation, but in SMILES no explicit indication is given as to whether [cH-]1cccc1 = 1cccc1[cH-].
Figure 1: The Cyclopentadienyl anion. Although the structure displays clear rotaional symmetry, this is not indicated by the SMILES representaion: [cH-]1cccc1. Image Credit: Oliver Goldstein (Author).
Attempts to canonicalize SMILES, and the proprietary nature of the Daylight canonicalisation algorithm, has led to inconsistencies between implementations and research teams [11], though the OpenSMILES initiative has mitigated this issue by providing a standardized grammar.
Importantly, as a data type, strings by themselves provide no guarantees about the structure of the validity of the data.
While SMILES and its extensions remains useful tools for encoding and sharing molecular structures, its limitations constrain its applicability in modern cheminformatics and machine learning workflows. Extensions to SMILES to fix problems, and the problems they themselves introduce, are discussed in a later section.
3.6 SELFIES
SELFIES (Self-Referencing Embedded Strings) [5] solves arguably the most pressing issue with SMILES: syntactic invalidity. They create a “100% robust” representation, which would translate to soundness in the terminology of logic or computer science. In particular, they present a Context Free Grammar (CFG) where all sequences of terminal symbols represent molecular graphs (soundness) and all molecular graphs can be represented (completeness – but only up to the limited CFG they have created, see Fig. 1). The CFG of SELFIES can be translated into sum and product notation as every CFG can be represented by an algebraic data type [26].
In attempting to retain semantic as well as syntactic validity after mutation, SELFIES uses a table of derivation rules, with overloading of terms to encode valence information [1]. Although successful in maintaining semantic validity after mutation (with the given set of rules, for small biomolecules), and as sucessfully demonstrated with GANs and VAEs[5], the overloading of terms makes the grammar more complicated. The authors claim that SELFIES may be more readable then SMILES for large molecules, although they acknowlede that “Read-ability is in the eye of the beholder”[5]. We personally behold SELFIES as not being human readable, but invite readers to draw their own conclusions.
Machine learning algorithms would also have to implicitly learn the rules of the grammar [11]. Despite the authors claim that SELFIES may be more “machine readable” this remains a conjecture.
The overloading of symbols also represents a challenge for readability. When discussing previous line notations in [4], Weininger criticized attempts to save space with overloaded terms: “Many of the pitfalls in other line notations can be attributed to overuse of symbols and hierarchical rules based on the length of the final notation. Advances in computer hardware have made these restrictions obsolete.” In this sense, SELFIES represents a retrograde step.
State $\epsilon$ [F] [=O] [#N] [O] [N] [=N] [C] [=C] [#C] [Branch1] [Branch2] [Branch3] [Ring]
$X_{0}$ $X_{0}$ F 0 $X_{2}$ N $X_{3}$ 0 $X_{2}$ $NX_{3}$ $CX_{4}$ $CX_{4}$ $CX_{4}$ ign $X_{0}$ ign $X_{0}$
$X_{1}$ $\epsilon$ F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ ign $X_{1}$ ign $X_{1}$ R(N) R(N)
$X_{2}$ $\epsilon$ F =O $NX_{2}$ 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ B(N, $X_{5}$) $X_{1}$ B(N, $X_{5}$) $X_{1}$ R(N) $X_{1}$ R(N) $X_{1}$
$X_{3}$ $\epsilon$ F =O #N 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ B(N, $X_{5}$) $X_{2}$ B(N, $X_{6}$) $X_{1}$ R(N) $X_{2}$ R(N)$X_{2}$
$X_{4}$ $\epsilon$ F =O #N 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ B(N, $X_{5}$) $X_{3}$ B(N, $X_{7}$) $X_{1}$ R(N) $X_{3}$ R(N)$X_{3}$
$X_{5}$ C F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ $X_{5}$ $X_{5}$ $X_{5}$ $X_{5}$
$X_{6}$ C F =O $NX_{2}$ 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$
$X_{7}$ C F =O #N 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$
$N$ 1 2 3 4 5 6 7 8 9 10 11 12 13 14
Table 1: Derivation rules of SELFIES for molecules in the QM9 dataset. Table formatted to fit within the page.
Again, like SMILES, resonant structures are not directly expressible in SELFIES, and delocalized bonding (other than aromaticity) cannot be expressed, along with tautomers.
Since publication, SELFIES has added the ability to represent chirality, through importation of symbols from SMILES, although some forms produce errors in decoding. However, three dimensional information is not included, and it is unclear how to extend the grammar whilst retaining it’s simplicity and soundness. The spatial arrangement of ligands in square-planar stereochemistry, like that of cisplatin and transplatin, cannot be fully captured without 3D information, and so in exploring the chemical space, SELFIES remains unable to represent potentially crucial information for novel drug discover and property prediction, despite being sound.
Furthermore, E-Z stereochemistry cannot provide accurate dihedral or torsional angles [27]. Proteins rely on these angles for many tasks, which may be critical for machine learning tasks [28]. As Dayalan explains ([21]) “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [29]. Although SELFIES solves the syntactic invalidity problem of SMILES, many challenges remain.
3.6.1 The future of SELFIES and of molecular representation - Extensions of SELFIES and future projects
In “SELFIES and the Future of Molecular String Representations,” Krenn et al. (2022) [1], review the limitations of string-based molecular representations in cheminformatics and propose future directions, primarily centered around extending string-based methods.
They highlight several representational difficulties with current molecular string formats, including their inability to effectively capture macromolecular structures, crystal lattices, complex bonding (e.g., organometallics), advanced stereochemistry, precise conformations, and non-covalent interactions such as hydrogen and ionic bonding. These limitations, they argue, are not only present in SMILES and SELFIES but are features of most modern, digital molecular representations in use.
The authors also identify challenges specific to machine learning applications, such as issues arising from the non-local encoding of molecular features within strings, and the impact this has on the latent space of VAEs, as well as the impact overloading of symbols has the efficiency of generative models, as well as human readability.
3.7 Macromolecules and crystals
For macromolecules, Krenn et al. discuss existing approaches such as CurlySMILES, BigSMILES, and HELM, each of which introduces distinct syntactic modifications to account (or not, as the case may be) for repeating subunits, complex connectivity, and stochastic relationships between subunits. They propose BigSELFIES and HELM-SELFIES as future extensions to enable the use of generative models which take strings as input for macromolecular synthesis and design.
In the case of crystals, the authors propose yet another syntax, Crystal-SELFIES, based on the labeled quotient graph of the crystal structure. The formalism of the approach is largely unexplored in the paper, however relies on representation of subunits, in a related manner as with macromolecules.
3.8 Barriers to a general, digital, molecular representation
The representational limitations of current string-based methods are said to rely on three core assumptions:
  1. 1.
    The assumption of localized bonding, with standard valances
  2. 2.
    The implicit or non-explicit representation of terminal hydrogen atoms.
  3. 3.
    The inability to encode advanced stereochemical relationships.
These limitations, the authors argue (and we agree), are an inherent problem with string-based representations in their ability to represent complex bonding (such as with diborane and ferrocene) and advanced stereochemical features. Cis- and trans-platin are also given as examples of molecules current string-based representations cannot represent.
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in the Kenn paper:
Thus, it should be stressed again that in d- and f-block chemistry, as well as main-group organometallic compounds, it is often impossible to assign any particular bond orders without high-level quantum chemical calculations, due to the highly de localized nature of the bonding, where electrons are often spread out over a significant number of atoms, including the metal cen ter itself, the immediately coordinated atoms, and additional ligand groups.
Despite recognizing the need for a general digital molecular representation, capable of handling all of the above issues. The authors do not reconsider the suitability of string-based models, and instaed, the solutions presented primarily involve introducing new grammars, expanding syntax rules, and further overloading symbols, leading to increasingly complex and specialized representations.
3.9 “Molecular programming languages”
In Future Project #8, Krenn et al. (2022) propose a “molecular programming language”. They reason that strings can encode the powerful computational algorithms: “Strings can store Turing-complete programming languages: In the most general case, one can store the source code of computer programs as strings. For example, a Python file is a ‘simple string’, which is executed by the Python interpreter. Python is, of course, a Turing-complete language, which means that strings can encode the most powerful computational algorithms.”
Although the source code of a programming language can be represented in files of text, they should be distinguished from string data types. For instance, they do not support string data type operations, such as insertions, [Concatenation, Contains,indexof, Lowercase, partition, replace, reverse].
Furthermore, if the programming language is represented by a string, the question of how functions within the language are interpreted is unanswered. If they are also strings, they would at some point need to be executed by something other than a string. If they are not strings, then it remains to be understood why the authors claim that strings alone are enough to represent a Turing-Complete language, let alone the utility of such a project.
Given that Turing complete programming languages with support for recursion and logical statements already exist, it is unclear why SELFIES should be augmented to become a programming language in order to represent molecules, rather than simply using a programming language which already exists.
Future Project #9 goes beyond molecular programming languages, and proposes a programming language which is “100% robust” by finding a syntax for a programming language that however you combine elements in the instruction set, a valid program is always produced. It seems this is an attempt to generalise their own robustness of mutation to the most general data structure: programs themselves. However, it remains to be seen what sense this would have, let alone any sensible denotational semantics.
3.10 Impact on machine learning models and empirical verification
From a machine learning perspective, improving representation by expanding the grammar, creating new symbols, and overloading of terms can introduce problems, as deep learning models must implicitly learn the larger, more complicated syntax.
Unlike SMILES, VAEs trained on SELFIES have a robust latent space, where any vector can be decoded into a valid molecule. Despite this, it is not a guarantee that the latent space will be smooth, especially given the non-local encoding of graph information on strings, which both SELFIES and SMILES share.
We welcome the empirical question of the effect of overloading of tokens on machine learning models, which was raised in Future Project #2, and, in light of the result that training string models on randomised representations of molecules has on model results, suggest an experiment to determine the impact of implicit learning of the grammar in the future work section.
3.11 The Algebraic Data Type (ADT) as an Alternative
We believe that the ADT later introduced will solve all of the above mentioned problems, for representation, and for the impact on machine learning models. and will represent the general, digital molecular representation searched for, by moving beyond string-based paradigms of SMILES, SELFIES, and their numerous extensions.
3.12 Extensions of SMILES
Despite the introduction of DeepSMILES, which fixed the bracketing issue in SMILES strings, the syntactic invalidity problem was not solved completely [30].
New syntaxes or grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with only SMILES. For instance, there are many syntaxes from just one vendor called Daylight222www.daylight.com.
SMILES was introduced and subsequently standardised in OpenSMILES. CurlySMILES was introduced to annotate SMILES strings. CHUCKLES, its extension, CHORTLES, as well as BigSMILES were introduced to represent polymeric molecules [11]. SMIRKS was introduced as an extension of SMARTS to represent chemical reactions [31] – used in the Open Reaction Database. InChi V1, not part of Daylight, was introduced and can represent basic prototautemerism [32]. SMIRKS allows for string representation of reactions. It does not represent reaction features like temperature or pressure or catalysts. One feature it lacks is the ability to indicate the directionality of the reaction, which took a new language to be implemented ten years after the release of SMIRKS, called RinChi [11], [12]. This was further extended years later, with another language called ProcAuxInfo, which enables the storage of metadata relating to yields, temperature, concentration, and other reaction conditions. Although, this represents definite progress, the profusion of new syntaxes and grammars, and their creative backronyms, create avoidable complexity for the cheminformatician.
[33] introduces Balsa, a fully-specified language designed as a subset of SMILES. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
3.13 Fingerprints
Fingerprints involve encoding key features of a molecule, typically into a vector representation and can offer convenience for machine learning applications. However, a significant drawback of fingerprints is their inability to consistently represent tautomers, which are molecules that have the same atomic composition but different structural arrangements. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [34]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
Moreover, while fingerprints can capture the presence of specific substructures within molecules, they are susceptible to bit collisions and vector sparsity, which limit their overall representational power [22]. Bit collisions occur when multiple distinct substructures are mapped to the same bit in the fingerprint, leading to ambiguity and potential false positives. Vector sparsity arises when the fingerprint representation is high-dimensional, but only a small fraction of the bits are actively used, resulting in inefficient memory usage and computational overhead.
A fundamental issue with numeric vector representations, such as fingerprints, is the lack of inherent semantics associated with the individual numbers or floating-point values. Without additional contextual information or a predefined mapping, the numeric values in the vector do not carry intrinsic meaning. This lack of semantics may hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
In contrast, algebraic data types (ADTs) provide a structured and type-safe representation that inherently encodes the semantics of the molecular data. By defining the data structure and its associated types, ADTs constrain the space of possible values and operations, ensuring that only meaningful and valid transformations can be applied. This type-driven approach helps prevent errors and inconsistencies that may arise from arbitrary manipulation of numeric vectors.
Furthermore, the type information associated with ADTs can guide and constrain the machine learning and software development process [^3] See Conor Mcbride talk about this issue: https://www.youtube.com/watch?v=v2yXrOkzt5w. The type constraints restrict the search space and focus learning algorithms on meaningful and plausible solutions. This may well lead to more efficient and effective learning, as the model can leverage the structure and semantics embedded in the ADT representation.
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [35], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint. These algebraic properties such as group theoretic properties can be used to predict molecular properties. Variational Auto Encoders (VAEs) also have problems, [6], whom claim that obtaining structured output to be a hard and currently unsolved problem, even for a simple graph data structure.
3.14 RDKit as representation
One might also ask whether existing and widely used Cheminformatics libraries support representation. Behemoth Cheminformatics programming libraries, such as RDKit, [10] and lightweight molecular manipulation libraries that exist on top of these e.g. datamol [^4] https://datamol.io/ and https://www.rdkit.org/ (which builds on NumPY [36], Pandas, RDKit and Python 3+), Datamol explicitly [^5] https://doc.datamol.io/stable/tutorials/Preprocessing.html don’t advocate themselves as providing a prototypical representation. These libraries parse representations to computational representations and then apply functions to them, before outputting them to strings like SMILES.
In the RDKit’s Mol classes, molecules are stored as a graph in the form of an adjacency list, with optional extra information in the form of properties and ring information. The internal representation therefore inherits the limitations of 2D graph based representations already discussed for computational purposes. Bonds in the graph can be labeled with 22 pre-defined bond types, and limited stereochemical features can be described, again with labeling.
The computational representation of RDKit in C++ is complex, taking up over 400 lines of highly complex code and the molecule class taking up 873 lines of code as of 2022. There are no clear invariants, or guarantees about the management of side effects in snippets shown below. The ubiquitous presence of null pointers and highly abstruse C++ keywords, scattered throughout the code, insinuates points of failure of the program. It is unclear, whether and when the program may crash and what the consequences might be. The following snapshot of lines $792$ to $821$ in ROMol.h [^6] https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html which represents the molecule class, provides a brief glimpse into the extreme complexity of code:
1 #ifdef RDK_USE_BOOST_SERIALIZATION
[…]
3 //! @{
4 template <class Archive>
5 void save(Archive &ar, const unsigned int version) const;
6 template <class Archive>
7 void load(Archive &ar, const unsigned int version);
8 BOOST_SERIALIZATION_SPLIT_MEMBER()
9 //! @}
10 #endif
11
12 private:
13 MolGraph d_graph;
14 ATOM_BOOKMARK_MAP d_atomBookmarks;
15 BOND_BOOKMARK_MAP d_bondBookmarks;
16 RingInfo *dp_ringInfo = nullptr;
17 CONF_SPTR_LIST d_confs;
18 std::vector<SubstanceGroup> d_sgroups;
19 std::vector<StereoGroup> d_stereo_groups;
20 std::unique_ptr<boost::dynamic_bitset<>> dp_delAtoms = nullptr;
21 std::unique_ptr<boost::dynamic_bitset<>> dp_delBonds = nullptr;
22
23 friend RDKIT_GRAPHMOL_EXPORT std::vector<SubstanceGroup> &getSubstanceGroups(
24 ROMol &);
25 friend RDKIT_GRAPHMOL_EXPORT const std::vector<SubstanceGroup>
26 &getSubstanceGroups(const ROMol &);
27 void clearSubstanceGroups() { d_sgroups.clear(); }
28
29 protected:
30 unsigned int numBonds{0};
Listing 1: This is just a small fragment of the molecule class in RDKit. It demonstrates the sheer complexity and obfuscation of a concept that can be represented more simply in the ADT proposed in this article.
We aim to demonstrate that our ADT-based representation of molecules in Haskell offers a more concise, expressive, and type-safe approach. By leveraging the power of algebraic data types and the strong static typing of Haskell, we can define molecular structures in a declarative and compositional manner. The resulting code is more readable, maintainable, and less prone to runtime errors. The use of pure functions and the explicit management of side effects in a pure functional language such as Haskell, will enable easier reasoning about the behavior of the program and facilitate formal verification techniques.
Moreover, we claim that the ADT-based representation provides a clear separation of concerns between the data structure and the operations performed on it, interface versus implementation. This allows for a separation of concerns between those who are designing better algorithms for the data type and those who only need to use the data type. The use of higher-order functions and type classes in Haskell further enhances the reusability and composability of the code.
While cheminformatics libraries like RDKit and datamol have been widely adopted and have demonstrated their utility in various applications, their complex often OOP-based representations can hinder the development and maintenance of robust and reliable software. The proposed ADT-based approach offers a compelling alternative that prioritizes simplicity, type safety, and functional mathematical rigor.
3.15 Molecules and programming languages
Inductive Logic Programming (ILP), has seen usage in representing basic molecules [37], but the focus was doing rule induction over graphical data structures, with molecules being an example, rather than presenting ILP as a representational format for cheminformaticians. It is perhaps for this reason that it has had little uptake within the cheminformatics community.
Molecular representations using Inductive Logic Programming (ILP) have been explored in previous research, such as the work by Srinivasan et al. [37]. However, the application of machine learning techniques in conjunction with ILP-based molecular representations has not been extensively discussed in this early work. This lack of emphasis on machine learning aspects may explain the absence of the paper from recent reviews of molecular representations, such as the comprehensive survey by Wigh et al. [12].
In 2018, de Meent et al. [38] argue that the recent successes of artificial intelligence have been caused by the emergence of programming language tools, such as NumPy, that are able to automatically engage in the “tedious” derivation and calculation of gradients. Probabilistic programming languages build on this idea to provide a generic way to write arbitrary models and then calculate the uncertainty of parameter choices.
In the approach presented by Srinivasan et al. [37], molecules are represented as logical rules using an extension of the Prolog programming language. Prolog is a declarative language that enables the expression of logical relationships between variables using first-order logic. Programs written in Prolog consist of a complex network of facts and rules, which are constructed using logical primitives such as conjunction, disjunction, and universal quantification. However, the data types supported by Prolog are limited to lists, integers, floats, and strings, with minimal support for user-defined data types.
Prolog’s inference mechanism allows for querying the knowledge base by providing examples and evaluating their truth values based on the presence or absence of counterexamples (negation as failure). Nevertheless, it is important to note that Prolog and its extensions are not purely declarative, as they incorporate side effects through built-in predicates such as read, get, and assert. These side effects can succeed or fail independently of the problem being solved, introducing challenges in reasoning about the program’s behavior, particularly when dealing with complex rule bases.
The specific extension of Prolog used in [37] is called Progol, which employs inductive logic programming techniques to derive rules that describe the given data. It is worth clarifying that Progol should not be confused with ProbLog, a probabilistic logic programming language that focuses on modeling the probabilities of examples being true based on the probabilities of the underlying relations.
While the work by Srinivasan et al. [37] demonstrates the feasibility of representing molecules using logical rules in an ILP framework, the integration of machine learning techniques with these representations was not explored.
This makes the proposal here the first to suggest to cheminformaticians to use constructs rooted in functional programming languages to be directly involved as representation. The notion that “molecular programming languages” could be possible, occurred in an article about the future of SELFIES [1]. However, the sense attached to that project has been questioned previous to this.
Investigations of type safe functional programming for cheminformatics has occurred before. Recent work has suggested the use of OCaml in writing software for cheminformatics [39, 40]. However, the emphasis in both of these articles as well as the Haskell library Ouch [41], was not on molecular representations but more a tutorial on functional programming. This is highlighted by the papers introduction of lengthy tutorials on functional programming. Furthermore they do not show how Bayesian Inference can be performed on drug structures over algebraic data types which is essential in motivating efficiency improvements in not requiring exploration of syntactically nonsensical regions of molecular space.
4 Probabilistic programming
As discussed in a recent thesis on modular probabilistic programming, and I quote from it through this and the next paragraph [42], the Bayesian approach to probabilistic reasoning is known as Bayesian Inference, and is an increasingly popular method to perform machine learning that enjoys strong mathematical foundations. Bayesian Inference has useful properties, compared to deep learning based machine learning, such as interpretability, explainability and expressiveness with respect to various forms of uncertainty [43, 44]. At the heart of the Bayesian paradigm is the notion of a generative model. Generative models can account for the expressiveness of uncertainty and interpretability. In order to perform probabilistic modelling, knowledge about a domain and the uncertainty associated with that knowledge is encoded into a generative model, using expert knowledge. Inference algorithms estimate parameters of the model, that generate the observed data, by conditioning models on associated observed data. For instance, a climate researcher may create a basic model of a city when studying urban heat islands. The climate researcher would condition the model of the city with empirical data, such as the observed temperatures of the city, in order to infer values of interest such as latent city temperatures or factors that influence the temperature of a city. Probabilistic programming languages provide a generic framework to perform inference.
To highlight the importance of probabilistic programming languages, large multinational technology companies such as Microsoft, Google and Uber have recently started to invest in their own probabilistic programming languages (Infer.NET [45], Edward [46] and Pyro [47] respectively). Increasingly, machine learning use cases find their way into areas as diverse as critical defense national infrastructure [48][49] or medical diagnostics [50]. Society is therefore increasingly in a position that relies on programmers to produce correct machine learning related code. As probabilistic programming becomes more important, the code that underlies probabilistic programming languages will face increasing scrutiny. Fortunately, there are two important programming language research goals which aim to ensure the correctness of code; modularity and type-safety. In particular, they ensure that programmers can be certain early on – at compile time – that composing code will not cause programs to crash or introduce unintended (inference) algorithmic bias. These research goals have two effects; as programming libraries become safer and more modular, machine learning models will become cleaner, simpler and easier to maintain, which will additionally make machine learning more trusted, and perhaps more widely adopted.
In the context of molecular representation, probabilistic programming offers a natural way to capture the inherent uncertainty and variability associated with molecular structures and properties. By combining the expressive power of functional programming with the ability to represent and reason about probability distributions, PPLs enable the development of rich, probabilistic models of molecules.
5 Methodology: Laziness and The Algebraic Data Type.
An Algebraic Data Type (ADT) is a composite data structure formed by combining simpler data types using sum types (representing alternatives) and product types (representing combinations), which obey algebraic laws [51]. It is a fundamental concept in functional programming, with links to universal algebra, that provide a way to define complex, hierarchical, or (iso)recursive structures, first introduced in 1980 with HOPE [52, 51]. They define the data structure, and its type with which typeclasses then define the operations that can be meaningfully performed on that data, through instantiation.
Our molecular Algebraic Data Type is written in Haskell[53], a lazy, purely functional, type-safe language, with type classes and algebraic data types.
In cheminformatics, we often deal with large datasets of chemical structures (e.g. millions of molecules) and sometimes expensive analyses (e.g. molecular property calculations, conformer generation, docking runs). A lazy programming language like Haskell can help manage these challenges by delaying computations until they are truly needed, and by automatically reusing results once they have been computed.
In many programming languages, as soon as you write an expression—for example, a function call or a chemical property calculation—the language runtime will typically evaluate it immediately. In a lazy language, however, expressions are turned into thunks: little bundles of code waiting to be executed. The language defers (delays) performing the actual computation until it is absolutely required, such as when the result is needed to proceed with another step in the program.
Suppose we define a huge list of candidate molecules, each annotated with an expensive-to-calculate property (e.g. predicted solubility). In a strict language, as soon as we construct this list, the program might try to calculate the property for every molecule upfront, even if we only want the first few results. In a lazy language, by contrast, the expensive property calculations are wrapped up in thunks. These thunks remain unevaluated until you actually iterate over the list and access the relevant element.
Thus, if you only access the first 10 molecules in the dataset, only 10 calculations take place; the remaining 999,990 molecular properties remain as unevaluated thunks.
A second important aspect of laziness is sharing:
once a thunk has been computed, its result is cached (or memoized) so that future accesses do not repeat the same work. If multiple parts of your cheminformatics workflow depend on the same expensive molecular descriptor, the lazy evaluation mechanism ensures that once you compute it once, the result is re-used everywhere else automatically.
  • On-Demand Computation: We might have large pipelines combining multiple steps (e.g. filtering by molecular weight, docking simulations, then clustering). If we never end up needing a particular property, we waste no effort computing it in a lazy setting.
  • Improved Compositionality: With lazy evaluation, we can write “infinite” or “very large” data streams and process them in a streaming fashion without ever fully materializing them in memory. This is particularly helpful when investigating structure-generation procedures or enumerations in a conceptually infinite chemical space, since only the portion of the stream you consume is actually computed.
  • Parallelization and Caching: Lazy languages can simplify certain parallelization strategies since once a thunk is forced in one thread, its result can be shared across other threads (in purely functional languages like Haskell).
Our Algebraic Data Type represents molecular constitution with the Dietz representation, three dimensional chemical information via atomic coordinate data, and provides a novel representation of quantum information via the Orbital type. Through the use of Haskell’s type classes, we specify allowed components and combinations of molecules, ensuring only semantically meaningful structures are represented. Additionally, algebraic data types can be shown to be a group (or a monoid etc), meaning rotational symmetry of molecules can be described.
By using the Dietz representation’s general model for bonding systems, the ADT is inherently able to represent delocalised bonding in general, without extensive labelling of bond types. Resonance structures, organometallics, multi-center and electron deficient bonds may all be represented. Fractional, and zero bond orders may also be calculated.
Using atomic coordinate data, all molecules with stereoisomerism may be represented and distinguished, and dihedral bond angles, torsion angles, and bond lengths may be calculated. Such information adds minimally to the space complexity, but yet can be crucial when predicting physical propreties of drugs.
The Dietz represenation of constitution, combined with atomic coordinate data then solves the three identified limitations of SELFIES for going beyond organic chemistry[1]: the ability to represent delocalised bonding, explicit hydrogen representation, and the ability to represent complex stereochemistry.
The novel orbital data type creates a data structure to model shells, subshells, and orbitals, completely representing the electronic structure of an atom. Quantum information can then be used for machine learning applications, and for quantum chemical calculations in Molecular Orbital theory, as can be required with organometallics.
Apart from the obvious benefits of being able to represent molecules which cannot currently be represented using string-based representations (even with the creation of new and complicated grammars), our ADT confers a crucial benefit. By directly synthesizing molecules in accordance with the grammar defined by the ADT, machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation into another format.
5.1 Haskell
Haskell’s type classes denote which data types support specific operations, and in some cases can also specify laws that the operations should satisfy. By making data types an instance of a type class, we specify that the data type supports the defined operations, and ensure the operations adhere to the operational laws.
What is novel is the integration of probabilistic programming with Haskell into a concrete use case. Probabilistic programming languages themselves in haskell have been created before[54, 55, 42], laying the groundwork for usage with chemical structures. The existence of libraries such as LazyPPL[56] (demonstrated later in the paper), makes the ADT possible to use with Bayesian machine learning applications, such as drug design, which have been attempted before [57, 58, 54], and is an emerging area of research. The “lazy” (non strict, or call-by-need) property of Haskell, as discussed earlier, can provide efficiency, for instance when sampling from expensive functions or performing simulations only when necessary.
The record syntax in Haskell also enables access of only parts of the molecule needed for a task , e.g. in machine learning. Non-necessary data is not accessed, making machine learning models potentially less wasteful. This is difficult to do with non-local encoding of information, such as with SMILES or SELFIES strings, where a whole molecule, as a string, must be parsed.
The Maybe and Nothing types provide a type-safe way to handle operations that may fail, distinguishing between empty and zero values, a concept that in other languages might be handled by null references or exceptions, which often introduce program crashing errors. This distinction is later used in the orbital data type, to denote the absence of a SubShell.
Additionally, Haskell’s type-safe type system offers compile-time safety. This can completely eliminate a class of runtime errors commonly encountered in dynamically-typed languages like Python. Machine learning applications written with Python, or the PyTorch framework, must contend with the possibility of type-related errors during execution, even with type-annotation.
It should be noted that concepts such as Algebraic Data Types, or Functional Purity, are not unique to Haskell, therefore Haskell is not strictly necessary for the implementation.
Record syntax is now briefly discussed, before introducing the ADT itself.
Furthermore, we have dihedral and bond angles already. We have the meta characters as operations in haskell such as folds for a for loop and filter or takeWhile for a while loop.
5.2 Record Syntax
Record syntax allows for defining named fields within data types, facilitating direct access to individual parts of a data structure. The use of these names acts as a way to prevent confusion about whether one has a specific atom or coordinate in a data structure representing atoms, for instance. Here is a simple example demonstrating record syntax:
1data Atom = AtomS { atomID :: Integer,
2 symbol :: AtomicSymbol }
3 | AtomC { atomID :: Integer,
4 symbol :: AtomicSymbol,
5 coordinate :: Coordinate }
6
7data AtomicSymbol = C | O | H
8
9data Coordinate = Coordinate { x :: Double,
10 y :: Double,
11 z :: Double }
Listing 2: Updating the coordinate of an atom
In this example, Atom is a sum type with two constructors: AtomS and AtomC. The AtomS constructor represents an atom with a symbol and has two fields: atomID of type Integer and symbol of type AtomicSymbol. The AtomC constructor represents an atom with a symbol and a coordinate, and it has three fields: atomID of type Integer, symbol of type AtomicSymbol, and coordinate of type Coordinate. Record syntax automatically generates getter functions for each field.
Example 1: Updating the x-coordinate of an atom
1updateXCoordinate :: Double -> Atom -> Atom
2updateXCoordinate newX (AtomC atomID symbol coordinate) =
3 AtomC atomID symbol coordinate { x = newX }
4updateXCoordinate _ atom = atom
Listing 3: Modifying only the x-coordinate of an atom
The setCoordinate function updates the coordinate of an AtomC. It takes a new Coordinate and an Atom as arguments. If the atom is an AtomC, it uses the @ pattern to bind the entire AtomC value to the variable atomWC. This allows us to refer to the original AtomC value when updating its fields. We then use record syntax to create a new AtomC by updating only the coordinate field of atomWC with the newCoordinate, while keeping the atomID and symbol fields unchanged. If the atom is an AtomS, it returns the original atom unchanged since AtomS does not have a coordinate field.
Example 2: Extracting the x-coordinate of an atom
1getXCoordinate :: Atom -> Maybe Double
2getXCoordinate (AtomC _ _ coordinate) = Just (x coordinate)
3getXCoordinate _ = Nothing
Listing 4: Retrieving the x-coordinate of an atom
The getXCoordinate function retrieves the x-coordinate of an AtomC. It uses pattern matching to extract the coordinate field from the AtomC constructor. If the atom is an AtomC, it uses record syntax to access the x field of the coordinate and returns it wrapped in Just. If the atom is an AtomS, it returns Nothing since AtomS does not have a coordinate field.
Record syntax in Haskell provides a basic level of extensibility for data types, allowing the addition of new fields without requiring changes to existing functions that operate on the data type. This capability enables the evolution of data models with minimal impact on the codebase. This extensibility is particularly valuable in the context of molecular representation, as it allows for the incorporation of new scientific knowledge and the adaptation of the representation to specific domains or applications. However, for more sophisticated and flexible data manipulation, lenses offer a powerful approach. Lenses are first-class getter & setter functions that can be composed to access and modify deeply nested fields or manipulate multiple fields simultaneously, simplifying the process of working with complex or nested data structures by focusing directly on the target fields, bypassing the need for explicit navigation.
Record syntax in Haskell offers a robust mechanism for defining algebraic data types. They simplify the process of working with data by automatically generating getter and setter functions, thus reducing boilerplate code. This feature not only enhances code readability but also supports the seamless extension of data types without impacting existing functionalities.
5.3 The Molecule ADT
The following presents the heart of this paper, the molecule data type:
1
2data Molecule = Molecule {
3 atoms :: [Atom],
4 bonds :: M.Map (Integer, Integer) BondType
5 } deriving (Show, Read)
6
7data Atom = Atom {
8 atomID :: Integer,
9 atomicAttr :: ElementAttributes,
10 coordinate :: Coordinate,
11 shells :: Shells
12 } deriving (Show, Read)
13
14data BondType = Bond {delocNum :: Integer, atomIDs :: Maybe [Integer]}
15 deriving (Eq, Read, Show)
16
17data AtomicSymbol = O | H | N | C | B | Fe | F | Cl | S | Br | P | I
18 deriving (Eq, Read, Show)
19
20data ElementAttributes = ElementAttributes
21 { symbol :: AtomicSymbol,
22 atomicNumber :: Integer,
23 atomicWeight :: Double
24 } deriving (Eq, Read, Show)
25
26type EquilibriumBondLength = Angstrom
27newtype Angstrom = Angstrom Double deriving (Read, Show, Eq)
Listing 5: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent Bonds (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over).
The Molecule data type consists of two fields: atoms and bonds. The atoms field is a list of Atom values, representing the individual atoms that make up the molecule. Each Atom is identified by a unique atomID of type Integer, ensuring that each atom can be distinctly referenced within the molecule. The Integer type in Haskell represents arbitrary-precision integers, allowing for the representation and manipulation of integers without any predefined size limits, thus avoiding the issue of overflow that can occur with fixed-size types like Int. Unlike Int, which is limited to a fixed range determined by the computer’s architecture, Integer can handle values as large as the available memory permits.
The Atom data type contains several important attributes. The atomicAttr field of type ElementAttributes captures the fundamental properties of an atom, such as its atomic symbol (AtomicSymbol), atomic number (atomicNumber), and atomic weight (atomicWeight). These attributes provide essential information about the nature and characteristics of each atom in the molecule.
In addition to the atomic attributes, each Atom also has a coordinate field of type Coordinate, which represents the spatial position of the atom within the molecule. Each Coordinate has an x, y and z field. This information is crucial for understanding the geometric structure and spatial arrangement of the atoms, enabling the analysis of molecular shape, conformation, and reactivity.
Furthermore, the shells field of type Shells represents the electronic configuration of the atom, capturing the arrangement of electrons in the various atomic orbitals. This information is useful for understanding the chemical behavior and bonding properties of the atom, as well as its participation in chemical reactions. This will be discussed in the section on the Orbital type.
The bonds field of the Molecule data type is represented using a Map, with the keys being tuples of Integer values representing the atom IDs of the bonded atoms, and the values being of type BondType. This representation acts as an adjacency matrix and allows for the efficient lookup and manipulation of bonds within the molecule.
The BondType data type includes delocNum and atomIDs, which represent the delocalization number and the optional ring structure associated with the covalent bond. This allows for the representation of delocalized electrons and cyclic structures within the molecule. This allows for the accurate modeling of the dynamic distribution of electrons across multiple bonds within a molecule, capturing the essence of the Dietz representation.
By incorporating this information, the data type can effectively represent resonant structures and account for the delocalization of electrons, which is crucial for understanding the electronic properties and reactivity of molecules. The atomIDs field, of type Maybe [Integer], provides a means to represent the interconnected bonds within a cyclic structure. It is an optional field, indicated by the Maybe type constructor, allowing for the presence or absence of a list of atomIDs across which electrons are delocalised associated with the covalent bond. This list of integers enables for instance the representation of cyclic structures within the molecule. This design choice of using the delocNum and atomIDs fields in the Bond constructor offers a flexible framework capable of accurately depicting complex molecules, such as organometallic compounds like ferrocene. It provides a foundation for computing fractional bond orders, which is essential for molecules with resonant bonding characteristics, such as ozone ($O_{3}$).
ADTs are composed of simple, atomic data types combined using constructs such as products (tuples or records) and sums (variants or unions).
5.3.1 Example of a Molecule : Benzene
1
2benzene :: Molecule
3benzene = Molecule
4 { atoms = [ Atom i (elementAttributes e) (Coordinate x y z)
5 | (i, e, x, y) <- atomsData ],
6 bonds = M.fromList \$ getSymmetricBonds bondsData
7 }
8
9atomsData :: [(Int, Element, Double, Double, Double)]
10atomsData =
11 [ (1, C, 0.0, 1.3948, 0.0), (2, C, 1.2079, 0.6974, 0.0),
12 (3, C, 1.2079, -0.6974, 0.0), (4, C, 0.0, -1.3948, 0.0),
13 (5, C, -1.2079, -0.6974, 0.0), (6, C, -1.2079, 0.6974, 0.0),
14 (7, H, 0.0, 2.4732, 0.0), (8, H, 2.1431, 1.2366, 0.0),
15 (9, H, 2.1431, -1.2366, 0.0), (10, H, 0.0, -2.4732, 0.0),
16 (11, H, -2.1431, -1.2366, 0.0), (12, H, -2.1431, 1.2366, 0.0)
17 ]
18
19bondsData :: [((Int, Int), Bond)]
20bondsData =
21 -- 6 single CC bonds between neighboring carbon atoms
22 [ ((1,2), single), ((1,6), single), ((2,3), single),
23 ((3,4), single), ((4,5), single), ((5,6), single),
24
25 -- 2c2e bonds between carbon and hydrogen atoms
26 ((1,7), single), ((2,8), single), ((3,9), single),
27 ((4,10), single), ((5,11), single), ((6,12), single),
28
29 -- delocalised bonding in the benzene ring
30 ((1,2), aromatic), ((1,6), aromatic), ((2,3), aromatic),
31 ((3,4), aromatic), ((4,5), aromatic), ((5,6), aromatic) ]
32 where
33 single = Bond {delocNum = 2, atomIDs = Nothing}
34 aromatic = Bond {delocNum = 6, atomIDs = Just [1,2,3,4,5,6]}
Listing 6: A representation of the molecule benzene expressed in the Molecule data type, written as a .hs file, and using list comprehension to ensure readability, excluding the Shells currently.
The provided code demonstrates the representation of the benzene molecule using the proposed algebraic data type (ADT) system. Benzene is a classic example of an aromatic compound, consisting of a ring of six carbon atoms with alternating single and double bonds (or delocalised electrons around a ring), and six hydrogen atoms attached to each carbon.
In this representation, the benzene molecule is defined as a value of type Molecule. The atoms field contains a list of 12 Atom values, representing the six carbon atoms and six hydrogen atoms in the molecule. Each atom is specified with its unique atomID, atomicAttr (including the AtomicSymbol, atomicNumber, and atomicWeight), coordinate (3D spatial position), and shells (electronic configuration).
The carbon atoms (atomIDs 1-6) are positioned in a hexagonal arrangement, with coordinates that define the vertices of the hexagon. The hydrogen atoms (atomIDs 7-12) are positioned radially outward from each carbon atom, with coordinates that place them at a typical carbon-hydrogen bond distance.
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as Bond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6], explicitly specifying the cyclic structure formed by the carbon atoms.
The bonds between the carbon atoms and their respective hydrogen atoms are also represented as Bond values, but with a delocNum of 2, indicating a typical single bond between carbon and hydrogen. The atomIDs field for these bonds is therefore set to Nothing, as they are not part of the aromatic ring.
The code also includes the definitions of elementAttributes for hydrogen (H), nitrogen (N), and carbon (C). These elementAttributes specify the AtomicSymbol, atomicNumber, and atomicWeight for each element. For example, elementAttributes C = ElementAttributes C 6 12.011 represents the carbon element with an atomic number of $6$ and an atomic weight of $12.011$. These predefined elementAttributes are used in the atomicAttr field of each Atom in the benzene molecule, providing the necessary information about the elemental properties of the atoms. The representation could be extended to represent reactions or polymers, without requiring a completely new and separate syntax like BigSMILES. Instead, lists whose syntax and semantics represents reflects repetition of subunits.
5.4 The Orbital ADT
The representation captures quantum mechanical properties of atoms and can be made to adhere to the principles governing the arrangement of electrons in an atom. Each shell is represented as a distinct entity, characterized by its principal quantum number. This aligns with the quantum mechanical description of atomic structure, where electrons occupy discrete energy levels or shells. By organizing the electronic configuration into separate shells, the representation accurately reflects the hierarchical nature of atomic orbitals and the increasing energy and complexity of the electronic structure as the principal quantum number increases.
The use of a list to represent the shells implies that there could be an arbitrarily large finite number of shells (or infinite), which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
Within each shell, there are subshells ($s,p,d$, and $f$), which correspond to the different types of atomic orbitals. Any interface to this data structure must respect the maximum number of electrons that can occupy each subshell, with $s$ subshells containing up to two electrons, $p$ subshells containing up to six electrons, $d$ subshells containing up to ten electrons, and $f$ subshells containing up to fourteen electrons. However, since the current implementation does not utilize dependent types, the correctness of the number of electrons in each subshell cannot be guaranteed at compile time and is left as future work. The responsibility of ensuring the proper electronic configuration lies with the user of the ADT.
1-- Orbital types for each subshell
2data So = So deriving (Show, Eq, Read)
3data P = Px | Py | Pz deriving (Show, Eq, Read)
4data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read)
5data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
6
7-- The Orbital has a type and a number of electrons.
8data Orbital subshellType = Orbital {
9 orbitalType :: subshellType
10 , electronCount :: Int
11 , orientation :: Maybe Coordinate
12 } deriving (Show, Eq, Read)
13
14-- Each SubShell consists of a list of Orbitals.
15-- We ensure that all orbitals within a subshell are of the same type
16-- by using the subshellType parameter.
17newtype SubShell subshellType = SubShell {
18 orbitals :: [Orbital subshellType]
19 } deriving (Show, Eq, Read)
20
21-- Each Shell has a principal quantum number and consists of a list of SubShells.
22data Shell = Shell {
23 principalQuantumNumber :: Int
24 , sSubShell :: Maybe (SubShell So)
25 , pSubShell :: Maybe (SubShell P)
26 , dSubShell :: Maybe (SubShell D)
27 , fSubShell :: Maybe (SubShell F)
28 } deriving (Show, Eq, Read)
Listing 7: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells.
Each atomic orbital, contains the subshell type, electron count, and spatial orientation of the orbital. These properties can be used to construct the actual basis functions, but they are not the basis functions themselves. Furthermore, the representation allows for the specification of the spatial orientation of each orbital within a subshell through the use of the Coordinate type. This information is crucial for describing the shape and symmetry of atomic orbitals, particularly for $p$, $d$, and $f$ orbitals that have distinct orientations. However, it is important to note that the actual spatial representation of the orbitals would require the use of the appropriate mathematical functions, such as Slater-type orbitals (STOs), spherical harmonics, or Gaussian functions. Slater-type orbitals are commonly used in quantum chemistry as they provide a more accurate representation of the electron density near the nucleus compared to Gaussian functions. However, these mathematical functions are not explicitly included in this ADT system, which focuses on the symbolic representation of the electronic configuration.
By designing the appropriate interface functions that operate on the Shells type, one can ensure that rules such as the Aufbau principle, Hund’s rule, and the Pauli exclusion principle are properly enforced when manipulating the electronic configuration. These interface functions would act as a layer of abstraction, encapsulating the logic and rules governing the distribution of electrons within the atom. By providing a well-defined set of operations that adhere to these principles, the interface can maintain the integrity and consistency of the electronic configuration, thereby accurately representing the charge of the molecule. The interface functions would handle the creation, modification, and querying of the Shells data structure, ensuring that any changes made to the electronic configuration are valid and compliant with the quantum mechanical principles. This approach allows for a clear separation of concerns, where the Shells type focuses on the data representation, while the interface functions handle the logic and enforcement of the principles governing the electronic configuration.
The Aufbau principle can be enforced by ensuring that electrons are added to orbitals in the correct order of increasing energy. Hund’s rule can be implemented by maximizing the number of unpaired electrons in degenerate orbitals before pairing them. The Pauli exclusion principle can be enforced by limiting the maximum number of electrons in each orbital to two, with opposite spins. The use of type information can greatly assist in enforcing these principles and ensuring the validity of the electronic configuration. For example, the SubShell data type could be parameterized by the maximum number of orbitals allowed for each subshell type. This type-level information would prevent the creation of invalid subshells with more orbitals than permitted by the quantum mechanical rules. Similarly, the electronCount field of the Orbital data type can be constrained to a range of values (e.g., 0 to 2) using type-level naturals or refined types. This ensures that the Pauli exclusion principle is upheld, and no orbital contains more than two electrons.
However, it does not explicitly account for the spin of electrons or the coupling between orbital and spin angular momenta, which play a significant role in determining the detailed electronic structure of atoms, particularly for elements with partially filled subshells.
. Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbitals with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
As mentioned, in Haskell, algebraic data types (ADTs) are constructed using sum and product type constructors, analogous to addition and multiplication in algebra. A sum type is a choice between different types (denoted by the | symbol), akin to an or; for example, a Maybe type represents a value that may be Nothing (like zero in addition) or Just a where a is any type. This is reflective of a disjoint sum in algebra, where a value can be one thing or another, but not both simultaneously. This allows the grammar to distinguish between empty values and zero values. As will be discussed, the orientation field of the Orbital data type is wrapped in a Maybe type, which allows for the natural expression of whether an orbital has a specific orientation or not. For $s$ orbitals, the orientation is Nothing because $s$ orbitals are spherically symmetric and have no spatial orientation. In contrast, $p$ orbitals have specific orientations, and their orientation can be represented as Just Coordinate, where Coordinate is a type that can represent orientation in the $x,y,z$ axes. On the other hand, a product type combines several values into one composite type, similar to how factors multiply to a product; a tuple (a, b) is an instance of a product type containing both a and b.
5.4.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
1hydrogen :: Shells
2hydrogen =
3 [ Shell
4 { principalQuantumNumber = 1
5 , sSubShell = Just (SubShell [Orbital So 1 Nothing]) -- 1s^1
6 , pSubShell = Nothing
7 , dSubShell = Nothing
8 , fSubShell = Nothing
9 }
10 ]
11
12carbon :: Shells
13carbon =
14 [ Shell
15 { principalQuantumNumber = 1
16 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 1s^2
17 , pSubShell = Nothing
18 , dSubShell = Nothing
19 , fSubShell = Nothing
20 }
21 , Shell
22 { principalQuantumNumber = 2
23 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 2s^2
24 , pSubShell = Just (SubShell [Orbital Px 1 (Just (Coordinate 1 0 0)),
25 Orbital Py 1 (Just (Coordinate 0 1 0))])
26 -- 2p^2
27 , dSubShell = Nothing
28 , fSubShell = Nothing
29 }
30 ]
31
32oxygen :: Shells
33oxygen =
34 [ Shell
35 { principalQuantumNumber = 1
36 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 1s^2
37 , pSubShell = Nothing
38 , dSubShell = Nothing
39 , fSubShell = Nothing
40 }
41 , Shell
42 { principalQuantumNumber = 2
43 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 2s^2
44 , pSubShell = Just (SubShell [Orbital Px 2 (Just (Coordinate 1 0 0)),
45 Orbital Py 1 (Just (Coordinate 0 1 0)),
46 Orbital Pz 1 (Just (Coordinate 0 0 1))])
47 -- 2p^4
48 , dSubShell = Nothing
49 , fSubShell = Nothing
50 }
51 ]
Listing 8: Examples of electronic configurations for hydrogen, carbon, and oxygen using the Orbital Algebraic Data Type representation in Haskell.
For the hydrogen atom, the electronic configuration is represented by a single shell with a principal quantum number of 1. The shell contains only an $s$ subshell (sSubShell) with one orbital occupied by a single electron, denoted as Orbital So 1 Nothing. The Nothing value for the orientation indicates that s orbitals have a spherical symmetry and do not have a specific orientation. This represents the ground state configuration of hydrogen, which is 1s1.
The carbon atom’s electronic configuration is represented by two shells. The first shell, with a principal quantum number of 1, consists of an $s$ subshell with two electrons (Orbital So 2 Nothing), corresponding to the 1s2 configuration. The second shell, with a principal quantum number of $2$, comprises an $s$ subshell with two electrons (Orbital So 2 Nothing) and a $p$ subshell (pSubShell) with two orbitals, each containing one electron. The p orbitals are specified as Orbital Px 1 (Just (Coordinate 1 0 0)) and Orbital Py 1 (Just (Coordinate 0 1 0)), representing the 2px and 2py orbitals, respectively. The Coordinate values (1, 0, 0) and (0, 1, 0) indicate the orientation of the p orbitals along the x and y axes, respectively. This configuration, 1s22s22p2, describes the ground state of carbon.
Similarly, the oxygen atom’s electronic configuration is represented by two shells. The first shell is identical to that of carbon, with an $s$ subshell containing two electrons. The second shell consists of an $s$ subshell with two electrons and a $p$ subshell with four electrons distributed across three orbitals: Orbital Px 2 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0)), and Orbital Pz 1 (Just (Coordinate 0 0 1)). The Coordinate values (1, 0, 0), (0, 1, 0), and (0, 0, 1) represent the orientation of the $p$ orbitals along the $x$, $y$, and $z$ axes, respectively. This describes the ground state configuration of oxygen, which is 1s22s22p4. The oxygen atom’s electronic configuration differs from carbon in the occupancy of the p subshell, where oxygen has four electrons distributed across three p orbitals, while carbon has two electrons in two p orbitals.
The inclusion of the Coordinate type in the representation allows for the specification of the spatial orientation of the orbitals. This is particularly relevant for $p$, $d$, and $f$ orbitals, which have distinct orientations in three-dimensional space. By providing the coordinates, the representation captures the directional nature of these orbitals, enabling a more accurate description of the electronic structure and the spatial distribution of electrons in the atom.
6 Experiments and extensions
6.1 Reaction Networks
The proposed algebraic data type for representing molecules can be extended to model chemical reactions and reaction networks. By defining a Reaction data type that captures the reactants, products, and conditions under which a reaction occurs, we can create a powerful tool for studying and simulating chemical processes. A chemical reaction network comprises a set of reactants, a set of products, and a set of reactions.
Reactions are commutative but not associative and given further time could be incorporated into the type. This would be called an abelian or commutative magma.
Here’s an example of how the Reaction data type can be defined: 1data Reaction = Reaction { 2 reactants :: [(Double, Molecule)] 3 , products :: [(Double, Molecule)] 4 , conditions :: [Condition] 5 , reactionRate :: Double 6} 7 8data Condition = TempCondition {temperature :: Double, times :: Times} 9 | PressureCondition {pressure :: Double, times :: Times} 10 11newtype Times = Times (StartTime Double, EndTime Double) Listing 9: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
In this representation, a Reaction consists of a list of reactants and products, each represented as a tuple of a stoichiometric coefficient (of type Double) and an Atom. The conditions field is a list of Conditions, which can be either temperature or pressure conditions, specified along with the time range over which they apply.
Using this Reaction data type, we can define a ChemicalReactionNetwork as a collection of reactions:
1newtype ChemicalReactionNetwork = ChemicalReactionNetwork {
2 reactions :: [Reaction]
3}
Listing 10: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
This allows for the modeling of complex chemical systems and the study of their behavior under different conditions. For example, consider the reaction for the formation of water from hydrogen and oxygen:
1exampleReaction :: Reaction
2exampleReaction = Reaction {
3 reactants = [(2.0, h2), (1.0, o2)],
4 products = [(2.0, h2o)],
5 conditions = [ TempCondition 500.0 (Times (0.0, 10.0)),
6 PressureCondition 1.0 (Times (0.0, 10.0)) ],
7 reactionRate = 0.1
8}
Listing 11: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate.
This reaction represents the formation of two water molecules from two hydrogen molecules and one oxygen molecule, under the conditions of a temperature of $500K$ and a pressure of $1$ atm, over a time range of 00 to $10$ seconds.
6.2 Group Theory
Moreover, the ADT can be instantiated as a group to incorporate group theory concepts, which play a crucial role in understanding molecular symmetry and its impact on chemical properties. Exploiting these group symmetries can aid neural networks by the use of group equivariance and group invariance [2]. Invariances can be used also with atomic IDs meaning that the same atom represented with different ID’s can be quotiented and so treated as the same, allowing less data to be needed in order to train a neural network, for instance. In particular, molecules can be made instances of rotation groups, such as the cyclic group $C3$, which represents the rotational symmetry of molecules with threefold rotational axes.
Here’s an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass: 1class Group g where 2 mul :: g -> g -> g -- Group operation: combining two operations 3 inv :: g -> g -- Inverse of an operation 4 e :: g -> g -- Identity operation 5 6data MolRot = MolRot Molecule Coordinate Double deriving (Show) 7 8instance Group MolRot where 9 mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) = 10 combineRotations mol1 axis1 angle1 axis2 angle2 11 12 inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle) 13 14 e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0 Listing 12: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate. In this representation, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e).
By making MolRot an instance of Group, we can combine rotations using the mul function, find the inverse of a rotation using inv, and obtain the identity rotation using e. This allows for the study of molecular symmetry and its implications on chemical properties and reactivity.
However, it’s important to note that proving that a data type satisfies all the group axioms (closure, associativity, identity, and inverse) would require the use of dependent types, which is beyond the scope of this representation. Nonetheless, by defining the necessary operations and providing sensible implementations, we can still leverage the power of typeclasses in representing these properties.
6.3 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the paper, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
Molecules in the ADT can be serialised and stored as .hs files, which may be used for storage, and for transmission of molecular data, perhaps via large chemical databases, through straightforward use of the Show and Read instances in Haskell, as shown in Listing 14.
Additionally, as .hs files will be interpreted with Haskell, molecules can be straightforwardly compressed as Listing 13 demonstrates by using foldr to compress the bondsData list of the benzene molecule. Although this may raise the issue of “canonicalisation” for storage of molecules, it should be noted the bondsData in Listing 13 and Listing 6 will be stored the same way inside of a Haskell program, as the corresponding data is the same in each case. This could be beneficial in the representation of large molecules, as only a small number of bonds or atoms need be explicitly described to encode the molecule for storage.
1
2bondsData :: [((Int, Int), Bond)]
[…]
6 where
7 single = Bond {delocNum = 2, atomIDs = Nothing}
8 aromatic = Bond {delocNum = 6, atomIDs = Just [1,2,3,4,5,6]}
Listing 13: Using foldr to compress the bondsData of the molecule benzene in the ADT representation. As Benzene is symmetric, the bonding system can be specified by only describing three bonding systems, and using those to construct the bondsData list.
1-- Writing the molecule to a file
2writeMoleculeToFile :: FilePath -> Molecule -> IO ()
3writeMoleculeToFile filePath molecule = writeFile filePath (show molecule)
4
5-- Reading the molecule from a file
6readMoleculeFromFile :: FilePath -> IO Molecule
7readMoleculeFromFile filePath = do
8 contents <- readFile filePath
9 return (read contents)
10
11-- Example usage
12main :: IO ()
13main = do
14 -- Write the methane molecule to a file
15 writeMoleculeToFile "methane.hs" methane
16
17 -- Read the methane molecule from the file
18 molecule <- readMoleculeFromFile "methane.hs"
19
20 -- Print the molecule read from the file
21 print molecule
Listing 14: Haskell code demonstrating the serialisation and writing to disk of a Molecule data type and the reading of a molecule stored in an .hs file, through use of the show and read functions.
An example of the usage is shown.
This code demonstrates the ease of serializing and deserializing molecules to and from files in Haskell. By leveraging the Show and Read instances defined for the Molecule data type and its related types, molecules can be seamlessly converted to and from their string representations, allowing them to be stored in files and retrieved later.
This functionality provides a convenient way to save and load molecular data, and is particularly useful when working with large molecules or when molecules are required across different program runs. Although storage and transmission may take place via serialisation into a human and machine-readable string of text, it is important to distinguish between the serialized format of the data and the underlying data type itself. While the ADT can be serialized into a text string, it is crucial to recognize that the ADT itself is not merely a string.
1-- Observed molecule (e.g., H2O)
2observedMolecule :: Molecule
3observedMolecule = h2o
4
5-- A generative model for a molecule
6moleculeModel :: Meas Molecule
7moleculeModel = do
8 let numAtoms = 3
9 atoms <- replicateM numAtoms \$ do
10 symbol <- sample \$ uniformD [C, N, O, H]
11 coord <- sampleCoordinate
12 let attr = elementAttributes symbol
13 let shellsVar = elementShells symbol
14 return \$ Atom { atomID = 0,
15 atomicAttr = attr,
16 coordinate = coord,
17 shells = shellsVar }
18
19 let molecule = Molecule { atoms = atoms, bonds = randomBonds 3 }
20
21 -- Score the model based on the distance to the observed molecule
22 let distance = distMolecule molecule observedMolecule
23 score \$ normalPdf 0 1 distance
24
25 return molecule
26
27-- Sample a coordinate from a normal distribution
28sampleCoordinate :: Meas Coordinate
29sampleCoordinate = do
30 x <- sample \$ normal 0 1
31 y <- sample \$ normal 0 1
32 z <- sample \$ normal 0 1
33 return \$ Coordinate x y z
34
35-- Calculate the Euclidean distance between two coordinates
36euclideanDistance :: Coordinate -> Coordinate -> Double
37euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) =
38 sqrt \$ (x1 - x2)^2 + (y1 - y2)^2 + (z1 - z2)^2
39
40-- Calculate the distance between two molecules
41distMolecule :: Molecule -> Molecule -> Double
42distMolecule mol1 mol2 =
43 let coords1 = map coordinate (atoms mol1)
44 coords2 = map coordinate (atoms mol2)
45 distances = [euclideanDistance c1 c2 | c1 <- coords1, c2 <- coords2]
46 in minimum distances
47
48-- Example usage
49main :: IO ()
50main = do
51 samples <- mh 0.1 moleculeModel
52 let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples
53 putStrLn \$ "Sampled-molecules:" ++ show molecules
54 putStrLn \$ "Weights:" ++ show weights
Listing 15: In a probabilistic programming context, trace Metropolis-Hastings (M-H) operates by sampling entire “traces” of latent variables (i.e., all random choices) from a model [42]. Each step proposes a small perturbation to the current trace, controlled by a jitter parameter, often the standard deviation of the proposal distribution. This parameter must be tuned carefully to ensure efficient exploration of the posterior while maintaining a reasonable acceptance rate. Typically, an initial burn-in period is applied, during which early samples (which have not yet converged) are discarded. This burn-in helps remove the influence of the initial point on the chain, leading to samples that more accurately reflect the stationary distribution of the target posterior. Here the jitter is set to 0.1 and the number of burnin samples is 1000.
6.4 A simple example of probablistic programming
The code block above can be explained as follows:
  • We define observedMolecule as the molecule we want to condition our model on. In this example, we use the predefined h2o molecule, but you can replace it with any other molecule of interest.
  • We define a generative model moleculeModel using the Meas monad. This model generates a Molecule by first sampling the number of atoms numAtoms from a discrete uniform distribution using uniformD. In this case, we set numAtoms to a fixed value of 3.
  • We then use replicateM to generate numAtoms number of Atoms. For each Atom, we sample the symbol uniformly from the list [C, N, O, H] using uniformD, and sample the coordinate using the sampleCoordinate function.
  • We use record syntax to create each Atom. The atomID is set to a default value of 0 (which can be updated later), the atomicAttr is set using the elementAttributes function based on the sampled symbol, and the shells are set using the elementShells function, also based on the symbol. The charge is assumed to be neutral at this point.
  • After generating the list of atoms, we create a Molecule using record syntax. The atoms field is set to the generated list of atoms, and the bonds field is set using the randomBonds function, which randomly initializes a 3x3 adjacency matrix.
  • The sampleCoordinate function is a helper function that samples a Coordinate by sampling the $x$, $y$, and $z$ values from normal distributions using normal.
  • The randomBonds function generates a random 3x3 adjacency matrix representing the bonds between atoms. Currently it is not implemented for brevity, but the implementation would be straightforward.
  • In the moleculeModel, after generating the molecule, we calculate the distance between the generated molecule and the observed molecule using the distMolecule function.
  • We use the score function to compute the likelihood of the distance given a normal distribution with mean 00 and standard deviation $1$. This scoring step conditions the model on the observed molecule, giving higher weights to generated molecules that are closer to the observed molecule.
  • The distMolecule function takes two molecules as input and calculates the distance between them. It extracts the coordinates of the atoms in each molecule and computes the Euclidean distance between all pairs of coordinates using euclideanDistance. The minimum distance among all pairs is returned as the distance between the molecules.
  • In the main function, we generate samples using Metropolis-Hastings and drop the first one thousand samples as a burn in. We then print the remaining molecules and weights.
By having an observed molecule and using the distMolecule function, we can condition our generative model on any molecule of interest. The scoring mechanism computes the distance between the generated molecule and the observed molecule, guiding the sampling process towards molecules that are more similar to the observed one.
The Prob monad is used for representing probability distributions, while the Meas monad is used for representing unnormalized probability distributions (measures). The sample function is used to sample from a probability distribution (Prob), and the score function is used to score the model based on some condition.
By conditioning the generative model on the observed molecule, we restrict the samples to only those molecules that are similar to the specified molecule. The Metropolis-Hastings algorithm then explores the space of possible molecules and generates samples according to their weights (probabilities) based on the conditioned model.
7 Discussion
7.1 Limitations
The storage efficiency of the should be further explored. The ADT syntax could result in large file sizes for molecules stored as .hs files, however, the use of list comprehension, and the fact that molfiles and the Molecule data type belong to the same complexity class may mean that the ADT representaion is fact smaller than some mol files, for the same degree of expressiveness. By storing a molecule as function, it is possible to significantly reduce the size of the molecule.
7.2 Future work
The ADT so far presented has the ability to be improved upon, extended, and utilised in empirical experiments in order to verify and test the representation’s utility across a number of domains in cheminformatics.
7.2.1 Augmentation of the ADT
To improve the representational ability of the ADT, it should be augmented to enable representation of tautomerism, ionic bonding, polymerism, and quantum information such as spin, none of which are represented in the current ADT.
Tautomers are in medications from sildenafil (Viagra) [59], to Warfarin, to Remdesivir (used for Covid 19), Tetracyclines and antibiotics. Representation of tautomers could allow exploration of a greater chemical space, and facilitate tasks such as prediction of ligand-protein binding, necessary for many drug design tasks[60, 61].
Polymers are important in many areas of chemistry and materials science. By using lists and recursive data structures, the representation could be augmented to handle polymeric molecules. This extension would enable the modeling and analysis of complex macromolecular systems, opening up new possibilities for polymer design and optimization.
The inclusion of ionic bonding to the ADT would be another significant enhancement. Ionic interactions are critical for the stability and functionality of nucleic acids, proteins, and membranes, as well as for processes like catalysis and ion transport [62, 63, 64]. Augmenting the ADT to represent ions and their bonding explicitly would enable the modeling of charged species and their interactions. This could also facilitate the study of electrochemical properties and enable the design of materials such as ionic liquids and solid electrolytes, which are essential for applications ranging from energy storage to green chemistry [65].
The extension of the ADT to include quantum mechanical properties such as spin and angular momentum represents a significant opportunity. Quantum mechanical features are crucial for accurately modeling the electronic structure and predicting molecular reactivity, as they account for phenomena like electron spin coupling and charge distribution, which play a key role in determining molecular properties [12, 66]. By incorporating these properties into the ADT, it could better capture the behavior of molecules in complex environments, including those encountered in catalysis and photochemistry. Such advancements would also enable the ADT to align more closely with the requirements of quantum chemistry simulations, which often rely on accurate spin and orbital information for predicting reaction mechanisms and designing new materials [11, 1].
7.2.2 Development of powerful, user-friendly library
To make the most of the ADT, a powerful, user-friendly library of functions should to be created. Functions should include the ability to convert between the ADT and other common representaions, which would enable use of the ADT with existing potentially vast chemical databases.
Canonicalisation is also a concern for the ADT, as molecules may have different indicies, and there may be differing atomic coordinates. The permuations of a set of indicies, would however form a group, and the group properties of the ADT can be leveraged for this purpose.
7.2.3 Futher enhancing the ADT
To further strengthen the type safety and invariance guarantees of the representation, exploring dependently typed programming languages, such as Agda, could be highly beneficial. Dependent types allow for the encoding of more sophisticated invariants and constraints directly into the type system, ensuring that only valid and consistent molecular structures are expressible. This increased level of type safety can help catch potential errors at compile-time and provide stronger guarantees about the correctness of the representation.
Further work could explore several avenues to enhance the proposed algebraic data type (ADT) representation and its applications. Integrating Haskell’s Lens library [67] could simplify the manipulation of complex molecular structures by providing a composable, modular and type-safe way to access and modify nested data fields.
To enhance the modularity and composability of the ADT representation, exploring techniques like modular syntax trees and the “Data Types a la Carte” approach [68] could prove highly valuable. By decomposing the molecular representation into smaller, reusable components, we can create a more flexible and adaptable framework. This modular design would allow for the easy integration of new features, the creation of domain-specific languages for molecular manipulation, and the development of reusable libraries and tools for cheminformatics.
By pursuing these avenues of future work, we can further strengthen the capabilities and thus impact the community, encouraging them to use the ADT-based molecular representation, hopefully accelerating the discovery of novel and useful molecules. Implementing backpropagation using differentiable programming [69] would enable using gradient based inference algorithms.
Integrating the ADT with quantum sampling techniques could lead to breakthroughs in predicting molecular properties by efficiently simulating complex quantum systems. However, realizing this goal would require close collaboration between experts in quantum computing, chemical informatics, and molecular modeling to develop efficient and correct quantum algorithms.
7.2.4 Experimental validation of the ADT
At present, the efficiency and efficacy of the ADT in cheminformatics tasks are conjectures. To validate the utility of the ADT as a molecular representation, empirical work is needed to benchmark its performance against representations such as SMILES and SELFIES, on tasks like molecular property prediction (e.g. calculation of logP and logS), virtual screening, and de novo drug design.
Another avenue for future work is utilising the proposed representation with geometric deep learning techniques to exploit symmetries and invariances in molecular structures [2]. By leveraging mathematical properties of ADTs, such as group invariances and equivariances and their verification in Haskell through typeclasses, one could develop neural network architectures that are specifically tailored to the unique characteristics of those molecules. For instance, the choice of index for each atomID used in the representation should not influence a model, given the same molecule but with a different ordering of atomic IDs.
An additional avenue is using equational reasoning to investiage algebraic approaches in molecular modelling [70].
Neural networks can be designed to be invariant to permutations of atomic indexes, ensuring that molecularly equivalent structures are treated identically regardless of labeling of indices or bonds. Similarly, molecular features such as the atomSymbol, are both translation and rotation invariant, while others may only be translation invariant. Incorporating these symmetries directly into the neural network architecture could significantly reduce the amount of training data required and improve the generalization capability of the models, but this would have to be demonstrated.
8 Conclusion
In this work, we presented a novel approach to molecular representation using algebraic data types (ADTs) embedded in the purely functional, type-safe, programming language Haskell.
We discussed contemporary representations, focusing primarily on the string based representations SMILES and SELFIES, and their short-comings for the use in machine learning.
By implementing the ADT using the Dietz representation for constitution of molecules, and atomic coordinate data for 3D information, the ADT greatly expands the chemical space which can be represented, and hence explored in cheminformatics tasks when compared to existing representations. In addition, the ADT allows the optional inclusion of quantum mechanical information, such as orbitals and electronic configurations, providing a level of detail unmatched in current representations.
As a general model for molecular representation, the ADT can represent complex, multi-centre, electron-deficient and delocalised bonding, as well as advanced stereoechemical features, and “zero order” bonds such as ionic and hydrogen bonds, without requiring extensive labeling for new bond types or oxidation states.
Through Haskell’s type classes and functional programming paradigm, we provide a rich and robust semantic context for molecular structure, and an efficient framework for computational tasks. Molecular data adheres to the constraints of the molecular representation, preventing nonsensical molecules from being formed, and going beyond the limitations of strings as a representation.
To demonstrate the ADT and how it can be extended, we implemented reaction networks and explore group based properties of molecular symmetry. We also showed integration of the ADT with probabilistic programming, showing that the ADT can be appropriately used for Bayesian inference, and that machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation to other formats.
The ADT presented was not intended to be a definitive representation, but a valuable representational concept for cheminformaticians. Nevertheless the ADT offers a powerful and flexible alternative to existing string-based representations like SMILES and SELFIES, for the representation of, manipulation of, and reasoning about molecular structures in cheminformatics.
9 Declarations
9.1 Availability of data and materials
All code used to implement the Algebraic Data Type and the experiments in this paper is available via github, at https://github.com/oliverjgoldstein/chemalgprog
9.2 Competing interests