### v3

# Representation of Molecules via Algebraic Data Types : Advancing Beyond SMILES & SELFIES

Abstract
We introduce a novel molecular representation through Algebraic Data Types (ADTs) - composite data structures formed through the combination of simpler types that obey algebraic laws. By explicitly considering how the datatype of a representation constrains the operations which may be performed, we ensure meaningful inference can be performed over generative models (programs with sample and score operations). This stands in contrast to string-based representations where string-type operations may only indirectly correspond to chemical and physical molecular properties, and at worst produce nonsensical output.
The ADT presented implements the Dietz representation for molecular constitution via multigraphs and bonding systems, and uses atomic coordinate data to represent 3D information and stereochemical features. This creates a general digital molecular representation which surpasses the limitations of the string-based representations and the 2D-graph based models on which they are based. In addition, we present novel support for quantum information through representation of shells, subshells, and orbitals, greatly expanding the representational scope beyond current approaches, for instance in Molecular Orbital theory.
The framework’s capabilities are demonstrated through key applications: Bayesian probabilistic programming is demonstrated through integration with LazyPPL, a lazy probabilistic programming library; molecules are made instances of a group under rotation, necessary for geometric learning techniques which exploit the invariance of molecular properties under different representations; and the framework’s flexibility is demonstrated through an extension to model chemical reactions.
After critiquing previous representations, we provide an open-source solution in Haskell – a type-safe, purely functional programming language.
keywords:
Molecular Representation, Cheminformatics, Algebraic Data Types
1 Introduction
The choice of molecular representation directly affects cheminformatics tasks such as novel molecule discovery and molecular property prediction [1]. More expressive representation enables the exploration of larger regions of chemical space, both as input to and output from machine learning models. However, representational ability alone does not determine the effectiveness of a molecular representation in ML applications. Computational efficiency, model convergence, and predictive accuracy are also important factors, along with the ease—or even possibility of integrating representations with techniques such as geometric deep learning [2] or Bayesian machine learning [3]. In cheminformatics, we believe that when it comes to molecular representation, “the medium is the message.”
Despite decades of research, string-based representations such as SMILES [4] and SELFIES [5] remain the most common representations for ML tasks [6, 7], despite well documented short-comings in representational ability and with syntaxes themselves. For instance, standard SMILES strings cannot represent complex bonding systems, resonance structures, or delocalized bonding, and do not include any 3D information, with only a limited ability to describe stereochemical features. In addition, not all SMILES strings correspond to valid molecules, meaning that the output of ML models is, in many cases, gibberish [5], and small changes in SMILES strings can result in large differences in molecular structure, making property prediction difficult. Attempts to build upon strings have led to a profusion of new grammars and syntaxes to represent molecular and chemical phenomena, a problem in itself.
While proposing future work to address the limitations of string-based representations such as SELFIES, Krenn et al (2022) [5] suggest developing a “molecular programming language”. Reasoning that strings can encode Turing-complete programming languages (e.g., Python source code), they suggest SELFIES could itself become a programming language, extended with a syntax that allows for recursion (for/while loops) and logical statements. Although this reasoning appears to conflate strings as a data structure with strings as a sequence of ASCII characters, we ask the question: if molecules can be represented as data within programming languages, we ask, why not simply use one?
In this article, we critique the string-based representations SMILES and SELFIES, along with their extensions, evaluating their structure, limitations, and suggested improvements. We assess molecular fingerprints, focusing on their role in similarity searching and cheminformatics workflows, and examine their applicability and challenges in machine learning contexts. Additionally, we evaluate how molecules are represented within programming environments, specifically considering RDKit as a representation, prior attempts to formalize molecular representations using Inductive Logic Programming (ILP), and previous work integrating functional programming and cheminformatics. We explore the mathematical primitives underlying these representations—vectors, graphs, and strings, including as a data type, and their implications for computational workflows. We consider existing representations in terms of their applicability to Bayesian Machine learning and Geometric Deep Learning, as well as reaction modeling. Finally, we conclude by presenting the ADT framework, implemented in the functional and strongly typed language, Haskell.
The ADT that we introduce uses the Dietz representation for constitution of molecules[8], using multigraphs of atomic valence information. It contains optional features such as 3D atomic coordinate data to provide functions such as conformation energy prediction and configuration of molecules, as well as quantum information in the form of orbitals.
The ADT is a general model for molecular representation, surpassing the limitations of 2D-graph based representations. It can represent complex bonding systems, such as the delocalised and multi-centre bonding within organometallics, and through atomic coordinates it can represent geometry beyond a limited range of labeled stereochemical features.
We demonstrate the flexibility of the ADT by showing an extension to model chemical reactions, and beyond representation, we demonstrate the ADTs suitability for Bayesian Machine Learning techniques with a probabilistic model written with Lazy PPL, a probabilistic programming library. We also demonstrate that ADTs can be made groups under permutation of atomic index, a necessary requirement for integration with Geometric Deep Leaning Techniques.
We also consider the impact the ADT has beyond computation, and consider the impact of the ADT as a representation for storage (and retrieval) of molecular information, and for transmission of molecular information between cheminformaticians.
The length of this article is reflective of the cross-disciplinary nature of the work, and of the requirement to explain concepts from functional programming and type-theory that may not be necessary for a more familiar audience. Although other work has linked functional programming and cheminformatics before [9, 10], this work was primarily focused on implementing existing cheminformatics workflows within functional programming languages, and the benefits of immutable state and type-saftey. Instead, we provide a molecular representation, and the surrounding context for the importance of considering data types when chosing a representaion.
While not intended as a definitive implementation of molecular representation, the ADT demonstrates the potential of algebraic approaches to reshape computational chemistry and invites the cheminformatics community to explore its capabilities and extend its utility.
1.1 Forms of Molecular Representation
In this article, we distinguish three different usages of the term “molecular representation”:
  1. 1.
    Storage representations refer to the format for storage of molecular data.
  2. 2.
    Transmission representations refer to the way molecules are sent between cheminformaticians, and depend on the medium of communication used.
  3. 3.
    Computational representations refer to how molecules are represented inside of a programming language or program, specifically, the data type or types which are used to represent the molecule.
For instance, a molecule may be stored in a database as a molfile, be processed with cheminformatics library such as RDKit[11] into another internal representation, and be then transformed into a molecular graph, or a fingerprint, for processing in a machine learning model [12, 13].
String based representations, such as the ubiquitous SMILES[4], may be used for all three. They are available in databases, may be downloaded and transmitted as strings, and also used as direct input for machine learning models[12, 13, 1], which as we later discuss, we regard as a mistake.
We regard the distinction between underlying data, and the data type to be crucial in cheminformatics, as it is the data type which constrains the operations which may be performed on data, reflects the structure and semantics of the data, and affects the efficiency and utility of computational tasks. Failure to make this distinction can lead ambiguity, inconsistency, and potential errors in interpretation and manipulation for meaning, as we later discuss in detail.
1.2 Data Types
Before continuing with the article, we need to define some concepts used throughout:
Definition 1 (Data Type).
A data type classifies and structures the values a program can manipulate, describing how they are constructed, stored, and what operations are valid on them. In many programming languages, a data type encapsulates both the possible shapes (or “constructors”) of the data and the rules for operating on those shapes.
In Haskell, a data type is introduced with the data keyword and defines one or more constructors that specify exactly how to build valid values of that type. This not only ensures consistency (i.e., that every value has a well-defined form) but also guides how functions can safely pattern-match and handle each variant of the type.
Definition 2 (Typeclass).
A typeclass (in languages such as Haskell) is a mechanism for defining a set of operations or methods that one or more data types can implement. By declaring a type as an instance of a typeclass, one provides implementations of the typeclass’s methods for that particular type.
Typeclasses enable polymorphic functions to operate uniformly over all types that satisfy the typeclass interface. An example might be equality, such that one can use == on that data type.
Definition 3 (Type Saftey).
A guarantee that operations in a language will only be performed on compatible types, preventing type errors at compile time (or runtime).
Languages which are not type-safe may utilise typeannotation, where types for variables and functions are explicitly specifying to guide the compiler and improve code clarity, or type inference , where a compiler automatically deduces the types of expressions without explicit type annotations.
Type annotations are considered practically weaker because they serve mainly as optional documentation or hints rather than as fully enforced contracts. In many languages (e.g., Python), type annotations are not strictly checked at compile time or runtime. meaning that:
  • They can be incomplete or even outdated, lacking the automatic consistency guarantees that come with full type inference.
  • They rely on external tools (like linters or type checkers) rather than being integrated into the language’s core execution semantics.
  • The annotations do not prevent all type errors at runtime, as the language may still perform implicit conversions or delay error detection until execution.
Thus, while they improve code readability and can catch some errors, type annotations do not enforce type correctness as rigorously as a fully statically and automatically inferred type system.
1.3 Molecular Graphs
Many common molecular representations are built around molecular graphs and, as such, inherit strengths and limitations of molecular graphs for molecular representation [12]. String based representations such as SMILES and SELFIES use a number of rules to decode ASCII strings which correspond to molecular graphs, and MDF molfiles have “blocks” which correspond to sets of atoms and sets of bonds.
Definition 4 (Graph).
Formally, in set-theoretic notation, a graph, $G=(V,E)$ is a pair of sets $V$ and $E$, where $V$ is a set of vertices and $E$ is a set of pairs of vertices {u,v}, called edges.
A molecular graph is an instance of a graph in which the vertices are atoms, and the edges are bonds between atoms. Molecular graphs are connected, undirected, and labeled, where vertices may be labeled with information such as chemical element, and charge; and edges may be labeled with bond type (e.g. single, double or aromatic.), bond order, or 3D information (e.g. R/S isometry) [8, 12].
One benefit of representing molecules computationally as graphs is that insights from chemical graph theory can be directly leveraged by machine learning models. However, not all molecules can be represented by graph models, including molecules with complex bonding, organometallics, resonant structures and tautomerism.
Edges in graphs are between two vertices only, so molecular graphs have a limited ability to represent delocalised and multi-centre bonding, including ionic bonds and organometalics [12, 13]. Although it is possible to label edges with, e.g. rational bond orders, the contribution of individual electrons to multiple covalent bonds would not be described by such a graph. Attempting to express delocalised bonding by labeling edges of a molecular graph with bond types would require a sufficient number of predefined bond types, which would need to be extended for each new bonding relation not described.
Evidence of this approach extends back decades, as Dietz said ([8]):
“Enhancing the expressiveness by including a new bond type for every exceptional case is certainly not a very elegant solution.”
Using a single structure, graph models cannot represent tautomers, so instead require a work-around such as using multiple structures to represent the different forms of molecules.
Hypergraph and multi-graph models have been proposed to solve some issues with graph models [14], but they have seen limited use [12].
1.4 Molecular Hypergraphs
A hypergraph, is a graph where edges may connect any number of vertices. In a molecular hypergraph, edges with two vertices represent localised bonding, and edges with more than two verticies represent delocalised bonding, and are labelled with the number of electrons shared between the corresponding atoms.
From Dietz:
“A hyperedge containing more than two atoms gives us no information about the binary neighborhood relationships between them. That means we have no information at our disposal conceming the way in which the electrons are delocalized over these atoms” [8].
Hypergraphs would indicate an electronic “soup” with out knowledge of the manner electrons are delocalised. Hypergraphs remain uncommon.
1.5 Multigraphs
Multigraphs are not widely used as a standard computational representation in cheminformatics or chemical graph theory. Existing tools and frameworks generally rely on labeled graphs or weighted graphs to encode bond multiplicities and types.
Theoretical discussions of multigraphs in chemical graph theory arise in the context of molecules with multiple bond types or delocalized bonding, but computational implementations of multigraph representations are not widespread.
Definition 5 (Multi-Set).
The set-theoretic definition of a multiset (or bag) is a generalization of a set where elements are allowed to appear more than once. Formally, a multiset $M$ over a universe $U$ is defined as a function:
$M:U\to\mathbb{N}$
where $M(x)$ denotes the multiplicity of element $x$, i.e., the number of times $x$ appears in the multiset. A finite multiset can be represented as:
$M=\{a_{1}^{m_{1}},a_{2}^{m_{2}},\dots,a_{n}^{m_{n}}\}$
where each element $a_{i}$ appears $m_{i}$ times.
The cardinality of a multiset $M$ is given by:
$|M|=\sum_{x\in U}M(x)$
which counts the total number of elements, including repetitions.
Definition 6 (Multi-Graph).
Formally, in set-theoretic notation, a multi-graph $G=(V,E)$ is an ordered pair of a set of vertices, $V$, and a multi-set, $E$, of edges.
By using a multigraph approach, it is possible to label bonds with information about the binary bonding relationships between atoms, as well as describing the delocalised electrons that persist over a subset of atoms, which is the approach taken by Dietz [8].
1.6 Dietz representation
The Dietz representation was designed to provide a computational framework capable of representing any chemical structure that can be described by a structural formula. It was developed to overcome the limitations of graph-based and hypergraph-based methods, including string representations that correspond to molecular graphs [8].
The paper describes the “constitution” of a structure in terms of the unshared valence electrons of atoms and the bonding relationships between them. Additionally, the terms “conformation” and “configuration” are used to describe different aspects of the relative spatial relationships between atoms. Conformation refers to the spatial arrangement of atoms that can change without breaking chemical bonds, such as rotations around single bonds, while configuration refers to the fixed spatial arrangement of atoms, including stereochemistry, which requires breaking and reforming bonds to alter. These spatial relationships are illustrated using a unique “paddle wheel” representation.
As the proposed ADT uses atomic coordinate data to describe 3D information of molecules, only the “constitution” of the Dietz representation is discussed here.
The constitution of a structure is represented in set theoretic notation, with two sets. The first set contains all atoms in the structure as triples, containing the number of unshared valence electrons of the atom, a numerical index, and the atomic symbol. The second set describes “bonding systems”, or bonding relationships between atoms, and contains pairs consisting of the number of shared valance electrons, and a set of edges representing bonds over which the electrons are shared.
Definition 7 (Constitution).
The constitution of a molecule is an ordered pair:
$C=(V,B),$
where $V$ is the set of vertices, representing the atoms in the molecule, and $B$ is the set of bonding systems, representing the bonding relationships between atoms.
Each vertex in V is a triple:
$V=\{(u,j,A)\mid u\in\mathbb{N}_{0},\,j\in\mathbb{N},\,A\in\text{Symbols}\},$
where:
  • u is the number of unshared valence electrons for the atom,
  • j is a unique index, and
  • A is the atomic symbol (e.g., $H,C,O$).
Each bonding system in B is a pair:
$B=\{(s,\{\{x,y\}\mid x,y\in V\})\mid s\in\mathbb{N}\},$
where:
  • s is the number of shared electrons in the bonding system, and
  • {{x, y}} is a set of unordered pairs of vertices representing bonds over which the electrons are shared.
Localised bonding is represented by a bonding relationship containing a single set of edges. For instance, a single covalent bond between two atoms would be represented as follows: $(2,\{\{a_{1},a_{2}\}\})$, where $a_{i}$ is the index of each atom.
Delocalised bonding is represented by a bonding relationship containing multiple set of edges. For instance, the delocalised pi electrons of a benzene ring would be described by a bonding relationship of the form $(6,\{\{c_{1},c_{2}\},\{c_{2},c_{3}\},\{c_{3},c_{4}\},\{c_{4},c_{5}\},\{c_{5},c_{6}\},\{c_{6},c_{1}\}\})$, where $c_{i}$ is the numerical index of each carbon atom in the ring. The localised bonding between carbon atoms would be represented by $(2,\{\{c_{i},c_{j}\})$.
This representation is equivalent to a multigraph with additional labelling or metadata to describe the groups of edges with shared electrons.
Example 1 (Constitution of Hydrogen).
The constitution of hydrogen
$V(H_{2})=\{(0,1,H),(0,2,H)\},$
$B(H_{2})=\{(2,\{\{1,2\}\})\}\vskip 6.0pt plus 2.0pt minus 2.0pt$
There are two hydrogen atoms, labeled 1 and 2, each with zero unshared valance electrons in the atomic core, and a one covalent bond with two shared valance electrons.
Example 2 (Constitution of Benzene).
The constitution of Benzene:
$V(Benzene)$ =$=\{(0,1,H),(0,2,H),(0,3,H),(0,4,H),(0,5,H),(0,6,H),$
[…]
The delocalised pi electrons of the benzene ring are described in a single bonding system, as the final element of the set, and separately from all other covalent bonds in the structure.
A major advantage of the Dietz representation is its general model for the representation of bonding. As a result, it is straightforward to describe complex bonding systems such as organometallics, with ferrocene and diborane being given as examples in the original article. The lack of labeling requirements for delocalised bonding stands in contrast to string based representations such as SMILES, which has explicit rules in the grammar (a form of labeling) for aromatic carbon, and two rules for different cases of aromatic nitrogen.
Despite satisfying the three requirements identified in [1] for representation beyond organic chemistry—namely, the ability to represent delocalized bonding, explicit labeling of terminal hydrogen atoms, and advanced stereochemistry such as in cumulenes—Krenn et al. describe the Dietz representation (apparently mischaracterized as using hypergraphs) as one that “leads to complicated nested sets of brackets that may be hard to comprehend” [1].
Although set-theoretic notation may be unfamiliar, the definition or “grammar” of the constitution of a molecule can be efficiently described. This stands in contrast to SMILES strings or SELFIES strings (which as an additional challenge contains overloading of symbols in the grammar [1]).
1.7 SMILES
SMILES (Simplified Molecular Input Line Entry System, [4]) encodes molecular graphs as ASCII strings through a series of rules, representing atoms and bonds as characters and symbols. Introduced in the 1980s, SMILES was designed to facilitate chemical database searching and to enhance portability, and has achieved significant popularity in cheminformatics [12, 13, 1, 15, 16, 17, 18]. However, increasing demands of modern machine learning and cheminformatics reveal challenges that extend beyond the scope of SMILES’ original design.
Corresponding to molecular graphs, SMILES inherit the limitations of this representational framework. While SMILES can represent aromaticity through specific syntax[4], it cannot encode delocalized bonding beyond such cases. Resonance and tautomerism must either be simplified into a single structure or require multiple distinct structures to represent each form.
The encoding of stereochemistry in SMILES has notable constraints. While tetrahedral centers are well-supported through isomeric SMILES, a common extension, the encoding of more complex stereochemical features, such as dynamic or rotational stereoisomerism seen in atropisomers, remains unsupported [19]. Isomeric SMILES can represent certain forms of axial chirality in cumulenes and handle non-tetrahedral stereocenters, including allene-like, trigonal-bipyramidal, and octahedral configurations [19]. However, these representations rely on 2D graph-based encoding and lack support for nuanced 3D spatial arrangements required for complex organometallic systems or multi-center bonding scenarios. The absence of three-dimensional (3D) spatial data further limits SMILES’ utility for applications requiring dihedral angles or precise molecular conformations, a critical aspect of protein-ligand modeling and other machine learning tasks such as property prediction [20, 21, 22].
Beyond these limitations, SMILES suffers from unique structural and functional shortcomings such as syntactic invalidity [5, 12, 13], non-local encoding of features, and non-unique representations of molecules, all of which can impair machine learning models as the models must implicitly learn the SMILES syntax, which can be difficult, and require larger training sets and wasted computation [12, 23].
“Syntactic invalidity” means that not all valid SMILES strings correspond to valid molecules. In Machine learning contexts, algorithms may output syntactic nonsense, and must re-sample or require larger training sets to ensure robustness [12, 24, 5]. Even though one paper claims that invalid SMILES should be seen as beneficial to chemical language models [25], it should be noted that the perceived improvement is only relative to comparisons between SMILES and SELFIES models, and likely reflects the models’ enhanced ability to infer SMILES syntax when generating valid molecules.
In SMILES, small changes in the string can lead to drastic changes in molecular structure, and vice versa – adjacent atoms in the molecule can be far apart in the string [23]. As well as impacting readability, the non-local encoding of structural information can make it difficult to write or even find a similarity metric of SMILES strings whose magnitude corresponds to a relevant similarity of molecules [23].
SMILES representations are non-unique: a single molecule can be expressed by multiple valid strings [4]. As well as complicating database searches [13], machine learning algorithms must discern the equality of different representations. Although some machine leaning approaches have attempted to utilize multiple SMILES for the same molecule in their training sets, and have claimed improvements to their models as a result [26, 27], this is again likely explained by the ability of the model to infer the SMILES syntax more easily. In other words, it exposes a weakness of SMILES for machine learning, not an advantage.
Representations in SMILES also do not give indications as to molecular symmetry. By example, the Cyclopentadienyl anion in SMILES is [cH-]1cccc1, but this does not indicate the rotational symmetry of the molecule, as is clearly shown in Fig. 1. The non-uniqueness of SMILES strings can be seen as 1cccc1[cH-] is also a valid representation, but in SMILES no explicit indication is given as to whether [cH-]1cccc1 = 1cccc1[cH-].
Figure 1: The Cyclopentadienyl anion. Although the structure displays clear rotaional symmetry, this is not indicated by the SMILES representaion: [cH-]1cccc1. Image Credit: Oliver Goldstein (Author).
Attempts to canonicalize SMILES, and the proprietary nature of the Daylight canonicalisation algorithm, have led to inconsistencies between implementations and research teams [12], though the OpenSMILES initiative has mitigated this issue by providing a standardized grammar.
Importantly, as a data type, strings by themselves provide no guarantees about the structure of the validity of the data.
While SMILES and its extensions remains useful tools for encoding and sharing molecular structures, its limitations constrain its applicability in modern cheminformatics and machine learning workflows. Extensions to SMILES to fix problems, and the problems they themselves introduce, are discussed in a later section.
1.8 SELFIES
SELFIES (Self-Referencing Embedded Strings) [5] solves arguably the most pressing issue with SMILES: syntactic invalidity. They create a “100% robust” representation, which would translate to soundness in the terminology of logic or computer science. In particular, they present a Context Free Grammar (CFG) where all sequences of terminal symbols represent molecular graphs (soundness) and all molecular graphs can be represented (completeness – but only up to the limited CFG they have created, see Fig. 1). The CFG of SELFIES can be translated into sum and product notation as every CFG can be represented by an algebraic data type [28].
In attempting to retain semantic as well as syntactic validity after mutation, SELFIES uses a table of derivation rules, with overloading of terms to encode valence information [1]. Although successful in maintaining semantic validity after mutation (with the given set of rules, for small biomolecules), and as sucessfully demonstrated with GANs and VAEs[5], the overloading of terms makes the grammar more complicated. The authors claim that SELFIES may be more readable then SMILES for large molecules, although they acknowlede that “Read-ability is in the eye of the beholder”[5]. We personally behold SELFIES as not being human readable, but invite readers to draw their own conclusions.
Machine learning algorithms would also have to implicitly learn the rules of the grammar [12]. Despite the authors claim that SELFIES may be more “machine readable” this remains a conjecture.
The overloading of symbols also represents a challenge for readability. When discussing previous line notations in [4], Weininger criticized attempts to save space with overloaded terms: “Many of the pitfalls in other line notations can be attributed to overuse of symbols and hierarchical rules based on the length of the final notation. Advances in computer hardware have made these restrictions obsolete.” In this sense, SELFIES represents a retrograde step.
State $\epsilon$ [F] [=O] [#N] [O] [N] [=N] [C] [=C] [#C] [Branch1] [Branch2] [Branch3] [Ring]
$X_{0}$ $X_{0}$ F 0 $X_{2}$ N $X_{3}$ 0 $X_{2}$ $NX_{3}$ $CX_{4}$ $CX_{4}$ $CX_{4}$ ign $X_{0}$ ign $X_{0}$
$X_{1}$ $\epsilon$ F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ ign $X_{1}$ ign $X_{1}$ R(N) R(N)
$X_{2}$ $\epsilon$ F =O $NX_{2}$ 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ B(N, $X_{5}$) $X_{1}$ B(N, $X_{5}$) $X_{1}$ R(N) $X_{1}$ R(N) $X_{1}$
$X_{3}$ $\epsilon$ F =O #N 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ B(N, $X_{5}$) $X_{2}$ B(N, $X_{6}$) $X_{1}$ R(N) $X_{2}$ R(N)$X_{2}$
$X_{4}$ $\epsilon$ F =O #N 0 $X_{1}$ $NX_{2}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ B(N, $X_{5}$) $X_{3}$ B(N, $X_{7}$) $X_{1}$ R(N) $X_{3}$ R(N)$X_{3}$
$X_{5}$ C F 0 N 0 $X_{1}$ $NX_{2}$ $NX_{2}$ $CX_{3}$ $CX_{3}$ $X_{5}$ $X_{5}$ $X_{5}$ $X_{5}$
$X_{6}$ C F =O $NX_{2}$ 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$ =C $X_{2}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$ $X_{6}$
$X_{7}$ C F =O #N 0 $X_{1}$ =N $X_{1}$ =C $X_{2}$C $X_{1}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$ $X_{7}$
$N$ 1 2 3 4 5 6 7 8 9 10 11 12 13 14
Table 1: Derivation rules of SELFIES [5] for molecules in the QM9 dataset.
Again, like SMILES, resonant structures are not directly expressible in SELFIES, and delocalized bonding (other than aromaticity) cannot be expressed, along with tautomers.
Since publication, SELFIES has added the ability to represent chirality, through importation of symbols from SMILES, although some forms produce errors in decoding. However, three dimensional information is not included, and it is unclear how to extend the grammar whilst retaining it’s simplicity and soundness. The spatial arrangement of ligands in square-planar stereochemistry, like that of cisplatin and transplatin, cannot be fully captured without 3D information, and so in exploring the chemical space, SELFIES remains unable to represent potentially crucial information for novel drug discover and property prediction, despite being sound.
Furthermore, E-Z stereochemistry does not in general provide accurate dihedral or torsional angles. Proteins rely on these angles for many tasks, which may be critical for machine learning tasks [29]. As Dayalan explains ([22]) “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [30]. Although SELFIES solves the syntactic invalidity problem of SMILES, many challenges remain.
1.9 The future of SELFIES and of molecular representation - Extensions of SELFIES and future projects
In “SELFIES and the Future of Molecular String Representations,” Krenn et al. (2022) [1], review the limitations of string-based molecular representations in cheminformatics and propose future directions, primarily centered around extending string-based methods.
They highlight several representational difficulties with current molecular string formats, including their inability to effectively capture macromolecular structures, crystal lattices, complex bonding (e.g., organometallics), advanced stereochemistry, precise conformations, and non-covalent interactions such as hydrogen and ionic bonding. These limitations, they argue, are not only present in SMILES and SELFIES but are features of most modern, digital molecular representations in use.
The authors also identify challenges specific to machine learning applications, such as issues arising from the non-local encoding of molecular features within strings, and the impact this has on the latent space of VAEs, as well as the impact overloading of symbols has the efficiency of generative models, as well as human readability.
1.10 Macromolecules and crystals
For macromolecules, Krenn et al. discuss existing approaches such as CurlySMILES, BigSMILES, and HELM, each of which introduces distinct syntactic modifications to account (or not, as the case may be) for repeating subunits, complex connectivity, and stochastic relationships between subunits. They propose BigSELFIES and HELM-SELFIES as future extensions to enable the use of generative models which take strings as input for macromolecular synthesis and design.
In the case of crystals, the authors propose yet another syntax, Crystal-SELFIES, based on the labeled quotient graph of the crystal structure. The formalism of the approach is largely unexplored in the paper, however relies on representation of subunits, in a related manner as with macromolecules.
1.11 Barriers to a general, digital, molecular representation
The representational limitations of current string-based methods are said to rely on three core assumptions:
  1. 1.
    The assumption of localized bonding, with standard valances
  2. 2.
    The implicit or non-explicit representation of terminal hydrogen atoms.
  3. 3.
    The inability to encode advanced stereochemical relationships.
These limitations, the authors argue (and we agree), are an inherent problem with string-based representations in their ability to represent complex bonding (such as with diborane and ferrocene) and advanced stereochemical features. Cis- and trans-platin are also given as examples of molecules current string-based representations cannot represent.
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in [5] article:
Thus, it should be stressed again that in d- and f-block chemistry, as well as main-group organometallic compounds, it is often impossible to assign any particular bond orders without high-level quantum chemical calculations, due to the highly de localized nature of the bonding, where electrons are often spread out over a significant number of atoms, including the metal cen ter itself, the immediately coordinated atoms, and additional ligand groups.
Despite recognizing the need for a general digital molecular representation, capable of handling all of the above issues. The authors do not reconsider the suitability of string-based models, and instaed, the solutions presented primarily involve introducing new grammars, expanding syntax rules, and further overloading symbols, leading to increasingly complex and specialized representations.
1.12 “Molecular programming languages”
In Future Project #8, Krenn et al. (2022) propose a “molecular programming language”. They reason that strings can encode the powerful computational algorithms: “Strings can store Turing-complete programming languages: In the most general case, one can store the source code of computer programs as strings. For example, a Python file is a ‘simple string’, which is executed by the Python interpreter. Python is, of course, a Turing-complete language, which means that strings can encode the most powerful computational algorithms.”
Although the source code of a programming language can be represented in files of text, they should be distinguished from string data types. For instance, they do not support string data type operations, such as insertions, [Concatenation, Contains,indexof, Lowercase, partition, replace, reverse].
Furthermore, if the programming language is represented by a string, the question of how functions within the language are interpreted is unanswered. If they are also strings, they would at some point need to be executed by something other than a string. If they are not strings, then it remains to be understood why the authors claim that strings alone are enough to represent a Turing-Complete language, let alone the utility of such a project.
Given that Turing complete programming languages with support for recursion and logical statements already exist, it is unclear why SELFIES should be augmented to become a programming language in order to represent molecules, rather than simply using a programming language which already exists.
Future Project #9 goes beyond molecular programming languages, and proposes a programming language which is “100% robust” by finding a syntax for a programming language that however you combine elements in the instruction set, a valid program is always produced. It seems this is an attempt to generalise their own robustness of mutation to the most general data structure: programs themselves. However, it remains to be seen what sense this would have, let alone any sensible denotational semantics.
1.13 Impact on machine learning models and empirical verification
From a machine learning perspective, improving representation by expanding the grammar, creating new symbols, and overloading of terms can introduce problems, as deep learning models must implicitly learn the larger, more complicated syntax.
Unlike SMILES, VAEs trained on SELFIES have a robust latent space, where any vector can be decoded into a valid molecule. Despite this, it is not a guarantee that the latent space will be smooth, especially given the non-local encoding of graph information on strings, which both SELFIES and SMILES share.
We welcome the empirical question of the effect of overloading of tokens on machine learning models, which was raised in Future Project #2, and, in light of the result that training string models on randomised representations of molecules has on model results, suggest an experiment to determine the impact of implicit learning of the grammar in the future work section.
1.14 Extensions of SMILES
Despite the introduction of DeepSMILES, which fixed the bracketing issue in SMILES strings, the syntactic invalidity problem was not solved completely [31].
New syntaxes or grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with only SMILES. For instance, there are many syntaxes from just one vendor called Daylight111www.daylight.com.
SMILES was introduced and subsequently standardised in OpenSMILES. CurlySMILES was introduced to annotate SMILES strings. CHUCKLES, its extension, CHORTLES, as well as BigSMILES were introduced to represent polymeric molecules [12]. SMIRKS was introduced as an extension of SMARTS to represent chemical reactions [32] – used in the Open Reaction Database. InChi V1, not part of Daylight, was introduced and can represent basic prototautemerism [33]. SMIRKS allows for string representation of reactions. It does not represent reaction features like temperature or pressure or catalysts. One feature it lacks is the ability to indicate the directionality of the reaction, which took a new language to be implemented ten years after the release of SMIRKS, called RinChi [12], [13]. This was further extended years later, with another language called ProcAuxInfo, which enables the storage of metadata relating to yields, temperature, concentration, and other reaction conditions. Although, this represents definite progress, the profusion of new syntaxes and grammars, and their creative backronyms, create avoidable complexity for the cheminformatician.
Balsa, a fully-specified language designed as a subset of SMILES, has been introduced [34]. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
1.15 Molecular Fingerprints
Molecular fingerprints encode structural or chemical features of a molecule, typically as a vector, and are used for comparison, searching, and modeling in cheminformatics[35, 12].
Different structural features may be encoded depending on the fingerprinting method. Some fingerprints capture topological information, as seen in Atom Pair and Topological Torsion fingerprints, which encode molecular connectivity and atomic relationships based on graph traversal [35, 36]. Others, such as Extended Connectivity Fingerprints (ECFP) and Functional Class Fingerprints (FCFP), encode local atomic environments, iteratively expanding circular neighborhoods around each atom [36, 13]. Pharmacophore fingerprints, including Pharmacophore Pairs (PH2) and Pharmacophore Triplets (PH3), emphasize functional group interactions, encoding features related to molecular recognition rather than direct connectivity [36]. Some fingerprints, such as MinHashed Atom Pair (MAP4) and LINGO, operate directly on molecular strings, such as SMILES, instead of deriving features from a molecular graph [35, 36].
Although often described as vectors, the data type of fingerprint representations which use hashing functions do not formally meet the definition of a vector over a vector space.
Definition 8 (Vector Space).
A vector space over a field $\mathbb{F}$ is a set $V$ equipped with two operations:
  • Vector addition: A binary operation $+:V\times V\to V$ that assigns to every pair of vectors $u,v\in V$ a vector $u+v\in V$.
  • Scalar multiplication: A binary operation $\cdot:\mathbb{F}\times V\to V$ that assigns to every scalar $\lambda\in\mathbb{F}$ and every vector $v\in V$ a vector $\lambda v\in V$.
These operations must satisfy the following axioms for all $u,v,w\in V$ and all scalars $\lambda,\mu\in\mathbb{F}$:
  1. 1.
    (Associativity of addition) $(u+v)+w=u+(v+w)$.
  2. 2.
    (Commutativity of addition) $u+v=v+u$.
  3. 3.
    (Existence of an additive identity) There exists an element $0\in V$ such that $v+0=v$ for all $v\in V$.
  4. 4.
    (Existence of additive inverses) For every $v\in V$, there exists an element $-v\in V$ such that $v+(-v)=0$.
  5. 5.
    (Distributivity of scalar multiplication over vector addition) $\lambda(u+v)=\lambda u+\lambda v$.
  6. 6.
    (Distributivity of scalar multiplication over field addition) $(\lambda+\mu)v=\lambda v+\mu v$.
  7. 7.
    (Associativity of scalar multiplication) $(\lambda\mu)v=\lambda(\mu v)$.
  8. 8.
    (Existence of a multiplicative identity) There exists a scalar $1\in\mathbb{F}$ such that $1v=v$ for all $v\in V$.
In hashed fingerprints, such as MAP4, the hashing algorithm SH1 is used to distribute data uniformly over a range of outputs, in a way whic is non-invertible, non-continuous, and irreversible. Since SHA-1 and MinHashing introduce non-linearity in the MAP4 fingerprinting process, the transformation from molecular structure to MAP4 does not preserve vector space properties such as addition, scalar multiplication, or linear combination.
Consider the following hypothetical MAP4 fingerprints of molecules A and B:
$\mathbf{u}$ $=MAP4(A)=(101,203,405,512)$
$\mathbf{v}$ $=MAP4(B)=(98,207,410,515)$
Operations such as vector addition, or scalar multiplication are not well-defined, and it is unclear what additive or multiplicate identities, or additive inverses would represent. For instance, are the following meaningful operations in the vector space, and do they correspond to valid molecular fingerprints?
$\mathbf{u}+\mathbf{v}$ $=(101+98,203+207,405+410,512+515)$
$=(199,410,815,1027)\quad?$
[…]
$=(0,0,0,0)\quad?$
$2\mathbf{u}$ $=(2\times 101,2\times 203,2\times 405,2\times 512)$
$=(202,406,810,1024)\quad?$
The term “vector” may be a terminological convenience, but it is not an accurate reflection of the data type. It may well be detrimental to cross-field communication to see simple mathematical terminology, which has a rigorous definition, carelessly used. However, it appears that cheminformaticians are more concerned with computing the Tanimoto similarity [37] or the dot product between two vectors, rather than taking into account the full definition of a vector.
There are also drawbacks to fingerprints such as their inability to consistently represent tautomers. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [38]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
Moreover, while fingerprints can capture the presence of specific substructures within molecules, they are susceptible to bit collisions and vector sparsity, which limit their overall representational power [23]. Bit collisions occur when multiple distinct substructures are mapped to the same bit in the fingerprint, leading to ambiguity and potential false positives. Vector sparsity arises when the fingerprint representation is high-dimensional, but only a small fraction of the bits are actively used, resulting in inefficient memory usage and computational overhead.
A potential issue with purely floating point or integer vectors, such as fingerprints, is the lack of corresponding semantics in the molecular world. Without additional contextual information or a predefined mapping, the numeric values may represent invalid or physically impossible molecules. This may well hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [35], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint.
1.16 RDKit as a Representation
RDKit is widely used in cheminformatics for molecular manipulation, property prediction, and conversion between molecular representations. However, its primary function is not to serve as a standalone molecular representation but rather to process, transform, and analyze existing representations such as SMILES strings and MOL/SDF files.
Internally, RDKit uses the Adjacency List from the Boost Graph Library to represent molecules as 2D graphs with additional metadata. Bonds are extensively labeled with predefined bond types, while stereochemical features are captured through limited labeling schemes. This approach inherits the fundamental weaknesses of graph-based representations previously discussed, including ambiguity in encoding stereochemistry and the necessity of extensive labeling to capture molecular nuances, although 3D information may be stored in the “Conformer” object.
While RDKit’s internal representation is based on adjacency lists, we refrain from formally describing its data type as such, as we lack proof of its correctness. The complexity of RDKit’s implementation makes it difficult to verify whether the adjacency list truly functions as intended. Moreover, as demonstrated in Listing 1 and 2, mutable state introduces the potential for undefined behavior, reinforcing concerns about the safety and correctness of the representation.
RDKit’s reliance on a non-functional paradigm means that operations on molecular graphs can lead to unintended side effects, making the system more prone to errors, although we acknowledge this criticism is not unique to RDKit, and is part of a broader criticism of the non-functional programming language paradigm.
The reliance on labeling in RDKit exemplifies the inherent issues of such an approach. For example, a bond order of 2.5 is labeled as TWOANDAHALF, an inelegant solution reminiscent of criticisms raised by Dietz on the limitations of labeled representations. Instead of a structural or algebraic description of bond order, RDKit resorts to a categorical label, which may introduce inefficiencies and inconsistencies when interfacing with other systems.
Datamol, a higher-level library built atop RDKit, further highlights the limitations of RDKit as a representational system. Primarily designed for molecular conversion and computation, Datamol features 45 “featurizers”, all of which take SMILES strings as input. The framework is not optimized for storage or transmission but rather for extracting graph-based molecular features for machine learning applications. This dependence on SMILES as an intermediary highlights a fundamental issue: Datamol and similar tools do not provide an independent molecular representation but instead rely on conversion mechanisms.
1Some care must be taken when changing the structure of a graph (via adding or removing edges). Depending on the type of adjacency_list and on the operation, some of the iterator or descriptor objects that point into the graph may become invalid. For example, the following code will result in undefined (bad) behavior:
2
3typedef adjacency_list<listS, vecS> Graph; // VertexList=vecS
4Graph G(N);
5// Fill in the graph...
6
7// Attempt to remove all the vertices. Wrong!
8graph_traits<Graph>::vertex_iterator vi, vi_end;
9for (tie(vi, vi_end) = vertices(G); vi != vi_end; ++vi)
10 remove_vertex(*vi, G);
11
12// Remove all the vertices. This is still wrong!
13graph_traits<Graph>::vertex_iterator vi, vi_end, next;
14tie(vi, vi_end) = vertices(G);
15for (next = vi; vi != vi_end; vi = next) {
16 ++next;
17 remove_vertex(*vi, G);
18}
Listing 1: Example description of undefined behaviour which can cause errors, taken from the Boost GraphLibrary documentation, illustraing nature of data type and operations which can be performed as of Wed 5 Feb 2025.
1// ROFIX: these will need to be readonly somehow?
2typedef std::map<int, ATOM_PTR_LIST> ATOM_BOOKMARK_MAP;
3typedef std::map<int, BOND_PTR_LIST> BOND_BOOKMARK_MAP;
Listing 2: Example listing as of Wed 5 Feb 2025.
Moreover, RDKit’s use of null pointers introduces potential points of failure, reflecting broader concerns about type safety in non-functional programming paradigms. The lack of strong guarantees about memory safety and the absence of robust type constraints can lead to errors that are difficult to detect and debug. While this criticism extends beyond RDKit to imperative and object-oriented programming more generally, it underscores the representational fragility of systems that rely on mutable state and loosely enforced data structures.
1.17 Molecules and programming languages
Inductive Logic Programming (ILP), has seen usage in representing basic molecules [39], but the focus was doing rule induction over graphical data structures, with molecules being an example, rather than presenting ILP as a representational format for cheminformaticians. It is perhaps for this reason that it has had little uptake within the cheminformatics community.
Molecular representations using Inductive Logic Programming (ILP) have been explored in previous research, such as the work by Srinivasan et al. [39]. However, the application of machine learning techniques in conjunction with ILP-based molecular representations has not been extensively discussed in this early work. This lack of emphasis on machine learning aspects may explain the absence of the paper from recent reviews of molecular representations, such as the comprehensive survey by Wigh et al. [13].
In 2018, de Meent et al. [40] argue that the recent successes of artificial intelligence have been caused by the emergence of programming language tools, such as NumPy, that are able to automatically engage in the “tedious” derivation and calculation of gradients. Probabilistic programming languages build on this idea to provide a generic way to write arbitrary models and then calculate the uncertainty of parameter choices.
In the approach presented by Srinivasan et al. [39], molecules are represented as logical rules using an extension of the Prolog programming language. Prolog is a declarative language that enables the expression of logical relationships between variables using first-order logic. Programs written in Prolog consist of a complex network of facts and rules, which are constructed using logical primitives such as conjunction, disjunction, and universal quantification. However, the data types supported by Prolog are limited to lists, integers, floats, and strings, with minimal support for user-defined data types.
Prolog’s inference mechanism allows for querying the knowledge base by providing examples and evaluating their truth values based on the presence or absence of counterexamples (negation as failure). Nevertheless, it is important to note that Prolog and its extensions are not purely declarative, as they incorporate side effects through built-in predicates such as read, get, and assert. These side effects can succeed or fail independently of the problem being solved, introducing challenges in reasoning about the program’s behavior, particularly when dealing with complex rule bases.
The specific extension of Prolog used in [39] is called Progol, which employs inductive logic programming techniques to derive rules that describe the given data. It is worth clarifying that Progol should not be confused with ProbLog, a probabilistic logic programming language that focuses on modeling the probabilities of examples being true based on the probabilities of the underlying relations.
While the work by Srinivasan et al. [39] demonstrates the feasibility of representing molecules using logical rules in an ILP framework, the integration of machine learning techniques with these representations was not explored.
1.18 Functional languages and cheminformatics
Work using functional programming in cheminformatics has been done before, however the focus of such efforts were usually related to ease of writing or safety of code, or efficiency of functional programming languages, rather than attempting to introduce a new representation by explicitly considering data types.
Ouch [41] is a chemical informatics toolkit written entirely in Haskell, released in 2010. It was released “in the hope that it will be useful”, but has seen little development or since 2013. Internally, molecules are represented with a molecular graph, implemented as an algebraic data type, where bond types are categorically labeled (e.g. “sigma”, “pi”, or “aromatic”), and without any support for stereochemical features or 3D information.
Chemf[10] is a toolkit written in Scala. Its purpose was not to create a molecular representation, but to permit existing cheminformatics workflows to utilise the functional programming paradigm, and to avoid the problems which can occur with mutable state and null pointers. The representation used is a 2D molecular graph written as a connectivity list.
“Radium” is partially complete Haskell library. It has support for atomic orbital information, but molecules may only be encoded as a linear notation based upon SMILES strings. Interestingly, the data types used for molecular representation are algebraic, and while being based upon SMILES syntax, are not in fact “strings”.
Other work [9] has suggested using OCaml for cheminformatics, however this work was primarily written as an introduction to chemists for functional programming, with a granular explanation of specific functions, and a focus on ease of writing code, type safety, and efficiency, such as the ease of parallelization.
This makes the proposal here the first to suggest to cheminformaticians use constructs rooted in functional programming languages directly in the representation of molecules themselves. Additionally, unlike previous work in functional programming, we attempt to expand the scope of molecular representation, and explicitly consider the role of the data type and how this affects machine learning, including with concepts such as Bayesian machine learning and Geometric deep learning.
1.19 Geometric Deep Learning on Molecular Representations
Geometric deep learning (GDL) extends deep learning techniques to non-Euclidean domains such as graphs, manifolds, and meshes[2]. In the context of molecular modeling, molecules can be represented as graphs, where atoms serve as nodes and chemical bonds as edges. This formulation allows GDL methods, particularly graph neural networks (GNNs), to learn representations that capture topological and geometric invariances or equivariances of molecules [12, 13].
A key advantage of GDL in cheminformatics is its ability to incorporate geometric priors, such as rotational and permutation invariance, ensuring that appropriate properties are invariant to the molecular representation at hand, in particular the ID of an Atom (by leveraging permutation invariance). These methods have already been applied successfully in deciphering interaction fingerprints from protein molecular surfaces [42].
The application of GDL is particularly relevant in addressing the limitations of traditional string-based molecular representations, such as SMILES, which encode chemical connectivity but neglect spatial conformation and stereochemistry [1]. By directly operating on molecular graphs and 3D structures, GDL-based approaches provide a more expressive and chemically meaningful representation, enhancing predictive performance in downstream tasks [13].
Our representation of the bond matrix, is indeed a graph structure, computationally implemented with a Haskell Map for the sake of speed. Thus, graph algorithms from the field of geometric deep learning can be quickly and simply applied to our representation. Given these advantages, GDL presents a compelling framework for molecular representation, motivating its adoption as an alternative to conventional symbolic encodings in cheminformatics.
1.20 Probabilistic programming
Typically, in conventional programming, one provides inputs to a function to compute outputs. However, in Bayesian inference, the process is reversed: given observed outputs, we infer the likely inputs and function parameters that could have produced them. For instance, given a model that outputs the solubility, for which we have the values, the inference process determines the parameter values that generated those data given in the training set, with the aim of finding parameter values that generalise the underlying data distribution. In the context of computational chemistry, the input parameters could represent any variable (parameter) that determines the property of a molecule, in a particular representation. This means the feasibility of sampling is fundamentally constrained by the operations permitted on the type of data at hand. If this was used with a String datatype, the notion of sampling and scoring becomes.
This approach, is known as Bayesian inference and is automatically performed by probabilistic programming languages. It is an increasingly popular method to perform machine learning that enjoys strong mathematical foundations [43]. Large technology companies such as Uber, Google and Microsoft, have invested in their own probabilistic programming languages (Pyro [44] Infer.NET [45], Edward [46],respectively). Academia has also focussed on building modular probabilistic programming languages which allow composing inference algorithms to build other inference algorithms [47, 48, 49].
Bayesian Inference has useful properties, compared to deep learning based machine learning, such as interpretability, explainability and expressiveness with respect to various forms of uncertainty [50, 51]. At the heart of the Bayesian paradigm is the notion of a generative model. Generative models can account for the expressiveness of uncertainty and interpretability. Probabilistic programming languages provide a generic framework to perform inference. Bayesian inference has already been used in drug design e.g. [52], where a Bayesian neural network is used to predict drug toxicity.
A Bayesian generative model is defined as a program that at least contains the following two operations:
  • The sample effect (or operation), equips models with an operation that returns a uniformly distributed probability floating point value from the unit interval, expressing probabilistic non-determinism. This is a natural operation in many areas of science which use probability density functions such as sampling from an Orbital or the radioactive decay of a particle (subject to building an appropriate distribution).
  • The score effect (or operation), provides a mechanism to calculate the likelihood of model samples, given observed data.
There are three well-established methods of Bayesian inference:
  • Monte-Carlo Inference – Uses sampling-based methods like MCMC to approximate posterior distributions in probabilistic programs.
  • Variational Inference – Converts inference into an optimization problem by approximating the posterior with a simpler distribution.
  • Exact Inference – Computes the posterior distribution analytically when closed-form solutions exist, often feasible only for simple models.
In this article, for the sake of simplicity, we will use Trace-Metropolis Hastings, although the example below uses rejection sampling which is another Monte-Carlo method. Rejection sampling becomes inefficient in high-dimensional spaces, where too many samples get rejected. In Trace-Metropolis Hastings, a sequence of random samples is generated by iteratively proposing new parameter values and accepting or rejecting them based on a probability ratio. This process explores the posterior distribution by permuting the sampled values to refine the parameter estimates that best generate the observed data.
In the context of molecular representation, probabilistic programming offers a natural way to capture the inherent uncertainty and variability associated with molecular structures and properties. By combining the expressive power of functional programming with the ability to represent and reason about probability distributions, PPLs enable the development of rich, probabilistic models of molecules. Consider a simple example involving inferring the bias of a coin in webPPL [53].
Example of inferring the bias of a coin with probabilistic programming:
1 // Inferring the bias of a coin using rejection sampling.
2 var heads = true, tails = false;
3 var observed_data = [heads, tails, heads, heads]
4
5 var model = function() {
6 var prior = sample(Uniform({a: 0, b: 1}))
7 var likelihood = function(p) { return bernoulli(p); };
8 var samples = mapData({data: [prior, prior, prior, prior]}, likelihood);
9 condition(samples.toString() $=$ observed_data.toString())
10 return prior
11 }
12
13 // Plot the result of inference.
14 viz(Infer({method: rejection’, samples:1000}, model))
Listing 3: Inferring the bias of a coin using rejection sampling in WebPPL
Figure 2: WebPPL inference graph of biased coin inference, based on the code. The most likely bias parameter is 0.75, which makes sense as 3/4 of the coins are heads (true).
The prior is sampled from the uniform interval $[0,1]$ and the likelihood function is Bernoulli, assuming values in the set $\{0,1\}$. The conditioning process is explicit with condition. Conditioning is the fundamental method in which statistical inference over parameters is done [54]. Only program traces that fulfill the condition are considered by the infer method, where a trace, defined by [54], is simply the tree data structure that captures the random choices encountered and the path taken by the interpreter while evaluating the program. The condition operation is a specialization of score to be either exactly the observed data or not. Some scoring functions use distance measures.
The random primitives are in this case real-valued random variables and Boolean random variables. The method of inference used here is a non-Markov Monte Carlo approach, termed rejection sampling. Intuitively, this can be understood by considering the traces of the program that progress past the condition statement, which will generally be dominated by prior values that tend to yield the observed data. Based on observed data, one could also include knowledge about scientific laws restricting the world, resulting in $\theta$ values that are most “realistic”.
In [55, 56], there are discussions of the importance of classifiers providing not just a single output, but two key pieces of information: a class label and an associated probability. This dual output allows for a more nuanced understanding of the classifier’s confidence in its predictions. For instance, in binary classification tasks, a classifier might assign a label (e.g., ‘spam’ or ‘not spam’) and also provide the probability of the instance belonging to that class. This probabilistic information is crucial for tasks such as ranking, where items are ordered based on the likelihood of belonging to a particular class, and for making informed decisions in applications where the costs of false positives and false negatives differ. [56] emphasizes that effective classification involves not only assigning the correct label but also accurately estimating the probability associated with that label to reflect the classifier’s certainty. In the case of SMILES, an estimate of 0.5 that it the molecule is soluble may mean exactly that or it may be saying that it is an example that is has never seen before. Without a grammar that tells one what is valid and what is invalid, it relegates the probability estimates to be ambiguous and potentially meaningless.
1.21 Laziness
In cheminformatics, we often deal with large datasets of chemical structures (e.g. millions of molecules) and sometimes expensive analyses (e.g. molecular property calculations, conformer generation, docking runs). A lazy programming language like Haskell can help manage these challenges by delaying computations until they are truly needed, and by automatically reusing results once they have been computed.
In many programming languages, as soon as you write an expression—for example, a function call or a chemical property calculation—the language runtime will typically evaluate it immediately. In a lazy language, however, expressions are turned into thunks: little bundles of code waiting to be executed. The language defers (delays) performing the actual computation until it is absolutely required, such as when the result is needed to proceed with another step in the program.
Suppose we define a huge list of candidate molecules, each annotated with an expensive-to-calculate property (e.g. predicted solubility). In a strict language, as soon as we construct this list, the program might try to calculate the property for every molecule upfront, even if we only want the first few results. In a lazy language, by contrast, the expensive property calculations are wrapped up in thunks. These thunks remain unevaluated until you actually iterate over the list and access the relevant element.
Thus, if you only access the first 10 molecules in the dataset, only 10 calculations take place; the remaining 999,990… molecular properties remain as unevaluated thunks.
A second important aspect of laziness is sharing:
Once a thunk has been computed, its result is cached (or memoized) so that future accesses do not repeat the same work. If multiple parts of your cheminformatics workflow depend on the same expensive molecular descriptor, the lazy evaluation mechanism ensures that once you compute it once, the result is re-used everywhere else automatically.
  • On-Demand Computation: We might have large pipelines combining multiple steps (e.g. filtering by molecular weight, docking simulations, then clustering). If we never end up needing a particular property, we waste no effort computing it in a lazy setting.
  • Improved Compositionality: With lazy evaluation, we can write “infinite” or “very large” data streams and process them in a streaming fashion without ever fully materializing them in memory. This is particularly helpful when investigating structure-generation procedures or enumerations in a conceptually infinite chemical space, since only the portion of the stream you consume is actually computed.
  • Parallelization and Caching: Lazy languages can simplify certain parallelization strategies since once a thunk is forced in one thread, its result can be shared across other threads (in purely functional languages like Haskell).
2 Methodology
2.1 Our Algebraic Data Type
An Algebraic Data Type (ADT) is a composite data structure formed by combining simpler data types using sum types (representing alternatives) and product types (representing combinations), which obey algebraic laws [57]. It is a fundamental concept in functional programming, with links to universal algebra, that provide a way to define complex, hierarchical, or (iso)recursive structures, first introduced in 1980 with HOPE [58, 57]. They define the data structure, and its type with which typeclasses then define the operations that can be meaningfully performed on that data, through instantiation.
Definition 9 (Algebraic Data Type).
An algebraic data type (ADT) over a set of base types is any type built up from:
  1. 1.
    Sums (coproducts): If $A$ and $B$ are types, then $A+B$ is the type of “either an $A$ or a $B$”;
  2. 2.
    Products (Cartesian products): If $A$ and $B$ are types, then $A\times B$ is the type of pairs $(a,b)$ with $a\in A$ and $b\in B$;
  3. 3.
    Recursion via a least fixpoint: Suppose $F$ is an operation that takes a type and returns a new type. The type $\mu X.\,F(X)$ is then the smallest (or simplest) solution to the equation $T\cong F(T)$. Intuitively, $T$ is built exactly in the shape that $F$ prescribes, by referencing itself over and over as needed.
    This construction gives a natural way to represent (iso)recursive data structures. Because $T$ can contain smaller copies of itself, you can define objects like lists, trees, and other self-referential shapes. Calling $\mu X.\,F(X)$ the “least fixpoint” emphasizes that it includes no extra elements beyond what arise from repeatedly applying $F$. In practice, this ensures that the type exactly captures the recursive shape you intend, without unnecessary complexity.
Concretely, any recursive ADT can be expressed as
$T\;=\;\mu X.\,F(X),$
where $F$ uses only sums, products, and previously defined base types.
For those, familiar with category theory, $\mu X.\,F(X)$ is the initial $F$-algebra, ensuring that it is the “smallest” type satisfying $T\cong F(T)$. Equirecursive type theories treat $\mu X.\,F(X)$ and $F(\mu X.\,F(X))$ as definitionally equal, while isorecursive theories regard them only as isomorphic via explicit fold/unfold operations.
Our molecular Algebraic Data Type is written in Haskell [59], a lazy, purely functional, type-safe language, with type classes and algebraic data types. In languages like Haskell, the recursion is effectively isorecursive: data definitions (e.g. data List a = Nil | Cons a (List a)) introduce an implicit fold/unfold mechanism (via constructors and pattern matching) even though it appears syntactically direct.
Our Algebraic Data Type represents molecular constitution with the Dietz representation, three dimensional chemical information via atomic coordinate data, and provides a novel representation of quantum information via the Orbital type. Through the use of Haskell’s type classes, we specify allowed components and combinations of molecules, ensuring only semantically meaningful structures are represented. Additionally, algebraic data types can be shown to be a group (or a monoid etc), meaning rotational symmetry of molecules can be described.
By using the Dietz representation’s general model for bonding systems, the ADT is inherently able to represent delocalised bonding in general, without extensive labelling of bond types. Resonance structures, organometallics, multi-center and electron deficient bonds may all be represented. Fractional, and zero bond orders may also be calculated.
Using atomic coordinate data, all molecules with stereoisomerism may be represented and distinguished, and dihedral bond angles, torsion angles, and bond lengths may be calculated. Such information adds minimally to the space complexity, but yet can be crucial when predicting physical propreties of drugs. In our ADT, we have dihedral and bond angles. We also have the ‘meta’ characters as operations in Haskell ([1]) such as folds, map and forM_ that together can capture the essence of a for loop and filter or takeWhile for a while loop (perhaps in conjunction with a State monad).
The Dietz represenation of constitution, combined with atomic coordinate data then solves the three identified limitations of SELFIES for going beyond organic chemistry[1]: the ability to represent delocalised bonding, explicit hydrogen representation, and the ability to represent complex stereochemistry.
The novel orbital data type creates a data structure to model shells, subshells, and orbitals, completely representing the electronic structure of an atom. Quantum information can then be used for machine learning applications, and for quantum chemical calculations in Molecular Orbital theory, as can be required with organometallics.
Apart from the obvious benefits of being able to represent molecules which cannot currently be represented using string-based representations (even with the creation of new and complicated grammars), our ADT confers a crucial benefit. By directly synthesizing molecules in accordance with the grammar defined by the ADT, machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation into another format.
2.2 Haskell
Haskell’s type classes denote which data types support specific operations, and in some cases can also specify laws that the operations should satisfy. By making data types an instance of a type class, we specify that the data type supports the defined operations, and ensure the operations adhere to the operational laws.
What is novel is the integration of probabilistic programming with Haskell into a concrete use case. Probabilistic programming languages themselves in haskell have been created before [60, 48, 47, 49], laying the groundwork for usage with chemical structures. The existence of libraries such as LazyPPL[61] (demonstrated later in the article), makes the ADT possible to use with Bayesian machine learning applications, such as drug design, which have been attempted before [62, 52, 60], and is an emerging area of research. The “lazy” (non strict, or call-by-need) property of Haskell, as discussed earlier, can provide efficiency, for instance when sampling from expensive functions or performing simulations only when necessary. Furthermore, it can represent arbitrarily large finite data structures, which sits between an extensional notion of infinite and finiteness. This is useful later on as [1] quotes Generally, a graph with an infinite number of edges cannot be described by a string of finite length. Fortunately, a net can be represented by a finite graph, known as its quotient graph.
The Maybe and Nothing types provide a type-safe way to handle operations that may fail, distinguishing between empty and zero values, a concept that in other languages might be handled by null references or exceptions, which often introduce program crashing errors. This distinction is later used in the orbital data type, to denote the absence of a SubShell.
Additionally, Haskell’s type-safe type system offers compile-time safety. This can completely eliminate a class of runtime errors commonly encountered in dynamically-typed languages like Python. Machine learning applications written with Python, or the PyTorch framework, must contend with the possibility of type-related errors during execution, even with type-annotation.
It should be noted that concepts such as Algebraic Data Types, or Functional Purity, are not unique to Haskell, therefore Haskell is not strictly necessary for the implementation.
Record syntax in Haskell also enables the access of only parts of the molecule needed for a task, e.g. only taking the atomicSymbol of the atoms in a molecule, for a machine learning task. Non-necessary data is not accessed, making machine learning models potentially less wasteful. This is difficult to do with non-local encoding of information, such as with SMILES or SELFIES strings, where a whole molecule, as a string, must be parsed.
Definition 10 (Record Types in Haskell).
A record type is a Haskell data type whose constructors include named fields, enabling direct access to components via automatically generated selector functions.
Record syntax is now briefly discussed, before introducing the ADT itself.
2.3 Record Syntax
Record syntax allows for defining named fields within data types, facilitating direct access to individual parts of a data structure. The use of these names acts as a way to prevent confusion about whether one has a specific atom or coordinate in a data structure representing atoms, for instance.
Listing 4 shows a simple example. Atom is a sum (coproduct) type with two constructors: AtomS and AtomC. The AtomS constructor represents an atom with a symbol and has two fields: atomID of type Integer and symbol of type AtomicSymbol. The AtomC constructor represents an atom with a symbol and a coordinate, and it has three fields: atomID of type Integer, symbol of type AtomicSymbol, and coordinate of type Coordinate.
1data Atom = AtomS { symbol :: AtomicSymbol }
2 | AtomC { symbol :: AtomicSymbol,
3 coordinate :: Coordinate }
4
5data AtomicSymbol = C | O | H
6
7data Coordinate = Coordinate { x :: Double,
8 y :: Double,
9 z :: Double }
Listing 4: Haskell record syntax showing a simple algebraic data type
Record syntax automatically generates getter functions for each field. Listing 5 shows the updateXCoordinate function which updates the coordinate of an AtomC. It takes a new Coordinate and an Atom as arguments. If the atom is an AtomC, it uses the @ pattern to bind the entire AtomC value to the variable atomWC. This allows us to refer to the original AtomC value when updating its fields. We then use record syntax to create a new AtomC by updating only the coordinate field of atomWC with the newCoordinate, while keeping the atomID and symbol fields unchanged. If the atom is an AtomS, it returns the original atom unchanged since AtomS does not have a coordinate field.
1updateXCoordinate :: Double -> Atom -> Atom
2updateXCoordinate newX (AtomC atomID symbol coordinate) =
3 AtomC atomID symbol coordinate { x = newX }
4updateXCoordinate _ atom = atom
Listing 5: Modifying only the x-coordinate of an atom
The getXCoordinate function shown in Listing 6 retrieves the x-coordinate of an AtomC. It uses pattern matching to extract the coordinate field from the AtomC constructor. If the atom is an AtomC, it uses record syntax to access the x field of the coordinate and returns it wrapped in Just. If the atom is an AtomS, it returns Nothing since AtomS does not have a coordinate field.
1getXCoordinate :: Atom -> Maybe Double
2getXCoordinate (AtomC _ _ coordinate) = Just (x coordinate)
3getXCoordinate _ = Nothing
Listing 6: Retrieving the x-coordinate of an atom
Record syntax in Haskell provides a basic level of extensibility for data types, allowing the addition of new fields without requiring changes to existing functions that operate on the data type. This capability enables the evolution of data models with minimal impact on the codebase. This extensibility is particularly valuable in the context of molecular representation, as it allows for the incorporation of new scientific knowledge and the adaptation of the representation to specific domains or applications. However, for more sophisticated and flexible data manipulation, lenses offer a powerful approach. Lenses are first-class getter & setter functions that can be composed to access and modify deeply nested fields or manipulate multiple fields simultaneously, simplifying the process of working with complex or nested data structures by focusing directly on the target fields, bypassing the need for explicit navigation.
Record syntax in Haskell offers a robust mechanism for defining algebraic data types. They simplify the process of working with data by automatically generating getter and setter functions, thus reducing boilerplate code. This feature not only enhances code readability but also supports the seamless extension of data types without impacting existing functionalities.
3 Results and discussion
3.1 The Molecule ADT
The Molecule data type is shown in Listing 7:
1data Molecule = Molecule {
2 atoms :: [Atom],
3 bonds :: M.Map (Integer, Integer) BondType
4 } deriving (Show, Read)
5
6data Atom = Atom {
7 atomID :: Integer,
8 atomicAttr :: ElementAttributes,
9 coordinate :: Coordinate,
10 shells :: Shells
11 } deriving (Show, Read)
12
13data BondType = Bond {delocNum :: Integer, atomIDs :: Maybe [Integer]}
14 deriving (Eq, Read, Show)
15
16data AtomicSymbol = O | H | N | C | B | Fe | F | Cl | S | Br | P | I
17 deriving (Eq, Read, Show)
18
19data ElementAttributes = ElementAttributes
20 { symbol :: AtomicSymbol,
21 atomicNumber :: Integer,
22 atomicWeight :: Double
23 } deriving (Eq, Read, Show)
24
25type EquilibriumBondLength = Angstrom
26newtype Angstrom = Angstrom Double deriving (Read, Show, Eq)
Listing 7: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent Bonds (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over).
The Molecule data type consists of two fields: atoms and bonds. The atoms field is a list of Atom values, representing the individual atoms that make up the molecule. Each Atom is identified by a unique atomID of type Integer, ensuring that each atom can be distinctly referenced within the molecule. The Integer type in Haskell represents arbitrary-precision integers, allowing for the representation and manipulation of integers without any predefined size limits, thus avoiding the issue of overflow that can occur with fixed-size types like Int. Unlike Int, which is limited to a fixed range determined by the computer’s architecture, Integer can handle values as large as the available memory permits.
The Atom data type contains several important attributes. The atomicAttr field of type ElementAttributes captures the fundamental properties of an atom, such as its atomic symbol (AtomicSymbol), atomic number (atomicNumber), and atomic weight (atomicWeight). These attributes provide essential information about the nature and characteristics of each atom in the molecule.
In addition to the atomic attributes, each Atom also has a coordinate field of type Coordinate, which represents the spatial position of the atom within the molecule. Each Coordinate has an x, y and z field. This information is crucial for understanding the geometric structure and spatial arrangement of the atoms, enabling the analysis of molecular shape, conformation, and reactivity.
Furthermore, the shells field of type Shells represents the electronic configuration of the atom, capturing the arrangement of electrons in the various atomic orbitals. This information is useful for understanding the chemical behavior and bonding properties of the atom, as well as its participation in chemical reactions. This will be discussed in the section on the Orbital type.
The bonds field of the Molecule data type is represented using a Map, with the keys being tuples of Integer values representing the atom IDs of the bonded atoms, and the values being of type BondType. This representation acts as an adjacency matrix and allows for the efficient lookup and manipulation of bonds within the molecule.
The BondType data type includes delocNum and atomIDs, which represent the delocalization number and the optional ring structure associated with the covalent bond. This allows for the representation of delocalized electrons and cyclic structures within the molecule. This allows for the accurate modeling of the dynamic distribution of electrons across multiple bonds within a molecule, capturing the essence of the Dietz representation.
By incorporating this information, the data type can effectively represent resonant structures and account for the delocalization of electrons, which is crucial for understanding the electronic properties and reactivity of molecules. The atomIDs field, of type Maybe [Integer], provides a means to represent the interconnected bonds within a cyclic structure. It is an optional field, indicated by the Maybe type constructor, allowing for the presence or absence of a list of atomIDs across which electrons are delocalised associated with the covalent bond. This list of integers enables for instance the representation of cyclic structures within the molecule. This design choice of using the delocNum and atomIDs fields in the Bond constructor offers a flexible framework capable of accurately depicting complex molecules, such as organometallic compounds like ferrocene. It provides a foundation for computing fractional bond orders, which is essential for molecules with resonant bonding characteristics, such as ozone ($O_{3}$).
3.1.1 Example of a Molecule : Benzene
1benzene :: Molecule
2benzene = Molecule { atoms = [Atom i (elementAttributes e) (Coordinate x y z)
3 | (i, e, x, y) <- atomsData],
4 bonds = M.fromList \$ getSymmetricBonds bondsData }
5
6atomsData :: [(Int, Element, Double, Double, Double)]
7atomsData = [ (1, C, 0.0, 1.3948, 0.0), (2, C, 1.2079, 0.6974, 0.0), (3, C, 1.2079, -0.6974, 0.0),
8 (4, C, 0.0, -1.3948, 0.0), (5, C, -1.2079, -0.6974, 0.0), (6, C, -1.2079, 0.6974, 0.0),
9 (7, H, 0.0, 2.4732, 0.0), (8, H, 2.1431, 1.2366, 0.0), (9, H, 2.1431, -1.2366, 0.0),
10 (10, H, 0.0, -2.4732, 0.0), (11, H, -2.1431, -1.2366, 0.0), (12, H, -2.1431, 1.2366, 0.0) ]
11
12bondsData :: [((Int, Int), Bond)]
13 -- 6 single CC bonds between neighboring carbon atoms
14bondsData = [ ((1,2), single), ((1,6), single), ((2,3), single),
15 ((3,4), single), ((4,5), single), ((5,6), single),
16
17 -- 2c2e bonds between carbon and hydrogen atoms
18 ((1,7), single), ((2,8), single), ((3,9), single),
19 ((4,10), single), ((5,11), single), ((6,12), single),
20
21 -- delocalised bonding in the benzene ring
22 ((1,2), aromatic), ((1,6), aromatic), ((2,3), aromatic),
23 ((3,4), aromatic), ((4,5), aromatic), ((5,6), aromatic) ]
24 where
25 single = Bond {delocNum = 2, atomIDs = Nothing}
26 aromatic = Bond {delocNum = 6, atomIDs = Just [1,2,3,4,5,6]}
Listing 8: A representation of the molecule benzene expressed in the Molecule data type, written as a .hs file, and using list comprehension to ensure readability, excluding the Shells currently.
The provided code demonstrates the representation of the benzene molecule using the proposed algebraic data type (ADT) system. Benzene is a classic example of an aromatic compound, consisting of a ring of six carbon atoms with alternating single and double bonds (or delocalised electrons around a ring), and six hydrogen atoms attached to each carbon.
In this representation, the benzene molecule is defined as a value of type Molecule. The atoms field contains a list of 12 Atom values, representing the six carbon atoms and six hydrogen atoms in the molecule. Each atom is specified with its unique atomID, atomicAttr (including the AtomicSymbol, atomicNumber, and atomicWeight), coordinate (3D spatial position), and shells (electronic configuration).
The carbon atoms (atomIDs 1-6) are positioned in a hexagonal arrangement, with coordinates that define the vertices of the hexagon. The hydrogen atoms (atomIDs 7-12) are positioned radially outward from each carbon atom, with coordinates that place them at a typical carbon-hydrogen bond distance.
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as Bond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6].
The bonds between the carbon atoms and their respective hydrogen atoms are also represented as Bond values, but with a delocNum of 2, indicating a typical single bond between carbon and hydrogen. The atomIDs field for these bonds is therefore set to Nothing, as they are not part of the aromatic ring.
The code also includes the definitions of elementAttributes for hydrogen (H), nitrogen (N), and carbon (C). These elementAttributes specify the AtomicSymbol, atomicNumber, and atomicWeight for each element. For example, elementAttributes C = ElementAttributes C 6 12.011 represents the carbon element with an atomic number of $6$ and an atomic weight of $12.011$. These predefined elementAttributes are used in the atomicAttr field of each Atom in the benzene molecule, providing the necessary information about the elemental properties of the atoms. The representation could be extended to represent reactions or polymers, without requiring a completely new and separate syntax like BigSMILES. Instead, lists whose syntax and semantics represents reflects repetition of subunits.
3.2 The Orbital ADT
The representation captures quantum mechanical properties of atoms and can be instantiated such that it adheres to the principles governing the arrangement of electrons in an atom. Each shell is represented as a distinct entity, characterized by its principal quantum number. This aligns with the quantum mechanical description of atomic structure, where electrons occupy discrete energy levels or shells. By organizing the electronic configuration into separate shells, the representation accurately reflects the hierarchical nature of atomic orbitals and the increasing energy and complexity of the electronic structure as the principal quantum number increases.
The use of a list to represent the shells implies that there could be an arbitrarily large finite number of shells (or even infinitely many), which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
Within each shell, there are subshells ($s$, $p$, $d$, and $f$), which correspond to the different types of atomic orbitals. We do not include $g$ orbitals due to their rarity. Any interface to this data structure must respect the maximum number of electrons that can occupy each subshell, with $s$ subshells containing up to two electrons, $p$ subshells up to six electrons, $d$ subshells up to ten electrons, and $f$ subshells up to fourteen electrons. However, since the current implementation does not utilize dependent types, the correctness of the number of electrons in each subshell cannot be guaranteed at compile time and is left as future work. The responsibility of ensuring the proper electronic configuration lies with the user of the ADT.
Real atoms fill orbitals according to the interplay of $n$ and $l$ (the Aufbau principle), but well-known exceptions exist (e.g., chromium and copper) that reflect subtler electron–electron interactions. Spin is also relevant: each orbital can hold two electrons due to spin ±$\pm\frac{1}{2}$, yet here the ADT simply stores a count instead of tracking each electron’s spin. Finally, the $f$-orbital enumerations (e.g., Fxxx) are reasonable labels for real combinations of eigenstates, though in practice chemists often use different conventions or simply label them by magnetic quantum numbers.
Orbital hybridisation is also explicitly supported in this representation through the inclusion of the hybridComponents field in the Orbital data type. This feature not only allows for the accurate modeling of hybrid orbitals (such as sp^3, sp^2, etc.) but also enables predictions of chemical properties like acidity, as hybridisation can influence the electron density distribution and thus the reactivity of atoms in molecules.
Each atomic orbital contains the subshell type, electron count, and spatial orientation of the orbital. These properties can be used to construct the actual basis functions, but they are not the basis functions themselves. Furthermore, the representation allows for the specification of the spatial orientation of each orbital within a subshell through the use of the Coordinate type. This information is crucial for describing the shape and symmetry of atomic orbitals, particularly for $p$, $d$, and $f$ orbitals that have distinct orientations. However, it is important to note that the actual spatial representation of the orbitals would require the use of appropriate mathematical functions, such as Slater-type orbitals (STOs), spherical harmonics, or Gaussian functions. Slater-type orbitals are commonly used in quantum chemistry as they provide a more accurate representation of the electron density near the nucleus compared to Gaussian functions. However, these mathematical functions are not explicitly included in this ADT system, which focuses on the symbolic representation of the electronic configuration.
By designing the appropriate interface functions that operate on the Shells type, one can ensure that rules such as the Aufbau principle, Hund’s rule, and the Pauli exclusion principle are properly enforced when manipulating the electronic configuration. These interface functions would act as a layer of abstraction, encapsulating the logic and rules governing the distribution of electrons within the atom. By providing a well-defined set of operations that adhere to these principles, the interface can maintain the integrity and consistency of the electronic configuration, thereby accurately representing the charge of the molecule. The interface functions would handle the creation, modification, and querying of the Shells data structure, ensuring that any changes made to the electronic configuration are valid and compliant with quantum mechanical principles. This approach allows for a clear separation of concerns, where the Shells type focuses on data representation while the interface functions handle the logic and enforcement of the principles governing the electronic configuration.
The Aufbau principle can be enforced by ensuring that electrons are added to orbitals in the correct order of increasing energy. Hund’s rule can be implemented by maximizing the number of unpaired electrons in degenerate orbitals before pairing them. The Pauli exclusion principle can be enforced by limiting the maximum number of electrons in each orbital to two, with opposite spins (which could be implemented given more time). The use of type information can greatly assist in enforcing these principles and ensuring the validity of the electronic configuration. For example, the SubShell data type could be parameterized by the maximum number of orbitals allowed for each subshell type. This type-level information would prevent the creation of invalid subshells with more orbitals than permitted by quantum mechanical rules. Similarly, the electronCount field of the Orbital data type can be constrained to a range of values (e.g., 0 to 2) using type-level naturals or refined types. This ensures that the Pauli exclusion principle is upheld and no orbital contains more than two electrons.
1-- Basic orbital types
2data So = So deriving (Show, Eq, Read)
3
4data P = Px | Py | Pz deriving (Show, Eq, Read)
5
6data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read)
7
8data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
9
10-- A type to wrap one of the pure orbital types,
11-- useful when describing the components of a hybrid orbital.
12data PureOrbital = PureSo So
13 | PureP P
14 | PureD D
15 | PureF F
16 deriving (Show, Eq, Read)
17
18-- The Orbital type includes an extra field for hybrid components.
19-- For a pure orbital, hybridComponents is Nothing.
20data Orbital subshellType = Orbital
21 { orbitalType :: subshellType
22 , electronCount :: Int
23 , orientation :: Maybe Coordinate
24 , hybridComponents :: Maybe [(Double, PureOrbital)]
25 } deriving (Show, Eq, Read)
26
27-- A SubShell is a list of Orbitals all having the same subshell type.
28newtype SubShell subshellType = SubShell
29 { orbitals :: [Orbital subshellType]
30 } deriving (Show, Eq, Read)
31
32-- A Shell has a principal quantum number and may contain s, p, d, and f subshells.
33data Shell = Shell
34 { principalQuantumNumber :: Int
35 , sSubShell :: Maybe (SubShell So)
36 , pSubShell :: Maybe (SubShell P)
37 , dSubShell :: Maybe (SubShell D)
38 , fSubShell :: Maybe (SubShell F)
39 } deriving (Show, Eq, Read)
40
41-- A Shells type is just a list of Shell.
42type Shells = [Shell]
Listing 9: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells. Note some of the types here are not in blue, for simplicity.
However, the current design does not explicitly account for the spin of electrons or the coupling between orbital and spin angular momenta, which play a significant role in determining the detailed electronic structure of atoms, particularly for elements with partially filled subshells.
Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbital with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
As mentioned, in Haskell, algebraic data types (ADTs) are constructed using sum and product type constructors, analogous to addition and multiplication in algebra (see Type Arithmetic). A sum type is a choice between different types (denoted by the | symbol), akin to an or; for example, a Maybe type represents a value that may be Nothing (like zero in addition) or Just a where a is any type. This is reflective of a disjoint sum in algebra, where a value can be one thing or another, but not both simultaneously. This allows the grammar to distinguish between empty values and zero values. As will be discussed, the orientation field of the Orbital data type is wrapped in a Maybe type, which naturally expresses whether an orbital has a specific orientation or not. For $s$ orbitals, the orientation is Nothing because $s$ orbitals are spherically symmetric and have no spatial orientation. In contrast, $p$ orbitals have specific orientations, and their orientation can be represented as Just Coordinate, where Coordinate is a type that can represent orientation in the $x$, $y$, and $z$ axes. On the other hand, a product type combines several values into one composite type, similar to how factors multiply to a product; a tuple (a, b) is an instance of a product type containing both a and b.
Furthermore, if the orbitals are not known, one can use the keyword undefined if they are unknown. Whilst risky, one could also implement an Unknown data type, which would be future work.
3.2.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
1-- Hydrogen atom: 1s^1
2hydrogen :: Shells
3hydrogen =
4 [ Shell
5 { principalQuantumNumber = 1
6 , sSubShell = Just (SubShell
7 [ Orbital { orbitalType = So
8 , electronCount = 1
9 , orientation = Nothing
10 , hybridComponents = Nothing
11 }
12 ])
13 , pSubShell = Nothing
14 , dSubShell = Nothing
15 , fSubShell = Nothing
16 }
17 ]
18
19-- Carbon atom: 1s^2, 2s^2, 2p^2 (one electron in each of two p orbitals)
20carbon :: Shells
21carbon =
22 [ Shell
23 { principalQuantumNumber = 1
24 , sSubShell = Just (SubShell
25 [ Orbital { orbitalType = So
26 , electronCount = 2
27 , orientation = Nothing
28 , hybridComponents = Nothing
29 }
30 ])
31 , pSubShell = Nothing
32 , dSubShell = Nothing
33 , fSubShell = Nothing
34 }
35 , Shell
36 { principalQuantumNumber = 2
37 , sSubShell = Just (SubShell
38 [ Orbital { orbitalType = So
39 , electronCount = 2
40 , orientation = Nothing
41 , hybridComponents = Nothing
42 }
43 ])
44 , pSubShell = Just (SubShell
45 [ Orbital { orbitalType = Px
46 , electronCount = 1
47 , orientation = Just (Coordinate 1 0 0)
48 , hybridComponents = Nothing
49 }
50 , Orbital { orbitalType = Py
51 , electronCount = 1
52 , orientation = Just (Coordinate 0 1 0)
53 , hybridComponents = Nothing
54 }
55 ])
56 , dSubShell = Nothing
57 , fSubShell = Nothing
58 }
59 ]
Listing 10: Examples of electronic configurations for hydrogen, carbon, and oxygen using the Orbital Algebraic Data Type representation in Haskell.
For the hydrogen atom, the electronic configuration is modeled using a single shell with a principal quantum number of 1. This shell contains only an $s$ subshell (represented by the sSubShell field) consisting of one orbital. The orbital is defined with the following properties:
  • orbitalType = So, indicating an $s$ orbital.
  • electronCount = 1, meaning it is occupied by one electron.
  • orientation = Nothing, which reflects the spherical symmetry of $s$ orbitals (they have no preferred spatial direction).
  • hybridComponents = Nothing, showing that the orbital is pure (i.e., it is not a hybrid orbital).
This precisely encodes the ground state configuration of hydrogen as 1s1.
For the carbon atom, the electronic configuration is represented by two shells. In the first shell (with principal quantum number 1), there is an $s$ subshell containing an orbital with:
  • orbitalType = So,
  • electronCount = 2 (representing 1s2),
  • orientation = Nothing, and
  • hybridComponents = Nothing.
In the second shell (with principal quantum number 2), the configuration comprises:
  • An $s$ subshell with an orbital defined similarly (with electronCount = 2 for the 2s electrons), and
  • A $p$ subshell (pSubShell) containing two orbitals:
    • The first $p$ orbital is defined as Orbital Px 1 (Just (Coordinate 1 0 0)), meaning it is of type Px, carries one electron, and has its orientation specified along the $x$-axis.
    • The second $p$ orbital is defined as Orbital Py 1 (Just (Coordinate 0 1 0)), meaning it is of type Py, carries one electron, and is oriented along the $y$-axis.
Both $s$ and $p$ orbitals in carbon have their hybridComponents fields set to Nothing, indicating that they are pure orbitals and not hybridized. This representation encodes the configuration 1s2 2s2 2p2, with one electron occupying each of the two available $p$ orbitals.
Overall, the inclusion of the orientation and hybridComponents fields in each orbital allows the data structure not only to capture the electron occupancy and energy level (via the principal quantum number) but also to describe the spatial characteristics of the orbitals. This detailed representation is fundamental for applications such as predicting molecular acidity, where orbital hybridisation and electron spatial distribution significantly influence chemical reactivity and intermolecular interactions.
3.3 Reactions
The proposed ADT can be extended to model chemical reactions. By defining a Reaction data type that captures the reactants, products, and conditions under which a reaction occurs, we can create a tool for studying and simulating chemical processes.
Listing 11 illustrates a Reaction data type, where reactants and products are represented as lists of pairs, containing a stoichiometric coefficient (a Double) and a Molecule. The conditions field captures reaction conditions such as temperature and pressure, defined with respect to a time range. Listing 12 provides an example reaction, illustrating the formation of water from hydrogen and oxygen under specific thermodynamic conditions.
Although these examples are simple, they demonstrate the flexibility of the ADT by extending it to model reaction processes. The compositional nature and modularity of the type system ensures that additional features not implemented here, such as catalysts or thermal energy, could be straightforwardly implemented. Stoichiometric or thermodynamic constraints could be enforced through use of dependent types, although this is left as future work.
1-- Reactions or transformations between chemical species
2data Reaction = Reaction
3 { reactants :: [(Double, Molecule)]
4 , products :: [(Double, Molecule)]
5 , conditions :: [Condition]
6 , rate :: Double
7 }
8
9-- Conditions under which a reaction occurs
10data Condition = TempCondition {temperature :: Double}
11 | PressureCondition {presssure :: Double}
12
13data Times = Times { startTime :: Double, endTime :: Double }
Listing 11: Haskell representation of a chemical reaction including reactants, products, conditions, and reaction rate. The Reaction data type is composed using the Molecule data type
1-- Example reaction: 2H2 + O2 -> 2H2O
2exampleReaction :: Reaction
3exampleReaction = Reaction
4 { reactants = [(2.0, undefined), (1.0, undefined)]
5 , products = [(2.0, undefined)]
6 , conditions =
7 [ TempCondition 500.0
8 , PressureCondition 1.0
9 ]
10 , rate = 0.1
11 }
Listing 12: Example reaction: formation of water from hydrogen and oxygen
In contrast, string-based representations such as reaction SMILES, SMARTS, and SMIRKS attempt to encode reactions as sequences of characters [12]. Reaction SMILES, for example, separates reactants and products using the “¿¿” operator, but does not inherently encode reaction conditions, catalysts, or rate dependencies [13]. SMARTS and SMIRKS provide pattern-matching capabilities for substructure searching and reaction templates, respectively, but lack a native mechanism for enforcing reaction validity [12].
String-based representations also inherit the limitations of the string data type itself. A lack of compositionality, and difficulties in formal reasoning make it unclear how one might extend SELFIES, for instance, to define a reaction SELFIES in which all syntactically valid strings correspond to chemically valid reactions [1]. Even if such an extension were possible, enforcing network-wide properties such as reversibility, equilibrium conditions, or pathway constraints would be nontrivial.
To generalise what is written here to adequately capture the semantics of reactions and reaction networks, it would be useful to derive an implementation off of a categorical semantics such as [63]. However, this is left as future work. Chemical reaction networks align naturally with category-theoretic formalisms, where reactions can be treated as morphisms between reactant and product sets [63]. This view has been used to model biochemical pathways, open dynamical systems, and Petri nets in systems chemistry [63], making an ADT in a typed functional language with support for categorical constructs such as Functors etc, an obvious choice for representation.
While constraints such as stoichiometric balance are not enforced in our implementation, the ADT provides a foundation upon which such constraints could be incorporated in future work through dependent types or additional verification layers.
However, the compositional nature of the ADT, its ability to be amended to incorporate reaction properties, and its alignment with category theoretic formalisms make it a natural framework cheminformatics applications.
3.4 Group Theory
Moreover, the ADT can be instantiated as a group to incorporate group theory concepts, which play a crucial role in understanding molecular symmetry and its impact on chemical properties. Exploiting these group symmetries can aid neural networks by the use of group equivariance and group invariance [2]. Invariances can be used also with atomic IDs meaning that the same atom represented with different ID’s can be quotiented and so treated as the same, allowing less data to be needed in order to train a neural network, for instance. In particular, molecules can be made instances of rotation groups, such as the cyclic group $C3$, which represents the rotational symmetry of molecules with threefold rotational axes.
In the following code snippet, a MolRot consists of a Molecule, a rotation axis specified by a Coordinate, and a rotation angle (in radians). The Group typeclass defines the necessary operations for a group, namely the group operation (mul), the inverse operation (inv), and the identity operation (e), and associativity can be proved but is implicit.
By making MolRot an instance of Group, we can combine rotations using the mul function, find the inverse of a rotation using inv, and obtain the identity rotation using e. This allows for the study of molecular symmetry and its implications on chemical properties and reactivity.
However, it’s important to note that proving that a data type satisfies all the group axioms (closure, associativity, identity, and inverse) would require the use of dependent types, which is beyond the scope of this representation. Nonetheless, by defining the necessary operations and providing sensible implementations, we can still leverage the power of typeclasses in representing these properties.
The following provides an example of how a MoleculeRotation data type can be defined and made an instance of the Group typeclass, though the full implementational details are not included due to their complexity.
1class Group g where
2 mul :: g -> g -> g -- Group operation: combining two operations
3 inv :: g -> g -- Inverse of an operation
4 e :: g -> g -- Identity operation
5
6data MolRot = MolRot Molecule Coordinate Double deriving (Show)
7
8instance Group MolRot where
9 mul (MolRot mol1 axis1 angle1) (MoleculeRotation mol2 axis2 angle2) =
10 combineRotations mol1 axis1 angle1 axis2 angle2
11
12 inv (MoleculeRotation mol axis angle) = MolRot mol axis (-angle)
13
14 e (MoleculeRotation mol axis angle) = MolRot mol (Coordinate 0 0 0) 0
Listing 13: An instance of making a MolRot a group which has a rotation axis and a rotation angle.
3.5 Storage, Transmission, and Readability
As well the computational purposes discussed elsewhere in the article, the ADT provides advantages with respect to storage of the molecule, and transmission between cheminformaticians.
Molecules in the ADT can be serialised and stored as .hs files, which may be used for storage, and for transmission of molecular data, perhaps via large chemical databases, through straightforward use of the Show and Read instances in Haskell, as shown in Listing 15.
Additionally, as .hs files will be interpreted with Haskell, molecules can be straightforwardly compressed as Listing 14 demonstrates by using foldr to compress the bondsData list of the benzene molecule. Although this may raise the issue of “canonicalisation” for storage of molecules, it should be noted the bondsData in Listing 14 and Listing 8 will be stored the same way inside of a Haskell program, as the corresponding data is the same in each case. This could be beneficial in the representation of large molecules, as only a small number of bonds or atoms need be explicitly described to encode the molecule for storage.
1bondsData :: [((Int, Int), Bond)]
[…]
5 where
6 single = Bond {delocNum = 2, atomIDs = Nothing}
7 aromatic = Bond {delocNum = 6, atomIDs = Just [1,2,3,4,5,6]}
Listing 14: Using foldr to compress the bondsData of the molecule benzene in the ADT representation. As Benzene is symmetric, the bonding system can be specified by only describing three bonding systems, and using those to construct the bondsData list.
1-- Writing the molecule to a file
2writeMoleculeToFile :: FilePath -> Molecule -> IO ()
3writeMoleculeToFile filePath molecule = writeFile filePath (show molecule)
4
5-- Reading the molecule from a file
6readMoleculeFromFile :: FilePath -> IO Molecule
7readMoleculeFromFile filePath = do
8 contents <- readFile filePath
9 return (read contents)
10
11-- Example usage
12main :: IO ()
13main = do
14 -- Write the methane molecule to a file
15 writeMoleculeToFile "methane.hs" methane
16
17 -- Read the methane molecule from the file
18 molecule <- readMoleculeFromFile "methane.hs"
19
20 -- Print the molecule read from the file
21 print molecule
Listing 15: Haskell code demonstrating the serialisation and writing to disk of a Molecule data type and the reading of a molecule stored in an .hs file, through use of the show and read functions.
This code demonstrates the ease of serializing and deserializing molecules to and from files in Haskell. By leveraging the Show and Read instances defined for the Molecule data type and its related types, molecules can be seamlessly converted to and from their string representations, allowing them to be stored in files and retrieved later.
This functionality provides a convenient way to save and load molecular data, and is particularly useful when working with large molecules or when molecules are required across different program runs. Although storage and transmission may take place via serialisation into a human and machine-readable string of text, it is important to distinguish between the serialized format of the data and the underlying data type itself. While the ADT can be serialized into a text string, it is crucial to recognize that the ADT itself is not merely a string.
3.6 A simple example of probablistic programming
Some code is excluded from listing 16 to conserve space, but is explained below.
This module implements a probabilistic generative model for molecules, where the goal is to generate a candidate molecule that is scored according to its similarity to an observed molecule. The observed molecule is obtained by reading an SDF file using the function observedMoleculeIO, which invokes parseDB1File to parse the file "./molecules/water.sdf". If no molecule is found, the function terminates with an error; otherwise, the first molecule is returned.
The core of the generative process is encapsulated in the function moleculeModel, which takes an observed Molecule as input and returns a computation in the Meas monad. Within this function, three atoms are generated by executing a monadic loop (via replicateM). For each atom, a chemical element is randomly selected from the set [C, N, O, H] using the uniformD distribution, and a three-dimensional coordinate is sampled from independent normal distributions by means of the helper function sampleCoordinate. The attributes associated with the chosen element, namely elementAttributes and elementShells, are then used to construct an Atom with a placeholder identifier (initially 0). After all atoms are generated, unique identifiers (1, 2, 3) are assigned to them.
Subsequently, the code computes all unique unordered pairs of atom identifiers (which, in the case of three atoms, yields the pairs (1,2), (1,3), and (2,3)) using list comprehensions and the function tails. For each pair, the code samples a Boolean value from a uniform distribution over [True, False] to decide whether a bond should be included between the corresponding atoms. When a bond is to be added, a bond order is sampled uniformly from the set [1, 2, 3] and is converted into a Bond record by setting its field delocNum to 2, 4, or 6, respectively. The bond is then stored in a map (of type M.Map (Integer, Integer) BondType) using a key that is the sorted pair of atom identifiers, thereby ensuring the symmetric nature of the bond representation.
Once both the atoms and bonds have been generated, a Molecule is constructed. Its quality is evaluated by computing the Hausdorff distance between the generated molecule and the observed molecule via the function hausdorffDistance. This distance is defined as the maximum over the minimum Euclidean distances between each atom of one molecule and the atoms of the other, where the Euclidean distance is computed by the function euclideanDistance. The resulting distance is then used to score the molecule by applying a normal probability density function (normalPdf) with mean 0 and standard deviation 1, through the score function.
The main function ties these components together. It first obtains the observed molecule using observedMoleculeIO, and then it executes a Metropolis–Hastings sampling procedure (via mh with a step size of 0.1) on the generative model instantiated with the observed molecule. After discarding an initial burn-in of 1000 samples, 1000 samples along with their corresponding weights are extracted and printed to standard output.
1moleculeModel :: Molecule -> Meas Molecule
2moleculeModel observed = do
3 let numAtoms = 3
4 atomsUnnumbered <- replicateM numAtoms \$ do
5 symbol <- sample \$ uniformD [C, N, O, H]
6 coord <- sampleCoordinate
7 let attr = elementAttributes symbol
8 shellsVar = elementShells symbol
9 return \$ Atom { atomID = 0, atomicAttr = attr, coordinate = coord, shells = shellsVar }
10
[…]
12 atomIDs = map atomID atoms
13
14 let possiblePairs = [(i, j) | (i:rest) <- tails atomIDs, j <- rest]
[…]
16 includeBond <- sample \$ uniformD [True, False]
17 if includeBond
18 then do
19 bondOrder <- sample \$ uniformD [1, 2, 3]
20 let bondType = case bondOrder of
21 1 -> Bond { delocNum = 2, atomIDs = Nothing }
22 2 -> Bond { delocNum = 4, atomIDs = Nothing }
23 3 -> Bond { delocNum = 6, atomIDs = Nothing }
24 _ -> error "Invalidbondorder"
25 return \$ Just (pair, bondType)
26 else return Nothing)
27 possiblePairs
28
29 let bondsList = catMaybes bondsMaybe
30 bondsMap = M.fromList bondsList
31 let molecule = Molecule { atoms = atoms, bonds = bondsMap }
32 let distance = hausdorffDistance molecule observed
33
34 score \$ normalPdf 0 1 distance
35 return molecule
36
37sampleCoordinate :: Meas Coordinate
38sampleCoordinate = do
39 x <- sample \$ normal 0 1
40 y <- sample \$ normal 0 1
41 z <- sample \$ normal 0 1
42 return \$ Coordinate x y z
43
44main :: IO ()
45main = do
46 observed <- observedMoleculeIO
47 samples <- mh 0.1 (moleculeModel observed)
48 let (molecules, weights) = unzip \$ take 1000 \$ drop 1000 samples
49 putStrLn \$ "Sampledmolecules:" ++ show molecules
50 putStrLn \$ "Weights:" ++ show weights
Listing 16: An example probabilistic program, where trace Metropolis-Hastings (M-H) operates by sampling entire “traces” of latent variables (i.e., all random choices) from a model [47]. Each step proposes a small perturbation to the current trace, controlled by a jitter parameter, often the standard deviation of the proposal distribution. This parameter must be tuned carefully to ensure efficient exploration of the posterior while maintaining a reasonable acceptance rate. Typically, an initial burn-in period is applied, during which early samples (which have not yet converged) are discarded. This burn-in helps remove the influence of the initial point on the chain, leading to samples that more accurately reflect the stationary distribution of the target posterior. Here the jitter is set to 0.1 and the number of burnin samples is 1000.
1euclideanDistance :: Coordinate -> Coordinate -> Double
2euclideanDistance (Coordinate x1 y1 z1) (Coordinate x2 y2 z2) =
3 sqrt \$ (x1 - x2)^2 + (y1 - y2)^2 + (z1 - z2)^2
4
5hausdorffDistance :: Molecule -> Molecule -> Double
6hausdorffDistance mol1 mol2 =
7 let coords1 = map coordinate (atoms mol1)
8 coords2 = map coordinate (atoms mol2)
9 d1 = [minimum [euclideanDistance c1 c2 | c2 <- coords2] | c1 <- coords1]
10 d2 = [minimum [euclideanDistance c2 c1 | c1 <- coords1] | c2 <- coords2]
11 in max (maximum d1) (maximum d2)
Listing 17: Euclidean and Hausdorff distance calculations for molecular coordinates
In summary, the code defines a complete pipeline in which a molecule is generated by randomly selecting atom types, coordinates, and bonds, and the quality of the generated molecule is measured against an observed molecule using the Hausdorff distance. The sampling procedure then explores the space of possible molecules according to the probabilistic model defined.
3.6.1 Meas & Prob Monads
A monad is an abstraction that represents computations as sequences of steps, encapsulating context such as side effects, state, or randomness. In Haskell, a monad is defined by the operations return (to inject a value into the monadic context) and bind (represented as >>=), which chains computations while automatically managing the underlying context. This structure allows for composability and abstraction of computational effects, enabling programmers to write clean, modular code that implicitly threads additional information—be it state, failure, or randomness—throughout a computation.
The Prob monad is used for representing probability distributions, while the Meas monad is used for representing unnormalized probability distributions (measures). The sample function is used to sample from a probability distribution (Prob), and the score function is used to score the model based on some condition.
In the presented code, the Prob monad encapsulates a probability distribution by representing a computation as a function that consumes an infinite, lazily generated Tree of random numbers to yield a result. This mechanism threads random choices through the computation by splitting the tree, thereby providing a deterministic view over randomness. Building on this, the Meas monad is defined as a wrapper around a WriterT transformer applied to Prob, which accumulates a log-weight (using the Product (Log Double) type) alongside the result. This dual nature enables the Meas monad to not only sample outcomes (via the sample function) but also to incorporate scoring (via the score function), which adjusts the unnormalized probability measure. As a result, Meas serves as a framework for building and manipulating probabilistic models that can be employed by inference algorithms such as likelihood weighted importance sampling and Metropolis-Hastings.
3.7 Limitations
The storage efficiency of the should be further explored. The ADT syntax could result in large file sizes for molecules stored as .hs files, however, the use of list comprehension, and the fact that molfiles and the Molecule data type belong to the same complexity class, may mean that the ADT representation may in fact be smaller than some mol files, for the same degree of expressiveness. By storing a molecule as function or properties as higher-order functions, it may be possible to significantly reduce the size of the molecule.
3.8 Future work
The ADT so far presented has the ability to be improved upon, extended, and utilised in empirical experiments in order to verify and test the representation’s utility across a number of domains in cheminformatics. It would be useful to have parsers from multiple other representations e.g. SMILES strings to this ADT. Currently only parsers for .sdf files have been written.
3.8.1 Augmentation of the ADT
To improve the representational ability of the ADT, it should be augmented to enable representation of tautomerism, ionic bonding, polymerism, and quantum information such as spin, none of which are represented in the current ADT.
Tautomers are in medications from sildenafil (Viagra) [64], to Warfarin, to Remdesivir (used for Covid 19), Tetracyclines and antibiotics. Representation of tautomers could allow exploration of a greater chemical space, and facilitate tasks such as prediction of ligand-protein binding, necessary for many drug design tasks[65, 66].
Polymers are important in many areas of chemistry and materials science. By using lists and recursive data structures, the representation could be augmented to handle polymeric molecules. This extension would enable the modeling and analysis of complex macromolecular systems, opening up new possibilities for polymer design and optimization.
The inclusion of ionic bonding to the ADT would be another significant enhancement. Ionic interactions are critical for the stability and functionality of nucleic acids, proteins, and membranes, as well as for processes like catalysis and ion transport [67, 68, 69]. Augmenting the ADT to represent ions and their bonding explicitly would enable the modeling of charged species and their interactions. This could also facilitate the study of electrochemical properties and enable the design of materials such as ionic liquids and solid electrolytes, which are essential for applications ranging from energy storage to green chemistry [70].
The extension of the ADT to include quantum mechanical properties such as spin and angular momentum represents a significant opportunity. Quantum mechanical features are crucial for accurately modeling the electronic structure and predicting molecular reactivity, as they account for phenomena like electron spin coupling and charge distribution, which play a key role in determining molecular properties [13, 71]. By incorporating these properties into the ADT, it could better capture the behavior of molecules in complex environments, including those encountered in catalysis and photochemistry. Such advancements would also enable the ADT to align more closely with the requirements of quantum chemistry simulations, which often rely on accurate spin and orbital information for predicting reaction mechanisms and designing new materials [12, 1].
3.8.2 Dependent Types
Furthermore, to enforce constraints, and guide programmers in development, dependently-typed functional programming languages, as previously defined, can be used. These languages (e.g. [72]) have many advantages including:
  • Eliminating Physically Impossible Molecules: By enforcing valency constraints at the type level, we can prevent the creation of molecules where atoms exceed their natural bonding capabilities (e.g., a carbon atom with five single bonds).
  • Ensuring Charge Balance: Dependent types can encode rules ensuring that molecules do not have unbalanced charges unless explicitly permitted (e.g., handling ionic compounds separately).
  • Constraining Molecular Topology: We can restrict bond formations to valid molecular configurations, preventing unphysical connectivity (e.g., an atom being bonded to itself).
  • Enforcing Electronic Configuration Validity: By parameterizing orbitals and electron shells with dependent types, we can ensure that electrons are correctly distributed within allowed quantum states.
3.8.3 Development of powerful, user-friendly library
To make the most of the ADT, a powerful, user-friendly library of functions should to be created. This could be parallelised with high performance Haskell libraries such as Accelerate (https://hackage.haskell.org/package/accelerate). Functions should include the ability to convert between the ADT and other common representations, which would enable use of the ADT with existing potentially vast chemical databases.
Canonicalisation is also a concern for the ADT, as molecules may have different indices, and there may be differing atomic coordinates. The permuations of a set of indices, would however form a group, and the group properties of the ADT can be leveraged for this purpose.
3.8.4 Futher enhancing the ADT
To further strengthen the type safety and invariance guarantees of the representation, exploring dependently typed programming languages, such as Agda, could be highly beneficial. Dependent types allow for the encoding of more sophisticated invariants and constraints directly into the type system, ensuring that only valid and consistent molecular structures are expressible. This increased level of type safety can help catch potential errors at compile-time and provide stronger guarantees about the correctness of the representation.
Further work could explore several avenues to enhance the proposed algebraic data type (ADT) representation and its applications. Integrating Haskell’s Lens library [73] could simplify the manipulation of complex molecular structures by providing a composable, modular and type-safe way to access and modify nested data fields.
To enhance the modularity and composability of the ADT representation, exploring techniques like modular syntax trees and the “Data Types a la Carte” approach [74] could prove highly valuable. By decomposing the molecular representation into smaller, reusable components, we can create a more flexible and adaptable framework. This modular design would allow for the easy integration of new features, the creation of domain-specific languages for molecular manipulation, and the development of reusable libraries and tools for cheminformatics.
By pursuing these avenues of future work, we can further strengthen the capabilities and thus impact the community, encouraging them to use the ADT-based molecular representation, hopefully accelerating the discovery of novel and useful molecules. Implementing backpropagation using differentiable programming [75] would enable using gradient based inference algorithms.
Integrating the ADT with quantum sampling techniques could lead to breakthroughs in predicting molecular properties by efficiently simulating complex quantum systems. However, realizing this goal would require close collaboration between experts in quantum computing, chemical informatics, and molecular modeling to develop efficient and correct quantum algorithms.
3.8.5 Experimental validation of the ADT
At present, the efficiency and efficacy of the ADT in cheminformatics tasks are conjectures. To validate the utility of the ADT as a molecular representation, empirical work is needed to benchmark its performance against representations such as SMILES and SELFIES, on tasks like molecular property prediction (e.g. calculation of logP and logS), virtual screening, and de novo drug design.
Another avenue for future work is utilising the proposed representation with geometric deep learning techniques to exploit symmetries and invariances in molecular structures [2]. By leveraging mathematical properties of ADTs, such as group invariances and equivariances and their verification in Haskell through typeclasses, one could develop neural network architectures that are specifically tailored to the unique characteristics of those molecules. For instance, the choice of index for each atomID used in the representation should not influence a model, given the same molecule but with a different ordering of atomic IDs.
An additional avenue is using equational reasoning to investigate algebraic approaches in molecular modeling [76].
Neural networks can be designed to be invariant to permutations of atomic indexes, ensuring that molecularly equivalent structures are treated identically regardless of labeling of indices or bonds. Similarly, molecular features such as the atomSymbol, are both translation and rotation invariant, while others may only be translation invariant. Incorporating these symmetries directly into the neural network architecture could significantly reduce the amount of training data required and improve the generalization capability of the models, but this would have to be demonstrated.
4 Conclusion
In this work, we presented a novel approach to molecular representation using algebraic data types (ADTs) embedded in the purely functional, type-safe, programming language Haskell.
We discussed contemporary representations, focusing primarily on the string based representations SMILES and SELFIES, and their short-comings for the use in machine learning.
By implementing the ADT using the Dietz representation for constitution of molecules, and atomic coordinate data for 3D information, the ADT greatly expands the chemical space which can be represented, and hence explored in cheminformatics tasks when compared to existing representations. In addition, the ADT allows the optional inclusion of quantum mechanical information, such as orbitals and electronic configurations, providing a level of detail unmatched in current representations.
As a general model for molecular representation, the ADT can represent complex, multi-centre, electron-deficient and delocalised bonding, as well as advanced stereoechemical features, and “zero order” bonds such as ionic and hydrogen bonds, without requiring extensive labeling for new bond types or oxidation states.
Through Haskell’s type classes and functional programming paradigm, we provide a rich and robust semantic context for molecular structure, and an efficient framework for computational tasks. Molecular data adheres to the constraints of the molecular representation, going beyond the limitations of strings as a representation.
To demonstrate the ADT and how it can be extended, we explored group based properties of molecular symmetry. We also showed integration of the ADT with probabilistic programming, showing that the ADT can be appropriately used for Bayesian inference, and that machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation to other formats.
The ADT presented was not intended to be a definitive representation, but a valuable representational concept for cheminformaticians. Nevertheless the ADT offers a powerful and flexible alternative to existing string-based representations like SMILES and SELFIES, for the representation of, manipulation of, and reasoning about molecular structures in cheminformatics.
5 Declarations
5.1 Availability of data and materials
All code used to implement the Algebraic Data Type and the experiments in this article is available via github, at https://github.com/oliverjgoldstein/Mol-Rep-Haskell-Bayes. In line with open source molecular software [77], we have submitted a pull request to https://opensourcemolecularmodeling.github.io/.
5.2 Competing interests
The authors declare no competing interests.
5.3 Funding
Oliver Goldstein is funded by an EPSRC DTP Scholarship.
Sam March is self-funded.
5.4 Authors’ contributions
5.4.1 Oliver Goldstein
Ideated the project and conceptualized the solution in the form of an Algebraic Data Type, wrote the code base of the library with all experiments and extensions and drafted the initial article
5.4.2 Samuel March
Made substantial revisions to the article’s form and content, contributed to the research, and made misc. contributions to the library
5.5 Acknowledgments
My thanks to my supervisor Jeremy, for putting up with me (easier said than done). We are also thankful for discussions at the Algebra of Programming group at Oxford University.